### v2

# Quantitative estimates for singularity for conjugate equations driven by linear fractional transformations

Abstract.
We consider the conjugate equation driven by two families of finite maps on the unit interval satisfying a compatibility condition. This framework contains de Rham’s functional equations. We give sufficient conditions for singularity of the solution with quantitative estimates in the case that the equation is driven by non-affine maps, in particular, linear fractional transformations.
1. Introduction
Let $N\geq 2$ and let $\mathcal{I}_{N}:=\{0,1,\cdots,N-1\}$. Let $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ be two families of $N$ maps on $[0,1]$ satisfying a compatibility condition. We consider the following conjugate equation for $\varphi:[0,1]\to[0,1]$:
$g_{i}\circ\varphi=\varphi\circ f_{i},\ i\in\mathcal{I}_{N}.$ (1.1)
See Section 2 below for the definition of the compatibility condition. An alternative expression is given by
$\varphi(x)=\begin{cases}g_{0}(\varphi(f_{0}^{-1}(x)))&0\leq x\leq f_{0}(1)\\ \vdots&\\ g_{N-1}(\varphi(f_{N-1}^{-1}(x)))&f_{N-1}(0)\leq x\leq 1\end{cases}.$ (1.2)
This is known as de Rham’s functional equation [7, 8]. This functional equation is generally fractal in nature, and solutions of (1.1) can be singular functions.
Assume that $N=2$, $10<p<$ 1, $f_{0}(x)=x/2$, $f_{1}(x)=(x+1)/2$, $g_{0}(x)=px$ and $g_{1}(x)=(1-p)x+p$. Then, (1.2) is expressed as
$\varphi(x)=\begin{cases}p\varphi(2x)&0\leq x\leq 1/2\\ (1-p)\varphi(2x-1)+p&1/2\leq x\leq 1\end{cases}.$
There exists a unique continuous strictly increasing solution $\varphi_{p}$ for this equation, and $\varphi_{p}$ is singular if $p\neq 1/2$ and $\varphi_{p}(x)=x$ if $p=1/2$. $\varphi_{p}$ is called the Lebesgue singular function. This function is also a solution of the functional equation by Matkowski and Wesołowski [16, 24, 25].
Minkowski’s question-mark function [23] is also obtained by a solution of (1.1) for the case that $(f_{0}(x),f_{1}(x),g_{0}(x),g_{1}(x))=\left(x/(x+1),1/(2-x),x/2,(x+1)/2\right)$. It can also be expressed in the form
$\varphi(x)=\begin{cases}\frac{1}{2}\varphi(\frac{x}{1-x})&0\leq x\leq 1/2\\ \frac{1}{2}\varphi\left(2-\frac{1}{x}\right)+\frac{1}{2}&1/2\leq x\leq 1\end{cases}.$
This function has appeared in various areas of mathematics such as number theory and dynamical systems. The real analytic properties of this function have been considered ([15, 19, 21]). The conjugate equation of the de Rham type has been investigated by many authors ([2, 4, 6, 11, 12, 13, 17, 28, 29, 30, 32]). We refer the reader to [27, Subsection 3.4] for a more detailed review of de Rham’s functional equations.
The main purpose of this paper is to investigate some properties of the solution of (1.1). Specifically, we give sufficient conditions for the solution to be singular. Since $\varphi$ is an increasing function on $[0,1]$, we deal with the probability measure $\mu_{\varphi}$, which is the Borel probability measure on $[0,1]$ whose distribution function is $\varphi$. We consider the case that both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are non-affine. This case is more difficult to analyze than the case that both $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are affine.
It is interesting to know a necessary and sufficient condition of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ for the singularity. This topic has been considered in [2, 3, 10, 13, 18, 25]. In [27], the author dealt with the case that $\{g_{i}\}_{i}$ are linear fractional transformations and $\{f_{i}\}_{i}$ are affine. Our results, Theorems 3.2 and 3.3 below, deal with the case that $\{f_{i}\}_{i}$ are not necessarily affine. We do not assume the differentiability of $\{f_{i}\}_{i}$. For the proof, we use a martingale approach as in [27], but here we need more delicate quantitative estimates.
The structure of this paper is as follows. In Section 2, we give some terminology and basic results of the conjugate equation on the unit interval. In Section 3, we deal with the singularity. Finally, we give an example in Section 4.
2. Preliminaries
2.1. Existence and uniqueness
We first discuss the existence and uniqueness of the solution of the conjugate equation (1.1).
We say that a family of maps $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ is a compatible system or satisfies a compatibility condition if they are strictly increasing continuous functions such that $h_{0}(0)=0$, $h_{i-1}(1)=h_{i}(0)$, $11\leq i\leq N-$ 1, $h_{N-1}(1)=1$, and, a compatible system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system on $[0,1]$ if
$\mathcal{D}_{h}:=\{h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j):i_{1},\cdots i_{n}\in\mathcal{I}_{N},j\in\{0,1\}\}$
is dense in $[0,1]$.
Lemma 2.1.
Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be a D-system on $[0,1]$. Then, for every $x\in[0,1]$, there exists an infinite sequence $(i_{n})_{n}$ of $\mathcal{I}_{N}$ such that
$x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1).$ (2.1)
If $x\in\mathcal{D}_{h}\setminus\{0,1\}$, then, there exist two such infinite sequences. Otherwise, there exists exactly one such infinite sequence.
Proof.
We first show that for every infinite sequence $(i_{n})_{n}$,
$\lim_{n\to\infty}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)|=0.$ (2.2)
Assume this fails. By the definition of $\{h_{i}\}_{i}$, $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right)_{n}$ is increasing and $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right)_{n}$ is decreasing. Since $\mathcal{D}_{h}$ is dense in $[0,1]$, there exists $k_{1},\cdots,k_{N}$ and $j\in\{0,1\}$ such that
$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)<h_{k_{1}}\circ\cdots\circ h_{k_{N}}(j)$
$<\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
This cannot occur. Hence (2.2) holds.
Case 1. Assume $x\in\mathcal{D}_{h}\setminus\{0,1\}$. Assume that $x=h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)$. Since $h_{0}(0)=0$ and $x>0$, we can assume that $i_{n}\geq 1$. Let $i_{k}:=0$ for every $k\geq n+1$. Then, (2.1) holds. By replacing $i_{n}$ with $i_{n}-1$ and let $i_{k}:=N-1$ for every $k\geq n+1$, then, (2.1) holds. Assume that $x=h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)$. Since $h_{N-1}(1)=1$ and $x<1$, we can assume that $i_{n}\leq N-2$. Let $i_{k}:=N-1$ for every $k\geq n+1$. Then, (2.1) holds. By replacing $i_{n}$ with $i_{n}+1$ and let $i_{k}:=0$ for every $k\geq n+1$, then, (2.1) holds.
Case 2. Assume $x\in\{0,1\}$. Assume $x=0$. Since each $h_{i}$ is strictly increasing, there is no other choice than $i_{k}=0$ for every $k\geq 1$ for (2.1). Assume $x=1$. Then, there is no other choice than $i_{k}=N-1$ for every $k\geq 1$ for (2.1).
Case 3. Assume $x\notin\mathcal{D}_{h}$. Then, for each $n\geq 1$, there exists a unique finite sequence $(i_{n,1},\cdots,i_{n,n})$ such that
$h_{i_{n,1}}\circ\cdots\circ h_{i_{n,n}}(0)<x<h_{i_{n,1}}\circ\cdots\circ h_{i_{n,n}}(1).$
Then, $i_{n-1,k}=i_{n,k}$ for every $k\leq n-1$. Let $i_{k}:=i_{n,k}$. By (2.2), (2.1) holds. If (2.1) holds, then, for every $N\geq 1$,
$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
Hence,
$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)<x<h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
Hence the uniqueness holds. ∎
Remark 2.2.
The existence and uniqueness of the bounded solutions of conjugate equations on more general spaces than $[0,1]$ have been discussed in [6, 26, 28, 30].
Lemma 2.3.
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are D-systems on $[0,1]$. Then, there exists a unique continuous strictly increasing solution of (1.1).
Hereafter, we let $\mu_{\varphi}$ be the continuous probability measure on $[0,1]$ whose distribution function is $\varphi$. This is the central object considered in this paper.
Proof.
We first see that
$\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(j)\right)=g_{i_{1}}\circ\cdots\circ g_{i_{n}}(j),\ i_{1},\cdots,i_{n}\in\mathcal{I}_{N},\ j\in\{0,1\},$ (2.3)
by induction in $n$. Hence, $\varphi(\mathcal{D}_{f})=\mathcal{D}_{g}$.
Step 1. We define $\varphi(x)$ for $x\notin\mathcal{D}_{f}$. By the assumption, there exists an infinite sequence $(i_{n})_{n}$ of $\mathcal{I}_{N}$ such that $x\in(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1))$. Since $\left[g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0),g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)\right]$ is decreasing with respect to $n$,
$\bigcap_{n\geq 1}\left[g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0),g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)\right]=\left[\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0),\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)\right].$
By the density of $\mathcal{D}_{g}$ in $[0,1]$, we can apply (2.2), and obtain that
$\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1).$ (2.4)
We let $\varphi(x):=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)$.
Step 2. We show $\varphi$ is strictly increasing on $[0,1]$. By (2.3), $\varphi$ is strictly increasing on $\mathcal{D}_{f}$. Let $x<y$. By (2.4), for some $n$,
$f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0)\leq x\leq f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)<f_{j_{1}}\circ\cdots\circ f_{j_{n}}(0)\leq y\leq f_{j_{1}}\circ\cdots\circ f_{j_{n}}(1).$
By this and Step 1,
$\varphi(x)\leq g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)<g_{j_{1}}\circ\cdots\circ g_{j_{n}}(0)\leq\varphi(y).$
Step 3. We show that $\varphi$ is continuous on $[0,1]$. Assume that there is a discontinuous point $x_{0}\in(0,1)$ of $\varphi$. Then,
$b_{0}:=\lim_{x\to x_{0}+0}\varphi(x)>\lim_{x\to x_{0}-0}\varphi(x)=:a_{0}.$
By Step 2, $\varphi([0,1])\cap(a_{0},b_{0})=\emptyset$, in particular, $\mathcal{D}_{g}\cap(a_{0},b_{0})=\emptyset$. This contradicts the density of $\mathcal{D}_{g}$ in $[0,1]$. The same argument is applied to the case that $x_{0}\in\{0,1\}$. ∎
It is natural to consider the inverse function of the solution $\varphi$. We easily see that
Lemma 2.4.
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are D-systems on $[0,1]$. Let $\varphi$ be a unique solution of (1.1). Then, the inverse function $\varphi^{-1}$ of $\varphi$ is well-defined and satisfies the following conjugate equation
$\varphi^{-1}\circ g_{i}=f_{i}\circ\varphi^{-1},\ i\in\mathcal{I}_{N},$
which is equivalent with (1.1) where $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are exchanged.
By [4, Proposition 3.1], if $\varphi$ is singular, then, $\varphi^{-1}$ is also singular.
2.2. Weak contraction
Hereafter, we assume more additional conditions for the D-systems $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$. We say that a function $h:[0,1]\to[0,1]$ is a contraction on $[0,1]$ if
$\textup{Lip}(h):=\sup_{x\neq y}\left|\frac{h(x)-h(y)}{x-y}\right|<1.$
$\textup{Lip}(h)$ is called the Lipschitz constant of $h$. We say that a function $h:[0,1]\to[0,1]$ is a weak contraction on $[0,1]$ if
$\phi_{h}(t):=\lim_{s\to t+0}\sup\left\{|h(x)-h(y)|:|x-y|\leq s\right\}<t,\ \ t>0.$
The following is well-known. See [5, 14].
Lemma 2.5.
Let $h$ be a weak contraction. Then, $\lim_{n\to\infty}\phi_{h}^{n}(t)=0$ for every $t>0$ and $|h(x)-h(y)|\leq\phi_{h}(|x-y|)$ for $x,y\in[0,1]$.
Proof.
Let
$\omega_{h}(s):=\sup\left\{|h(x)-h(y)|:|x-y|\leq s\right\}.$
Since this function is increasing, $\phi_{h}$ is right-continuous and increasing on $[0,1]$. Hence, $|h(x)-h(y)|\leq\lim_{s\to|x-y|+0}\phi_{h}(s)=\phi_{h}(|x-y|)$.
By the assumption, for every $t>0$, $(\phi_{h}^{n}(t))_{n}$ is strictly decreasing. There exists the limit $\phi^{\infty}_{h}(t):=\lim_{n\to\infty}\phi_{h}^{n}(t)$. By the right-continuity of $\phi_{h}$, $\phi_{h}\left(\phi^{\infty}_{h}(t)\right)=\phi^{\infty}_{h}(t)$. Hence, $\phi^{\infty}_{h}(t)=0$. ∎
Lemma 2.6.
Let $\{h_{i}\}_{i}$ be a compatible system on $[0,1]$ such that each $h_{i}$ is a weak contraction on $[0,1]$. Then, $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system.
We say that a compatible system $\{h_{i}\}_{i}$ on $[0,1]$ is an (weakly) contractive system on $[0,1]$ if each $h_{i}$ is an (weak) contraction on $[0,1]$. Every weakly contractive system is a D-system.
Proof.
Let $\phi:=\max_{i\in\mathcal{I}_{N}}\phi_{h_{i}}$. Then, $\phi$ is right-continuous [^1] The assumption that $N$ is finite would be important. on $[0,1]$, $\phi(t)<t$ for every $t>0$, and $\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\phi(|x-y|)$, $x,y\in[0,1]$. Then, we see that $\lim_{n\to\infty}\phi^{n}(t)=0$ for every $t>0$. Hence,
$\left|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right|\leq\phi^{n}(1).$
$\lim_{n\to\infty}\max_{i_{1},\cdots,i_{n}\in\mathcal{I}_{N}}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)|=0.$
By the definition, for each $n\geq 1$,
$[0,1]=\bigcup_{i_{1},\cdots,i_{n}\in\mathcal{I}_{N}}\left[h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0),h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right],$
and their interiors are disjoint. ∎
2.3. Linear fractional transforms
Definition 2.7.
We say that a family of functions $\{h_{i}\}_{i}$ is an LF system on $[0,1]$ if $h_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, for each $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers such that
(1) (well-defined) $c_{i}+d_{i}>0$, $d_{i}>0$, $i\in\mathcal{I}_{N}$.
(2) (strictly increasing, weak contraction)
$0<a_{i}d_{i}-b_{i}c_{i}\leq\min\{d_{i}^{2},(c_{i}+d_{i})^{2}\},\ i\in\mathcal{I}_{N}.$
(3) (compatibility) $b_{0}=0$, $\dfrac{a_{i}+b_{i}}{c_{i}+d_{i}}=\dfrac{b_{i+1}}{d_{i+1}}$, $i\in\mathcal{I}_{N}$, where we let $\dfrac{b_{N}}{d_{N}}:=1$.
For condition (2), we may consider a stronger condition that
(2’) $0<a_{i}d_{i}-b_{i}c_{i}<\min\{d_{i}^{2},(d_{i}+c_{i})^{2}\},\ i\in\mathcal{I}_{N}$.
Since $h_{i}$ is invariant if we multiply a positive constant by each of $a_{i},b_{i},c_{i}$ and $d_{i}$, we may assume that $d_{i}=1$, or, $a_{i}d_{i}-b_{i}c_{i}=1$.
Lemma 2.8.
Every LF system is a weakly contractive system. If (2’) holds instead of (2), then, it is a contractive system.
Proof.
Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF-system. We assume that $d_{i}=1$ for every $i$. By Lemma 2.6, it suffices to show that each $h_{i}$ is a weak contraction. We first remark that
$|h_{i}(x)-h_{i}(y)|=\frac{a_{i}-b_{i}c_{i}}{(c_{i}x+1)(c_{i}y+1)}|x-y|.$
We consider three cases: (i) $c_{i}>0$, (ii) $c_{i}<0$, and (iii) $c_{i}=0$.
(i) Assume $c_{i}>0$. Then, $a_{i}-b_{i}c_{i}\leq 1$ and
$|h_{i}(x)-h_{i}(y)|\leq\frac{a_{i}-b_{i}c_{i}}{c_{i}|x-y|+1}|x-y|\leq\frac{|x-y|}{c_{i}|x-y|+1}.$
Hence,
$\omega_{h_{i}}(s)\leq\frac{s}{c_{i}s+1},\ \ s\geq 0.$
By this and $c_{i}>0$,
$\phi_{h_{i}}(t)\leq\frac{t}{c_{i}t+1}<t,\ t>0.$
(ii) Assume $c_{i}<0$. Then, $a_{i}-b_{i}c_{i}\leq(c_{i}+1)^{2}$ and
$|h_{i}(x)-h_{i}(y)|\leq\frac{a_{i}-b_{i}c_{i}}{(c_{i}+1)(c_{i}(1-|x-y|)+1)}|x-y|\leq\frac{(1+c_{i})|x-y|}{-c_{i}|x-y|+1+c_{i}}.$
Hence,
$\omega_{h_{i}}(s)\leq\frac{(1+c_{i})s}{-c_{i}s+1+c_{i}},\ \ s\geq 0.$
By this and $c_{i}<0$,
$\phi_{h_{i}}(t)\leq\frac{(1+c_{i})t}{-c_{i}t+1+c_{i}}<t,\ t>0.$
(iii) Assume $c_{i}=0$. Then, $0<a_{i}\leq 1$ and $|h_{i}(x)-h_{i}(y)|=a_{i}|x-y|$. By the assumption for $a_{i},b_{i},c_{i}$, $h_{j}(1)>h_{j}(0)$ for every $j$. Hence, by $N\geq 2$, $a_{i}=h_{i}(1)-h_{i}(0)<1$. ∎
Remark 2.9.
The technique used above is related to a characterization of Rakotch contractions. See [20, Lemma 2.10 (iii)].
Let $(\mathcal{I}_{N})^{*}:=\cup_{n\in\mathbb{N}}(\mathcal{I}_{N})^{n}$. We denote the length of $\sigma\in(\mathcal{I}_{N})^{*}$ by $|\sigma|$. For an LF system $\{h_{i}\}_{i}$ and $\sigma=(\sigma_{1},\cdots,\sigma_{n})\in(\mathcal{I}_{N})^{*}$, we let
$h_{\sigma}(x)=h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
Lemma 2.10.
Let $\{h_{i}\}_{i}$ be an LF system. Then, there exists $c\in(0,1)$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
$h_{\sigma}(1)-h_{\sigma}(0)\geq c^{|\sigma|}.$
Proof.
Let $A_{i}:=\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}$, $i\in\mathcal{I}_{N}$. We assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i$. Let $(i_{n})_{n}\in(\mathcal{I}_{N})^{\mathbb{N}}$. Let
$\begin{pmatrix}p_{n}&q_{n}\\ r_{n}&s_{n}\end{pmatrix}=A_{i_{1}}\cdots A_{i_{n}},\ n\geq 1.$
Let $C_{1}:=\max\{|a_{i}|,|b_{i}|,|c_{i}|,|d_{i}|:i\in\mathcal{I}_{N}\}$. Then, for every $n$, $|r_{n+1}|+|s_{n+1}|\leq C_{1}(|r_{n}|+|s_{n}|)$. Hence, $|r_{n}|+|s_{n}|\leq 2C_{1}^{n}$ for every $n$.
Assume that $\sigma=(i_{1},\cdots,i_{n})$. Then, $|\sigma|=n$ and
$h_{\sigma}(1)-h_{\sigma}(0)=\frac{1}{s_{n}(r_{n}+s_{n})}\geq\frac{1}{(|r_{n}|+|s_{n}|)^{2}}\geq\frac{1}{4|C_{1}|^{2n}}.$
We give an upper bound for $h_{\sigma}(1)-h_{\sigma}(0)$.
Lemma 2.11.
Let $\{h_{i}\}_{i}$ be an LF system. Then, there exists $C>0$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
$h_{\sigma}(1)-h_{\sigma}(0)\leq\frac{C}{|\sigma|}.$
Proof.
We first remark that
$(c_{i}x+d_{i})(c_{i}y+d_{i})\geq 1+|c_{i}||x-y|,\ x,y\in[0,1].$ (2.5)
We assume that $x\leq y$. Assume $c_{i}>0$. Then,
$c_{i}y+d_{i}\geq c_{i}x+d_{i}\geq\min_{x\in[0,1]}c_{i}x+d_{i}=\min\{d_{i},c_{i}+d_{i}\}=d_{i}\geq 1.$
Hence,
$(c_{i}x+d_{i})(c_{i}y+d_{i})=(c_{i}x+d_{i})^{2}+c_{i}(y-x)(c_{i}x+d_{i})\geq 1+c_{i}(y-x).$
Assume $c_{i}<0$. Then,
$(c_{i}x+d_{i})(c_{i}y+d_{i})=(c_{i}y+d_{i})^{2}-c_{i}(y-x)(c_{i}x+d_{i})\geq 1-c_{i}(y-x).$
If $c_{i}=0$, then, $(c_{i}x+d_{i})(c_{i}y+d_{i})=d_{i}^{2}\geq 1=1+|c_{i}||x-y|$. Thus we see (2.5).
By (2.5),
$|h_{i}(x)-h_{i}(y)|\leq\frac{|x-y|}{1+|c_{i}||x-y|}.$
If $c_{i}=0$, then,
$|h_{i}(x)-h_{i}(y)|=\frac{a_{i}}{d_{i}}|x-y|.$
Since $h_{i}(1)-h_{i}(0)<1$, $a_{i}<d_{i}$. Let $C_{1}:=\min\{|c_{i}|:c_{i}\neq 0\}$ and $C_{2}:=\max\left\{\dfrac{a_{i}}{d_{i}}:c_{i}=0\right\}$. Then, $C_{1}>0$, $C_{2}<1$ and
$\max_{i}|h_{i}(x)-h_{i}(y)|\leq\max\left\{\frac{|x-y|}{1+C_{1}|x-y|},C_{2}|x-y|\right\}.$
We can take $C_{3}\in(0,C_{1})$ such that
$\max_{i}|h_{i}(x)-h_{i}(y)|\leq\frac{|x-y|}{1+C_{3}|x-y|},\ x,y\in[0,1].$
Let $a_{n}:=\max\left\{h_{\sigma}(1)-h_{\sigma}(0):\sigma\in(\mathcal{I}_{N})^{*},|\sigma|=n\right\}$. Then, $a_{1}\leq\dfrac{1}{1+C_{3}}$ and $a_{n+1}\leq\dfrac{a_{n}}{1+C_{3}a_{n}},n\geq 1$. Hence,
$a_{n}\leq\frac{1}{nC_{3}+1}\leq\frac{1}{nC_{3}}.$
Proposition 2.12 (Weak version of regularity).
Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are both LF systems and let $\varphi$ be the solution of (1.1). Then, there exists a constant $C>0$ such that for every $y\in(0,1)$ and $r>0$ satisfying that $0\leq y-r<y+r\leq 1$,
$\mu_{\varphi}\left([y-r,y+r]\right)\geq\exp(-C/r).$
This might be regarded as a weak version of the regularity of $\mu_{\varphi}$ in the sense of Ullman-Stahl-Totik. See [22, Criterion 1.1]. However, the following proof does not yield the regularity.
Proof.
We apply Lemma 2.10 to $\{g_{i}\}_{i}$ and Lemma 2.11 to $\{f_{i}\}_{i}$. Let $C$ be the constant appearing in Lemma 2.11. Take $y\in(0,1)$ and $r>0$ such that $0\leq y-r<y+r\leq 1$ arbitralily. Let $n$ be a natural number such that $\frac{C}{n}<r\leq\frac{C}{n-1}$. Then, by the compatibility condition and Lemma 2.11, there exist $i_{1},\cdots,i_{n}\in\mathcal{I}_{N}$ such that
$[f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)]\subset[y-r,y+r].$
Hence, by (2.3),
$\mu_{\varphi}([y-r,y+r])\geq g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0).$
Let $c$ be the constant appearing in Lemma 2.10. By Lemma 2.10,
$\mu_{\varphi}([y-r,y+r])\geq c^{n}\geq c^{1+C/r}.$
3. Singularity
In this section, we consider (1.1) under the setting that $\{g_{i}\}_{i}$ is an LF-system on $[0,1]$, and, $\{f_{i}\}_{i}$ is a weakly contractive D-system on $[0,1]$. By Lemma 2.8, both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are D-systems. Hence, by Lemma 2.3, there exists a unique continuous strictly increasing solution $\varphi$ of (1.1).
We consider singularity for the solution $\varphi$. We will give sufficient conditions for $\dim_{H}\mu_{\varphi}<1$. It is known that if $\dim_{H}\mu_{\varphi}<1$, then, $\varphi$ is a singular function on $[0,1]$, specifically, $\varphi$ is a strictly increasing and continuous function on $[0,1]$ such that its derivative is equal to zero almost everywhere.
The main purpose of this section is to extend [27, Theorem 43] to the case that $f_{i}$ can be non-linear. In [27], the author dealt with many objects other than the conjugate functional equation, so the arguments were not specialized for the functional equation. Some parts dealing with the functional equation in [27] are somewhat sketchy, so we give more details here. We introduce some new notations and the formulations of our main results and some of the arguments in the proofs are different from [27, Theorem 43]. These changes will make arguments more transparent.
We assume that an LF-system $\{g_{i}\}_{i}$ is defined by $g_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers satisfying the conditions (1), (2) and (3) in Definition 2.7. In this section, we assume $d_{i}=1$ for each $i$.
Let
$\alpha:=\min\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\middle|1\leq i\leq N-1\right\},$
where we delete [^2] This is different from [27]. $\frac{c_{0}}{1-a_{0}}$ if $a_{0}=1$. Let
$\beta:=\max\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\middle|1\leq i\leq N-1\right\},$
where $\dfrac{c_{0}}{1-a_{0}}:=+\infty$ if $a_{0}=1$.
Let $Y:=[\alpha,\beta]$. Let
$G_{i}(y):=\frac{(a_{i}-b_{i}c_{i})(y+1)}{(b_{i}y+1)((a_{i}+b_{i})y+(c_{i}+1))}=\frac{(b_{i+1}-b_{i})(y+1)}{(b_{i+1}y+1)(b_{i}y+1)},\ y\in[\alpha,\beta].$
Let
$H_{i}(y):=\frac{a_{i}y+c_{i}}{b_{i}y+1},\ y\in[\alpha,\beta].$
Let $G_{i}(+\infty)$ and $H_{i}(+\infty)$ be the limits. Specifically,
$G_{0}(+\infty):=1,G_{i}(+\infty):=0,H_{0}(+\infty):=+\infty,H_{i}(+\infty):=\frac{a_{i}}{b_{i}},1\leq i\leq N-1.$
The following is the same as [27, Lemma 27], so we will not repeat the proof.
Lemma 3.1.
(i) $0\in Y\subset[-1,+\infty]$.
(ii) Each $H_{i}$ is well-defined on $Y$.
(iii) $\alpha\leq H_{i}(\alpha)\leq H_{i}(\beta)\leq\beta$, that is, $H_{i}(Y)\subset Y$.
(iv) $\sum_{i\in\Sigma_{N}}G_{i}(y)=1,\ y\in Y.$
By (iii), we can compose $H_{i}$ freely on $Y$. We consider the compactification when $\beta=+\infty$. Then, $Y=[\alpha,\beta]$ is compact, and $G_{i}:Y\to[0,1]$ and $H_{i}:Y\to Y$ are continuous. We remark that $Y,G_{i},H_{i}$ are defined by $(g_{i})_{i}$, not by $(f_{i})_{i}$.
$\textup{Fix}(H_{i})$ denotes the set of fixed points of $H_{i}$ on $[\alpha,\beta]$. This is uniquely determined and
$\textup{Fix}(H_{0})=\left\{\dfrac{c_{0}}{1-a_{0}}\right\},\ \textup{Fix}(H_{i})=\left\{\dfrac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\right\},\ i\geq 1.$
Let
$\mathcal{P}_{N}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}>0,\sum_{i}p_{i}=1\right\}$
and
$\overline{\mathcal{P}_{N}}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}\geq 0,\sum_{i}p_{i}=1\right\}.$
Let
$\mathfrak{p}_{i}:=\left(G_{j}(\textup{Fix}(H_{i}))\right)_{j}\in\overline{\mathcal{P}_{N}},\ i\in\mathcal{I}_{N}.$
Let
$\dim_{H}\mu_{\varphi}:=\inf\left\{\dim_{H}A:A\in\mathcal{B}([0,1]),\mu_{\varphi}(A)=1\right\}.$
The following results extend [27, Theorem 43 (i)] to the case that $\{f_{i}\}_{i}$ may be non-affine.
Theorem 3.2.
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$, where we let $r_{i}:=\textup{Lip}(f_{i})$.
The probability vectors $\mathfrak{p}_{i},i\in\mathcal{I}_{N}$, are determined only by $(\alpha,\beta,\{G_{i}\}_{i},\{H_{i}\}i)$ and date back to $\{g_{i}\}_{i}$. $\{f_{i}\}_{i}$ is not related to this. This contains the affine case that $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i$. The reason why we have introduced $Y,G_{i}$ and $H_{i}$ will be clarified in the proof.
We also deal with the case that |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$. Let the total distance between probability vectors be
$\|{\bf p}-{\bf q}\|:=\sum_{i\in\mathcal{I}_{N}}|p_{i}-q_{i}|,\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
Theorem 3.3.
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$, and $\|{\bf r}-\mathfrak{p}_{0}\|\geq\epsilon$, where we let $r_{i}:=\textup{Lip}(f_{i})$ and ${\bf r}:=(r_{0},\cdots,r_{N-1})$.
If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$. If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i$, and $\mathfrak{p}_{0}={\bf q}$, then, the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$. See [27, Theorem 43 (ii)].
3.1. Proofs
We first give notations for symbolic dynamics. Let :=$\mathbb{N}:=\{1,2,\cdots\}$. We define a map $\pi:(\mathcal{I}_{N})^{\mathbb{N}}\to[0,1]$ by
$\pi(i_{1}i_{2}\cdots):=\lim_{k\to\infty}f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0).$
This is well-defined because each $f_{i}$ is a contraction and $\left(f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0)\right)_{k}$ is a Cauchy sequence.
Let $\xi_{j}:(\mathcal{I}_{N})^{\mathbb{N}}\to\mathcal{I}_{N}$ be the projection to the $j$-th coordinate. Let
[…]
Then, by Lemma 2.1,
$\pi\left(I(i_{1},\cdots,i_{k})\right)=\left[f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0),f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(1)\right].$
Hence, by (2.3),
$\mu_{\varphi}\left(\pi(I(i_{1},\cdots,i_{k}))\right)$ $=\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(1)\right)-\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(0)\right)$
$=g_{i_{1}}\circ\cdots\circ g_{i_{k}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{k}}(0).$
For $y\in Y$, let $\nu_{y}$ be a probability measure on $(\mathcal{I}_{N})^{\mathbb{N}}$ such that $\nu_{y}(I(i))=G_{i}(y)$, $i\in\mathcal{I}_{N}$, and,
$\frac{\nu_{y}(I(i_{1},\cdots,i_{n-1},i_{n}))}{\nu_{y}(I(i_{1},\cdots,i_{n-1}))}=G_{i_{n}}\circ H_{i_{n-1}}\circ\cdots\circ H_{i_{1}}(y),\ n\geq 2.$ (3.1)
Since $\sum_{i\in\mathcal{I}_{N}}G_{i}(y)=1$ for each $y\in Y$, the consistency condition holds. Therefore, a probability measure on $(\mathcal{I}_{N})^{\mathbb{N}}$ satisfying (3.1) uniquely exists by the Kolmogorov extension theorem.
For $i\in\mathcal{I}_{N}$, let $F_{i}:(\mathcal{I}_{N})^{\mathbb{N}}\to(\mathcal{I}_{N})^{\mathbb{N}}$ be a map defined by $F_{i}(z)=iz$. Now we see that
$\nu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\nu_{H_{i}(y)}\circ F_{i}^{-1}.$ (3.2)
Let $\mu_{y}:=\nu_{y}\circ\pi^{-1}$. Then, $\mu_{y}$ is a Borel probability measure on $[0,1]$. By $\pi\circ F_{i}=f_{i}\circ\pi$ and (3.2),
$\mu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\mu_{H_{i}(y)}\circ f_{i}^{-1}.$ (3.3)
We recall that $0\in Y=[\alpha,\beta]$. Then, by (3.3),
$\mu_{0}=\sum_{i\in\mathcal{I}_{N}}G_{i}(0)\mu_{H_{i}(0)}\circ f_{i}^{-1}.$
Since $H_{i}(0)\neq 0$ can occur, this does not define an iterated function system. (3.3) is in the framework of a place-dependent iterated function system [9].
By [27, Lemma 30],
$\nu_{0}\circ\pi^{-1}=\mu_{0}=\mu_{\varphi}.$ (3.4)
Let $\epsilon\in(0,1/2)$. We assume that $\epsilon\leq\min_{i}r_{i}\leq\max_{i}r_{i}\leq 1-\epsilon$. Since $\{f_{i}\}_{i}$ is a compatible system, $\sum_{i}r_{i}\geq 1$ and $r_{i}>0$ for each $i$. Let $s\geq 1$ is a constant such that $\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$.
As the following assertion shows, the constant $s$ is close to $1$ if $\delta>0$ is sufficiently small.
Lemma 3.4.
Let $\epsilon_{2}>0$. Then, for $\delta>0$ such that
$\frac{\log N}{\log(N/(1+\delta))}<1+\frac{\epsilon_{2}}{\log 1/\epsilon},$ (3.5)
it holds that $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$ for every choice of $(r_{i})_{i}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$.
Proof.
Let $u_{i}:=r_{i}/\sum_{i}r_{i}$. Then, $\sum_{i}u_{i}=1$. We see that $\sum_{i}r_{i}^{s}=1$ if and only if $\sum_{i}u_{i}^{s}=(\sum_{i}r_{i})^{-s}$. Hence, if $\sum_{i}r_{i}\leq 1+\delta$, then,
$\frac{\log(\sum_{i}u_{i}^{s})}{s}\leq-\log(1+\delta).$
$\sum_{i}u_{i}^{s}$ takes its minimum at $u_{0}=\cdots=u_{N-1}=\dfrac{1}{N}$. We see this by the method of Lagrange multipliers. The minimum is $N^{1-s}\leq 1$. Hence, $-\dfrac{s-1}{s}\log N\geq-\log(\sum_{i}r_{i})$, which is equivalent to $s<\dfrac{\log N}{\log(N/(1+\delta))}$. Therefore, if (3.5) holds, then, $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$. ∎
Let
$R_{n}({\bf q};z):=\frac{\nu_{0}(I_{n}(z))}{q_{\xi_{1}(z)}\cdots q_{\xi_{n}(z)}},\ \ {\bf q}=(q_{i})_{i}\in\mathcal{P}_{N},z\in(\mathcal{I}_{N})^{\mathbb{N}}.$
By a measure-theoretic argument, we can show that
Lemma 3.5.
Let $\widetilde{q}_{i}:=r_{i}^{s}>0$ and ${\bf\widetilde{q}}:=(\widetilde{q}_{i})_{i}\in\mathcal{P}_{N}$. If there exists $\epsilon_{2}>0$ such that
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf\widetilde{q}};z)}{n}\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$ (3.6)
then,
$\dim_{H}\mu_{\varphi}\leq s-\frac{\epsilon_{2}}{\log 1/\epsilon}.$
We omit the proof of this assertion. See the proof of [27, Lemma 7]. (3.4) plays an important role.
By Lemmas 3.4 and 3.5, we obtain that
Proposition 3.6.
Assume that there exists $\epsilon_{2}>0$ such that for every ${\bf q}\in\mathcal{P}_{N}$,
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$ (3.7)
Then, for $\delta=\delta(N,\epsilon,\epsilon_{2})>0$ satisfying (3.5), $\dim_{H}\mu_{\varphi}<1$.
Now it suffices to establish (3.7). We give some definitions for the relative entropy. Let
$s_{N}\left({\bf p}|{\bf q}\right):=\sum_{i\in\mathcal{I}_{N}}p_{i}\log\frac{q_{i}}{p_{i}},\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
If $p_{i}=0$, then, $p_{i}\log\frac{q_{i}}{p_{i}}:=0$. $s_{N}$ is continuous on $\overline{\mathcal{P}_{N}}$. Then, by Jensen’s inequality, $s_{N}({\bf p}|{\bf q})\leq 0$. It holds that $s_{N}\left({\bf p}|{\bf q}\right)=0$ if and only if ${\bf p}={\bf q}$.
For $y\in Y,z\in(\mathcal{I}_{N})^{\mathbb{N}}$, let
$p_{j}(y;z):=\left(G_{i}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(y)\right)_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}},\ j\geq 1.$
and
$p_{0}(y;z):=(G_{i}(y))_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}}.$
By a martingale argument, we can show that
Proposition 3.7.
It holds that for every ${\bf q}\in\mathcal{P}_{N}$,
$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}=\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}(p_{j}(0;z)|{\bf q}),\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
We omit the proof of this assertion. See the proof of [27, Proposition 6]. It is important to assume that ${\bf q}\in\mathcal{P}_{N}$, not ${\bf q}\in\overline{\mathcal{P}_{N}}$.
For $\epsilon\in(0,1)$, let $\overline{\mathcal{P}_{N}(\epsilon)}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}\geq\epsilon,\sum_{i}p_{i}=1\right\}$. Then, $\overline{\mathcal{P}_{N}(\epsilon)}\subset\mathcal{P}_{N}\subset\overline{\mathcal{P}_{N}}$.
Recall Lemma 3.4. If
$\frac{\log N}{\log(N/(1+\delta))}\leq 1+\frac{\log 2}{\log(1/\epsilon)},$
then, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. In other words, ${\bf\widetilde{q}}$ is apart from the boundary $\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$
Assume that
Proposition 3.8.
If |$|\{\mathfrak{p}_{i}\}_{i}|\geq 2$, then, there exists $\epsilon_{2}>0$ such that for every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
Theorem 3.2 follows from Propositions 3.6, 3.7 and 3.8. We now show Proposition 3.8. We give a strategy of the proof. By the definition, $s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq 0$. Hence, it suffices to show the number of $j$ satisfying that $s_{N}\left(p_{j}(0;z)|{\bf q}\right)$ is strictly negative has positive density in $\mathbb{N}$.
We first give some further definitions for quantities of the relative entropy.
Lemma 3.9.
Let
$V\left({\bf q};\epsilon\right):=\max\left\{s_{N}\left({\bf p}|{\bf q}\right)\middle|{\bf p}\in\overline{\mathcal{P}_{N}},\|{\bf p}-{\bf q}\|\geq\epsilon\right\},\ \epsilon>0.$
Then,
(i) $V\left({\bf q};\epsilon\right)<0$.
(ii) $V\left({\bf q};\epsilon\right)$ is decreasing with respect to $\epsilon$.
(iii) $V\left({\bf q};\epsilon\right)$ is right-continuous with respect to $\epsilon$.
(iv) $V\left({\bf q};\epsilon\right)$ is continuous with respect to ${\bf q}$.
Proof.
(i) and (ii) are easy to see. (iii) Assume that for some $\epsilon_{0}\in(0,1/3)$, $V\left({\bf q};\epsilon\right)$ is not right-continuous. Then, $V\left({\bf q};\epsilon_{0}\right)>\lim_{\epsilon\to\epsilon_{0}+0}V\left({\bf q};\epsilon\right)$. Assume that $V\left({\bf q};\epsilon_{0}\right)=s_{N}\left({\bf p}|{\bf q}\right)$. If $\|{\bf p}-{\bf q}\|>\epsilon_{0}$, then, this cannot occur. Assume $\|{\bf p}-{\bf q}\|=\epsilon_{0}$. Then, there exists a sequence $\{{\bf p}^{(n)}\}_{n}$ in $\overline{\mathcal{P}_{N}}$ such that $\|{\bf p}^{(n)}-{\bf q}\|\geq\epsilon_{0}+N^{-n}$ for each $n$ and $\lim_{n\to\infty}\|{\bf p}^{(n)}-{\bf q}\|=\epsilon_{0}$. Hence,
$\liminf_{n\to\infty}V\left({\bf q};\epsilon_{0}+N^{-n}\right)\geq\lim_{n\to\infty}s_{N}\left({\bf p}^{(n)}|{\bf q}\right)=s_{N}\left({\bf p}|{\bf q}\right)=V\left({\bf q};\epsilon_{0}\right).$
This is a contradiction.
(iv) follows from (iii) and
$V\left({\bf q};\epsilon+\|{\bf p}-{\bf q}\|\right)\leq V\left({\bf p};\epsilon\right)$
By Lemma 3.9, we can let
$V(\epsilon):=\max_{{\bf q}\in\overline{\mathcal{P}_{N}}}V\left({\bf q};\epsilon\right)<0.$
We now give a quantity defined by $Y,\{G_{i}\}_{i}$ and $\{H_{i}\}_{i}$.
Lemma 3.10.
Let
$W_{i}\left({\bf q},y\right):=\max\left\{\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|,\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|\right\},\ {\bf q}\in\overline{\mathcal{P}_{N}},y\in Y,i\in\mathcal{I}_{N}.$
Then,
(i) $W_{i}({\bf q},y)$ is continuous on $\overline{\mathcal{P}_{N}}\times Y$.
(ii) $W_{i}({\bf q},y)=0$ if and only if ${\bf q}=\mathfrak{p}_{i}$ and $y=\textup{Fix}(H_{i})$.
This assertion is easily shown. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then, $s_{N}\left((G_{k}(y))_{k}|{\bf q}\right)<0$. However, in our setting, we cannot exclude the possibility that $|G_{k}(y)-q_{k}|=0$ for every $k\in\mathcal{I}_{N}$. Assume that |$\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$. If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|$ is close to 00, then, $\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|$ is strictly positive. Hence, $W_{i}\left({\bf q},y\right)>0$.
The following corresponds to [27, Lemma 8]. Informally speaking, if $W_{i}({\bf q},y)$ is positive uniformly to $y$, then, a sum of two successive terms of $s_{N}$ is negative.
Lemma 3.11.
Let ${\bf q}\in\mathcal{P}_{N}$. Let $i_{1}\in\mathcal{I}_{N}$. Assume that for some $\epsilon_{1}>0$,
$\inf_{y\in Y}W_{i_{1}}\left({\bf q},y\right)\geq\epsilon_{1}.$ (3.8)
Assume that $X_{j+1}(z)=i_{1}$. Then,
$s_{N}\left(p_{j}(0;z)|{\bf q}\right)+s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)\leq V(\epsilon_{1})<0.$
Proof.
The proof is essentially the same as the proof of [27, Lemma 8] [^3] It contains a typo.. Since $X_{j+1}(z)=i_{1}$,
$p_{j}(0;z)=\left(G_{k}(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))\right)_{k\in\mathcal{I}_{N}},$
and
$p_{j+1}(0;z)=\left(G_{k}(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))\right)_{k\in\mathcal{I}_{N}}.$
Let ${\bf q}=(q_{k})_{k}$. Then, by (3.8), either
$\left|G_{k_{1}}(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))-q_{k_{1}}\right|\geq\epsilon_{1},\textup{ for some }k_{1}\in\mathcal{I}_{N},$
or,
$\left|G_{k_{2}}(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))-q_{k_{2}}\right|\geq\epsilon_{1},\textup{ for some }k_{2}\in\mathcal{I}_{N},$
holds. If the former holds, then, $s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$. If the latter holds, then, $s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$. Now it suffices to recall that $s_{N}$ takes non-positive values. ∎
Let :=|$\ell:=|\{\mathfrak{p}_{i}\}_{i}|\geq 2$. We assume that $\{\mathfrak{p}_{i}\}_{i}=\{\mathfrak{p}_{j_{k}}:1\leq k\leq\ell\}$ where $10\leq j_{1}<\cdots<j_{\ell}\leq N-$ 1. Let $\delta_{0}:=\min\{\|\mathfrak{p}_{j_{k_{1}}}-\mathfrak{p}_{j_{k_{2}}}\|:k_{1}\neq k_{2}\}>0$. Let $B_{j}:=\left\{{\bf p}\in\overline{\mathcal{P}_{N}}:\|{\bf p}-\mathfrak{p}_{j}\|\leq\delta_{0}/4\right\}$. $B_{j}=B_{k}$ can occur even if $j\neq k$. Define a coloring map $\mathscr{A}:\overline{\mathcal{P}_{N}}\to\mathcal{I}_{N}$ by
$\mathscr{A}\left({\bf p}\right):=\begin{cases}j_{2}&{\bf p}\in B_{j_{1}}\\ \vdots&\\ j_{\ell}&{\bf p}\in B_{j_{\ell-1}}\\ j_{1}&{\bf p}\in B_{j_{\ell}}\\ 0&\textup{ otherwise}\end{cases}.$
Under these definitions, we see that
Lemma 3.12.
$\widetilde{\epsilon_{1}}:=\inf\left\{W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\middle|{\bf q}\in\overline{\mathcal{P}_{N}},y\in Y\right\}>0.$
Proof.
Fix ${\bf q}\in\overline{\mathcal{P}_{N}}$. By Lemma 3.10,
$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{2}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|.$
If $(G_{j}(y))_{j}\notin B_{\mathscr{A}({\bf q})}$, then, $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}=\left(G_{j}(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)}))\right)_{j}$ and hence, $y$ is not contained in an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$. Since $G_{j}$ and $H_{\mathscr{A}\left({\bf q}\right)}$ are continuous, $\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$ is continuous with respect to $y$ on $Y$. Hence,
$\inf_{(G_{j}(y))_{j}\notin B_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
$\geq\inf_{y\notin J_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|=:\epsilon_{\mathscr{A}\left({\bf q}\right)}>0.$
Assume that $(G_{j}(y))_{j}\in B_{\mathscr{A}({\bf q})}$. Since ${\bf q}\notin B_{\mathscr{A}({\bf q})}$,
$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{N}\sum_{j}|G_{j}(y)-q_{j}|\geq\frac{\delta_{0}}{4N}.$
Hence,
$\inf_{y\in Y}W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\min\left\{\frac{\delta_{0}}{4N},\epsilon_{j_{k}}\middle|1\leq k\leq\ell\right\}=:\tilde{\epsilon_{1}}.$
We discuss the frequency of occurrence of some elements in $\mathcal{I}_{N}$ in $(\xi_{n})_{n}$.
Lemma 3.13.
Let $I\subset\mathcal{I}_{N}$ be a non-empty set. Let $c_{I}:=\inf_{y\in Y}\sum_{i\in I}G_{i}(y)$. Then,
$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n:\xi_{j}(z)\in I\}|}{n}\geq c_{I},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
This corresponds to [27, Lemma 9], but the formulation differs.
Proof.
If $c_{I}=0$, then, the assertion is obvious. Assume $c_{I}>0$. We denote the expectation with respect to $\nu_{0}$ by $E^{\nu_{0}}$. Let $\mathcal{F}_{m}:=\sigma(\xi_{1},\cdots,\xi_{m})$ for $m\geq 1$ and $\mathcal{F}_{0}$ be the trivial $\sigma$-algebra on $(\mathcal{I}_{N})^{\mathbb{N}}$. We denote the indicator function of $A\subset(\mathcal{I}_{N})^{\mathbb{N}}$ by ${\bf 1}_{A}$.
Let $M_{0}:=0$ and
$M_{n}-M_{n-1}:={\bf 1}_{\{\xi_{n}\in I\}}-E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{n}\in I\}}|\mathcal{F}_{n-1}\right],n\geq 1.$
Then, $\{M_{n},\mathcal{F}_{n}\}_{n}$ is a martingale.
By (3.1), $E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{n}\in I\}}|\mathcal{F}_{n-1}\right]\geq c_{I}$, $\nu_{0}$-a.s. Therefore, for every $\eta\in(0,1)$,
$\nu_{0}\left(\frac{|\{1\leq j\leq n:\xi_{j}\in I\}|}{n}\leq(1-\eta)c_{I}\right)\leq\nu_{0}\left(M_{n}\leq-n\eta c_{I}\right).$
We remark that $|M_{n}-M_{n-1}|\leq 1$ for each $n\geq 1$. By Azuma’s inequality ([1], [31, E14.2]),
$\nu_{0}\left(M_{n}\leq-n\eta c_{I}\right)\leq\exp\left(-\frac{n\eta^{2}c_{I}^{2}}{2}\right).$
Using $c_{I}>0$ and Borel-Cantelli’s lemma,
$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n:\xi_{j}(z)\in I\}|}{n}\geq(1-\eta)c_{I},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
The assertion follows if $\eta\to+0$. ∎
We proceed to the proof of Proposition 3.8. The proof is divided into two cases. The following is easily shown.
Lemma 3.14.
(i) $a_{0}\leq 1$ and $b_{N-1}+c_{N-1}\geq 0$.
(ii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then, $\alpha=-1,\beta=+\infty$.
(iii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$, then, $\alpha>-1,\beta=+\infty$.
(iv) If $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$, then, $\alpha=-1,\beta<+\infty$.
(v) If $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$, then, $\alpha>-1,\beta<+\infty$.
(vi) If $a_{0}=1$, then, $\textup{Fix}(H_{0})=\{+\infty\}$ and $\mathfrak{p}_{0}=(1,0,\cdots,0)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
(vii) If $b_{N-1}+c_{N-1}=0$, then, $\textup{Fix}(H_{N-1})=\{-1\}$ and $\mathfrak{p}_{N-1}=(0,\cdots,0,1)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
$\alpha=-1$ or $\beta=+\infty$ occurs according to the values of $a_{0}$ and $b_{N-1}+c_{N-1}$.
Case 1. We first deal with the case that $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$.
In this case, we do not need to assume ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
Proof of Proposition 3.8, Case 1.
By the assumption that $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$ and Lemma 3.14 (v), $-1<\alpha\leq\beta<+\infty$. Hence,
$10<\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)\leq\sup_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)<$ 1 .
Fix ${\bf q}\in\overline{\mathcal{P}_{N}}$. Let $\widetilde{c}:=\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)$. Then, $c_{\{\mathscr{A}({\bf q})\}}\geq\widetilde{c}>0$ and by Lemma 3.13 for $I=\{\mathscr{A}({\bf q})\}$,
$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}{\bf 1}_{\{\xi_{k}=\mathscr{A}({\bf q})\}}(z)\geq\widetilde{c},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
The number $i_{1}\in\mathcal{I}_{N}$ in Lemma 3.11 will be chosen according to the position of ${\bf q}$. By applying Lemma 3.11 to $i_{1}=\mathscr{A}({\bf q})$ and $\epsilon_{1}=\widetilde{\epsilon_{1}}$ where $\widetilde{\epsilon_{1}}$ is the constant in Lemma 3.12, we see that if $X_{i+1}(z)=\mathscr{A}({\bf q})$, then,
$s_{N}\left(p_{i}(0;z)|{\bf q}\right)+s_{N}\left(p_{i+1}(0;z)|{\bf q}\right)\leq V(\widetilde{\epsilon_{1}})<0.$
Now the assertion follows if we let $\epsilon_{2}:=\dfrac{\widetilde{c}\ V(\widetilde{\epsilon_{1}})}{2}$. ∎
Case 2. We now deal with the case that $a_{0}=1$ or $b_{N-1}+c_{N-1}=0$.
If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then, $c_{\{0\}}=c_{\{N-1\}}=0$, so we consider $c_{\{0,N-1\}}$. In the following, we do not use Lemma 3.12 or the assumption that |$|\{\mathfrak{p}_{i}\}_{i}|\geq 2$. On the other hand, we assume ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
Proof of Proposition 3.8, Case 2.
Case 2.1. We first consider the case that $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$.
By Lemma 3.14, $\alpha>-1$ and $\beta=+\infty$, $H_{0}(y)=y+c_{0}$, and $\textup{Fix}(H_{0})=\{+\infty\}$. $\mathfrak{p}_{0}=(1,0,\cdots,0)\notin\mathcal{P}_{N}$. On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\geq 1$.
Since $G_{0}(+\infty)=G_{0}(H_{0}(+\infty))=1$ and $q_{0}<1-\epsilon/2$,
$\inf_{y\in Y}W_{0}({\bf q},y)>c_{0}(\epsilon/2)>0,$
where the lower bound $c_{0}(\epsilon/2)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
Since $\alpha>-1$, $\inf_{y\in Y}G_{0}(y)\geq c>0$. By using Lemma 3.11 and Lemma 3.13 for $I=\{0\}$, we have the assertion.
Case 2.2. We secondly consider the case that $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$.
We recall that $\alpha=-1$ and $\beta<+\infty$, By Lemma 3.14, $H_{N-1}(y)=\dfrac{(1-2b_{N-1})y+b_{N-1}}{1-b_{N-1}}$, and $\textup{Fix}(H_{N-1})=\{-1\}$. $\mathfrak{p}_{N-1}=(0,\cdots,0,1)\notin\mathcal{P}_{N}$. On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\leq N-2$.
Since $G_{N-1}(-1)=G_{N-1}(H_{N-1}(-1))=1$ and $q_{N-1}<1-\epsilon/2$,
$\inf_{y\in Y}W_{N-1}({\bf q},y)>c_{N-1}(\epsilon/2)>0,$
where the lower bound $c_{N-1}(\epsilon)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. Since $\beta<+\infty$, $\inf_{y\in Y}G_{N-1}(y)\geq c>0$. The rest of the argument is the same as in the above case.
Case 2.3. We finally consider the case that $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$.
Since $\textup{Fix}(H_{0})=\{+\infty\}$, $\textup{Fix}(H_{N-1})=\{-1\}$, $q_{0}<1-\epsilon/2$ and $q_{N-1}<1-\epsilon/2$, it holds that
$\inf_{y\in Y}W_{0}({\bf q},y)>c_{0}(\epsilon/2)>0$
and
$\inf_{y\in Y}W_{N-1}({\bf q},y)>c_{N-1}(\epsilon/2)>0.$
Since
$G_{0}(y)+G_{N-1}(y)=\frac{b_{1}b_{N-1}y^{2}+2b_{1}y1+b_{1}-b_{N-1}}{(b_{1}y+1)(b_{N-1}y+1)},$
it holds that
$\inf_{y\in Y}G_{0}(y)+G_{N-1}(y)>0.$
By using Lemma 3.11 and Lemma 3.13 for $I=\{0,N-1\}$, we have the assertion. ∎
Now the proof of Theorem 3.2 completes.
Remark 3.15.
(i) In the proof of [27, Theorem 43], the details of Case 2 were not given.
(ii) Assume that $\{g_{i}\}_{i}$ are affine functions such that $g_{i}(x)=b_{i}x+(b_{0}+\cdots+b_{i-1})$. Since $Y=\{0\}$, there is no need to introduce $H_{i}$, and $W_{i}({\bf q},y)=\max_{k}|G_{k}(y)-q_{k}|=\max_{k}|b_{k}-q_{k}|$. However, $\mathscr{p}_{0}=\cdots=\mathscr{p}_{N-1}=(b_{i})_{i}$, and hence this is not in the scope of Theorem 3.2. This case is considered in Theorem 3.3.
We now proceed to the proof of Theorem 3.3.
By (3.6), if we take $\delta>0$ such that
$\frac{\log N}{\log(N/(1+\delta))}\leq 1+\frac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)},$ (3.9)
then, $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$.
Assume that $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$. Since $\min_{i}r_{i}\geq\epsilon$, $\sum_{i}|r_{i}-r_{i}^{s}|\leq\epsilon^{1-s}-1\leq\epsilon/2$. Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$, where we let $\widetilde{q}_{i}:=r_{i}^{s}$ for each $i$ and ${\bf\widetilde{q}}:=(\widetilde{q}_{i})_{i}$. Since $\epsilon\leq 1/2$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\epsilon/2$ for each $i$. Hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
By Lemma 3.10 (i) and (ii), it holds that
$\epsilon_{3}:=\inf\left\{W_{i}({\bf q},y)\middle|i\in\mathcal{I}_{N},y\in Y,{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2\right\}>0.$
By Lemma 3.11, we see that for every $j\geq 0$, every $z\in(\mathcal{I}_{N})^{\mathbb{N}}$, and every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$ with $\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$,
$s_{N}\left(p_{j}(0;z)|{\bf q}\right)+s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{3}\right).$
By Lemma 3.9, we can let
$V_{\epsilon}(\epsilon_{3}):=\max\left\{V\left({\bf q};\epsilon_{3}\right)\middle|{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2\right\}<0.$
Then,
$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\frac{V_{\epsilon}(\epsilon_{3})}{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
Assume that $\delta>0$ satisfies not only (3.9) but also (3.5) for $\epsilon_{2}:=V_{\epsilon}(\epsilon_{3})/2$. Then, by Propositions 3.6 and 3.7, $\dim_{H}\mu_{\varphi}<1$.
Remark 3.16.
(i) We do not need to use Lemma 3.13. For every small $\epsilon>0$ and every $i\in\mathcal{I}_{N}$,
$\min\left\{W_{i}({\bf q},y)\middle|y\in Y,{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}\right\}=0,$
by Lemma 3.10 (ii).
(ii) The case that both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are affine is easily handled by the dimension formula. However, if some of the functions $\{f_{i},g_{i}\}_{i}$ are non-affine, then, it is difficult to judge the singularity by a generalized dimension formula in [9].
4. Example
We give an example that $\{g_{i}\}_{i}$ is an LF system and $\{f_{i}\}_{i}$ is non-affine and the solution $\varphi$ is singular. The result depends on the quantitative estimates in Section 3.
Let $N=2$. Then,
$G_{0}(y)=\frac{b_{1}(y+1)}{b_{1}y+1},\ H_{0}(y)=a_{0}y+c_{0},\ H_{1}(y)=\frac{(1-b_{1}+c_{1})y+c_{1}}{b_{1}y+1},$
Let
$g_{0}(x):=\frac{x}{-x+6},\ g_{1}(x):=\frac{3x+1}{5-x}.$
Then, $g_{0}$ is contractive, and $g_{1}$ is weakly contractive but not contractive. It holds that $a_{0}=1/6$, $c_{0}=-1/6$, $a_{1}=3/5$, $b_{1}=1/5$, and $c_{1}=-1/5$. Hence,
$\alpha=\min\{0,-1/5,-1\}=-1,\beta=\max\{0,-1/5,-1\}=0.$
Hence,
$\sup_{y\in[-1,0]}G_{0}(y)=1-\inf_{y\in[-1,0]}G_{1}(y)=\frac{1}{5}.$
Let $\mathcal{P}[4/5]:=\left\{(p_{0},p_{1})\in\mathcal{P}_{2}:p_{0}\geq 4/5\right\}$. Let ${\bf q}=(q_{0},q_{1})\in\mathcal{P}[4/5]$. Then,
$|G_{0}(y)-q_{0}|=|G_{1}(y)-q_{1}|\geq\frac{3}{5},\ y\in[-1,0].$
Hence,
$s_{2}\left(p_{j}(0,z)|{\bf q}\right)\leq V\left({\bf q};\frac{6}{5}\right),\ j\geq 0,z\in(\mathcal{I}_{N})^{\mathbb{N}}.$ (4.1)
Since $s_{2}({\bf p}|{\bf q})=s_{2}\left((p_{0},1-p_{0})|(q_{0},1-q_{0})\right)$ is increasing with respect to $p_{0}$ on $[0,q_{0}]$,
$V\left({\bf q};\frac{6}{5}\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|(q_{0},1-q_{0})\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|\left(\frac{4}{5},\frac{1}{5}\right)\right)\leq-\frac{6}{5}\log 2\leq-0.83.$
By Proposition 3.6, it holds that
$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.83,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\mathcal{P}[4/5]$.}$ (4.2)
Let $\epsilon\in(0,1/2)$. Let $\{f_{0},f_{1}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1$. Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}=1$. Assume that $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$. Then, by (4.2) and Lemma 3.5,
$\dim_{H}\mu_{\varphi}\leq s-\frac{0.83}{\log(1/\epsilon)}.$
We see that $\dfrac{0.83}{\log 10}\geq 0.36$ and $\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$. Hence, (3.5) holds for $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$. By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, $r_{0}+r_{1}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$, then,
$\dim_{H}\mu_{\varphi}\leq s-\frac{0.83}{\log(1/\epsilon)}\leq\dfrac{\log 2}{\log 2-\log 1.2}-\frac{0.83}{\log 10}<1.$
Now we can let $\{f_{0},f_{1}\}$ be a compatible system such that $0.85\leq r_{0}\leq 0.9$ and $0.1\leq r_{1}\leq 0.3$. Then, $r_{0}+r_{1}\leq 1.2$. By the compatibility condition, $r_{0}+r_{1}\geq 1$. The maps $f_{0}$ and $f_{1}$ may not be affine. Since $s\leq\frac{\log 2}{\log 2-\log 1.2}\leq 1.357$, $r_{0}^{s}\geq(0.85)^{1.357}\geq 0.8$. Hence, $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$.
Remark 4.1.
We see that $a_{0}<1$ and $b_{1}+c_{1}=0$. Hence, this case corresponds to Case 2.2. However, we do not need to apply Lemma 3.11 since we have the upper bound for the relative entropy in (4.1).

### v3

@@
 In Section 3, we deal with the singularity.
-Finally, we give an example in Section 4.
+Finally, we give three numerical examples in Section 4.
 2. Preliminaries
@@
 For an LF system $\{h_{i}\}_{i}$ and $\sigma=(\sigma_{1},\cdots,\sigma_{n})\in(\mathcal{I}_{N})^{*}$, we let
-$h_{\sigma}(x)=h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
+$h_{\sigma}(x):=h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
 Lemma 2.10.
@@
 Let $A_{i}:=\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}$, $i\in\mathcal{I}_{N}$.
-We assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i$.
+We assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i\in\mathcal{I}_{N}$.
 Let $(i_{n})_{n}\in(\mathcal{I}_{N})^{\mathbb{N}}$.
@@
 In [27], the author dealt with many objects other than the conjugate functional equation, so the arguments were not specialized for the functional equation.
-Some parts dealing with the functional equation in [27] are somewhat sketchy, so we give more details here.
 We introduce some new notations and the formulations of our main results and some of the arguments in the proofs are different from [27, Theorem 43].
@@
 We assume that an LF-system $\{g_{i}\}_{i}$ is defined by $g_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers satisfying the conditions (1), (2) and (3) in Definition 2.7.
-In this section, we assume $d_{i}=1$ for each $i$.
+In this section, we assume $d_{i}=1$ for each $i\in\mathcal{I}_{N}$.
 Let
@@
 Let
-$\mathfrak{p}_{i}:=\left(G_{j}(\textup{Fix}(H_{i}))\right)_{j}\in\overline{\mathcal{P}_{N}},\ i\in\mathcal{I}_{N}.$
+${\bf G}(y):=(G_{0}(y),\dots,G_{N-1}(y))\in\overline{\mathcal{P}_{N}},\ y\in Y.$
+Let
+$\mathfrak{p}_{i}:={\bf G}\left(\textup{Fix}(H_{i})\right)\in\overline{\mathcal{P}_{N}},\ i\in\mathcal{I}_{N}.$
 Let
@@
 Theorem 3.2.
-If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$, where we let $r_{i}:=\textup{Lip}(f_{i})$.
+If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, where we let $r_{i}:=\textup{Lip}(f_{i})$.
 The probability vectors $\mathfrak{p}_{i},i\in\mathcal{I}_{N}$, are determined only by $(\alpha,\beta,\{G_{i}\}_{i},\{H_{i}\}i)$ and date back to $\{g_{i}\}_{i}$.
 $\{f_{i}\}_{i}$ is not related to this.
-This contains the affine case that $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i$.
+This contains the affine case that $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$.
 The reason why we have introduced $Y,G_{i}$ and $H_{i}$ will be clarified in the proof.
@@
 Theorem 3.3.
-If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$, and $\|{\bf r}-\mathfrak{p}_{0}\|\geq\epsilon$, where we let $r_{i}:=\textup{Lip}(f_{i})$ and ${\bf r}:=(r_{0},\cdots,r_{N-1})$.
+If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, and $\|{\bf r}-\mathfrak{p}_{0}\|\geq\epsilon$, where we let $r_{i}:=\textup{Lip}(f_{i})$ and ${\bf r}:=(r_{0},\cdots,r_{N-1})$.
 If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$.
-If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i$, and $\mathfrak{p}_{0}={\bf q}$, then, the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$.
+If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$, and $\mathfrak{p}_{0}={\bf q}$, then, the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$.
 See [27, Theorem 43 (ii)].
@@
 Let $\epsilon\in(0,1/2)$.
-We assume that $\epsilon\leq\min_{i}r_{i}\leq\max_{i}r_{i}\leq 1-\epsilon$.
-Since $\{f_{i}\}_{i}$ is a compatible system, $\sum_{i}r_{i}\geq 1$ and $r_{i}>0$ for each $i$.
+We assume that $\epsilon\leq\min_{i\in\mathcal{I}_{N}}r_{i}\leq\max_{i\in\mathcal{I}_{N}}r_{i}\leq 1-\epsilon$.
+Since $\{f_{i}\}_{i}$ is a compatible system, $\sum_{i\in\mathcal{I}_{N}}r_{i}\geq 1$ and $r_{i}>0$ for each $i\in\mathcal{I}_{N}$.
 Let $s\geq 1$ is a constant such that $\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$.
@@
 $\frac{\log N}{\log(N/(1+\delta))}<1+\frac{\epsilon_{2}}{\log 1/\epsilon},$ (3.5)
-it holds that $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$ for every choice of $(r_{i})_{i}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i$.
+it holds that $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$ for every choice of $(r_{i})_{i\in\mathcal{I}_{N}}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$.
 Proof.
@@
 and
-$p_{0}(y;z):=(G_{i}(y))_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}}.$
+$p_{0}(y;z):={\bf G}(y)\in\overline{\mathcal{P}_{N}}.$
 By a martingale argument, we can show that
@@
 This assertion is easily shown.
-If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then, $s_{N}\left((G_{k}(y))_{k}|{\bf q}\right)<0$.
+If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then, $s_{N}\left({\bf G}(y)|{\bf q}\right)<0$.
 However, in our setting, we cannot exclude the possibility that $|G_{k}(y)-q_{k}|=0$ for every $k\in\mathcal{I}_{N}$.
@@
 The proof is essentially the same as the proof of [27, Lemma 8] [^3] It contains a typo.. Since $X_{j+1}(z)=i_{1}$,
-$p_{j}(0;z)=\left(G_{k}(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))\right)_{k\in\mathcal{I}_{N}},$
+$p_{j}(0;z)={\bf G}\left(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0)\right),$
 and
-$p_{j+1}(0;z)=\left(G_{k}(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))\right)_{k\in\mathcal{I}_{N}}.$
+$p_{j+1}(0;z)={\bf G}\left(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0)\right).$
 Let ${\bf q}=(q_{k})_{k}$. Then, by (3.8), either
@@
 $W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{2}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|.$
-If $(G_{j}(y))_{j}\notin B_{\mathscr{A}({\bf q})}$, then, $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}=\left(G_{j}(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)}))\right)_{j}$ and hence, $y$ is not contained in an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$.
+If ${\bf G}(y)\notin B_{\mathscr{A}({\bf q})}$, then, $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}={\bf G}\left(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})\right)$ and hence, $y$ is not contained in an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$.
 Since $G_{j}$ and $H_{\mathscr{A}\left({\bf q}\right)}$ are continuous, $\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$ is continuous with respect to $y$ on $Y$.
 Hence,
-$\inf_{(G_{j}(y))_{j}\notin B_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
+$\inf_{{\bf G}(y)\notin B_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
 $\geq\inf_{y\notin J_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|=:\epsilon_{\mathscr{A}\left({\bf q}\right)}>0.$
-Assume that $(G_{j}(y))_{j}\in B_{\mathscr{A}({\bf q})}$. Since ${\bf q}\notin B_{\mathscr{A}({\bf q})}$,
+Assume that ${\bf G}(y)\in B_{\mathscr{A}({\bf q})}$. Since ${\bf q}\notin B_{\mathscr{A}({\bf q})}$,
 $W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{N}\sum_{j}|G_{j}(y)-q_{j}|\geq\frac{\delta_{0}}{4N}.$
@@
 Since $\min_{i}r_{i}\geq\epsilon$, $\sum_{i}|r_{i}-r_{i}^{s}|\leq\epsilon^{1-s}-1\leq\epsilon/2$.
-Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$, where we let $\widetilde{q}_{i}:=r_{i}^{s}$ for each $i$ and ${\bf\widetilde{q}}:=(\widetilde{q}_{i})_{i}$.
-Since $\epsilon\leq 1/2$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\epsilon/2$ for each $i$.
+Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$, where we let $\widetilde{q}_{i}:=r_{i}^{s}$ for each $i\in\mathcal{I}_{N}$ and ${\bf\widetilde{q}}:=\left(\widetilde{q}_{i}\right)_{i\in\mathcal{I}_{N}}$.
+Since $\epsilon\leq 1/2$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\epsilon/2$ for each $i\in\mathcal{I}_{N}$.
 Hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
@@
 4. Example
-We give an example that $\{g_{i}\}_{i}$ is an LF system and $\{f_{i}\}_{i}$ is non-affine and the solution $\varphi$ is singular. The result depends on the quantitative estimates in Section 3.
-Let $N=2$. Then,
-$G_{0}(y)=\frac{b_{1}(y+1)}{b_{1}y+1},\ H_{0}(y)=a_{0}y+c_{0},\ H_{1}(y)=\frac{(1-b_{1}+c_{1})y+c_{1}}{b_{1}y+1},$
+We give three examples that $\{g_{i}\}_{i}$ is an LF system and $\{f_{i}\}_{i}$ is non-affine and the solution $\varphi$ is singular. The result depends on the quantitative estimates in Section 3.
+We first deal with the case that $N=2$. Then,
+$G_{0}(y)=\frac{b_{1}(y+1)}{b_{1}y+1},\ H_{0}(y)=a_{0}y+c_{0},\ H_{1}(y)=\frac{(1-b_{1}+c_{1})y+c_{1}}{b_{1}y+1}.$
+Example 4.1.
 Let
-$g_{0}(x):=\frac{x}{-x+6},\ g_{1}(x):=\frac{3x+1}{5-x}.$
-Then, $g_{0}$ is contractive, and $g_{1}$ is weakly contractive but not contractive. It holds that $a_{0}=1/6$, $c_{0}=-1/6$, $a_{1}=3/5$, $b_{1}=1/5$, and $c_{1}=-1/5$. Hence,
-$\alpha=\min\{0,-1/5,-1\}=-1,\beta=\max\{0,-1/5,-1\}=0.$
+$g_{0}(x):=\frac{x}{-x+6},\textup{ and }g_{1}(x):=\frac{3x+1}{5-x}.$
+Then, $g_{0}$ is contractive, and $g_{1}$ is weakly contractive but not contractive. We see that
+$\alpha=\min\left\{0,-\frac{1}{5},-1\right\}=-1,\beta=\left\{0,-\frac{1}{5},-1\right\}=0.$
+We see that
+$G_{0}(y)=\frac{y+1}{y+5},\textup{ and }G_{1}(y)=1-G_{0}(y)=\frac{4}{y+5}.$
 Hence,
 $\sup_{y\in[-1,0]}G_{0}(y)=1-\inf_{y\in[-1,0]}G_{1}(y)=\frac{1}{5}.$
-Let $\mathcal{P}[4/5]:=\left\{(p_{0},p_{1})\in\mathcal{P}_{2}:p_{0}\geq 4/5\right\}$. Let ${\bf q}=(q_{0},q_{1})\in\mathcal{P}[4/5]$. Then,
-$|G_{0}(y)-q_{0}|=|G_{1}(y)-q_{1}|\geq\frac{3}{5},\ y\in[-1,0].$
+It also holds that
+$H_{0}(y)=\frac{y-1}{6}\textup{ and }H_{1}(y)=\frac{3y-1}{y+5}.$
+Hence, $\textup{Fix}(H_{0})=\{-1/5\}$ and $\textup{Fix}(H_{1})=\{-1\}$. Hence,
+${\bf G}(\textup{Fix}(H_{0}))=\left(\frac{1}{6},\frac{5}{6}\right)\textup{ and }{\bf G}(\textup{Fix}(H_{1}))=(0,1).$
+Let $\widetilde{\mathcal{P}_{2}}:=\left\{(p_{0},p_{1})\in\mathcal{P}_{2}:p_{0}\geq 4/5\right\}$. Then,
+$|G_{0}(y)-q_{0}|=|G_{1}(y)-q_{1}|\geq\frac{3}{5},\ y\in[-1,0],\ {\bf q}=(q_{0},q_{1})\in\widetilde{\mathcal{P}_{2}}.$
 Hence,
-$s_{2}\left(p_{j}(0,z)|{\bf q}\right)\leq V\left({\bf q};\frac{6}{5}\right),\ j\geq 0,z\in(\mathcal{I}_{N})^{\mathbb{N}}.$ (4.1)
+$s_{2}\left({\bf G}(y)|{\bf q}\right)\leq V\left({\bf q};\frac{6}{5}\right),\ y\in[-1,0],\ {\bf q}\in\widetilde{\mathcal{P}_{2}}.$ (4.1)
 Since $s_{2}({\bf p}|{\bf q})=s_{2}\left((p_{0},1-p_{0})|(q_{0},1-q_{0})\right)$ is increasing with respect to $p_{0}$ on $[0,q_{0}]$,
@@
 By Proposition 3.6, it holds that
-$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.83,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\mathcal{P}[4/5]$.}$ (4.2)
-Let $\epsilon\in(0,1/2)$.
+$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.83,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{2}}$.}$ (4.2)
+Let $\epsilon\in(0,1/4)$.
 Let $\{f_{0},f_{1}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1$.
 Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}=1$.
-Assume that $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$.
+Assume that $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$.
 Then, by (4.2) and Lemma 3.5,
@@
 We see that $\dfrac{0.83}{\log 10}\geq 0.36$ and $\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$.
-Hence, (3.5) holds for $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$.
-By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, $r_{0}+r_{1}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$, then,
+Hence, (3.5) holds for $N=2$, $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$.
+By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, $r_{0}+r_{1}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$, then,
 $\dim_{H}\mu_{\varphi}\leq s-\frac{0.83}{\log(1/\epsilon)}\leq\dfrac{\log 2}{\log 2-\log 1.2}-\frac{0.83}{\log 10}<1.$
-Now we can let $\{f_{0},f_{1}\}$ be a compatible system such that $0.85\leq r_{0}\leq 0.9$ and $0.1\leq r_{1}\leq 0.3$.
+Now we can let $\{f_{0},f_{1}\}$ be a compatible system such that $0.85\leq r_{0}\leq 0.9$, $0.1\leq r_{1}\leq 0.3$ and $r_{0}+r_{1}\geq 1$.
 Then, $r_{0}+r_{1}\leq 1.2$.
-By the compatibility condition, $r_{0}+r_{1}\geq 1$.
-The maps $f_{0}$ and $f_{1}$ may not be affine.
+The maps $f_{0}$ and $f_{1}$ can be non-affine.
 Since $s\leq\frac{\log 2}{\log 2-\log 1.2}\leq 1.357$, $r_{0}^{s}\geq(0.85)^{1.357}\geq 0.8$.
-Hence, $(r_{0}^{s},r_{1}^{s})\in\mathcal{P}[4/5]$.
-Remark 4.1.
-We see that $a_{0}<1$ and $b_{1}+c_{1}=0$. Hence, this case corresponds to Case 2.2. However, we do not need to apply Lemma 3.11 since we have the upper bound for the relative entropy in (4.1).
+Hence, $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$.
+In this case, we see that $a_{0}<1$ and $b_{1}+c_{1}=0$.
+Hence, this case corresponds to Case 2.2.
+However, we do not need to apply Lemma 3.11 since we have the upper bound for the relative entropy in (4.1).
+Example 4.2.
+Let $N=3$. Let
+$g_{0}(x):=\frac{x}{8-x},\ g_{1}(x):=\frac{5x+1}{7},\textup{ and }g_{2}(x):=\frac{6}{7-x}.$
+Then,
+$\alpha=\min\left\{0,-\frac{1}{7},0,-\frac{1}{6}\right\}=-\frac{1}{6}\textup{ and }\beta=\max\left\{0,-\frac{1}{7},0,-\frac{1}{6}\right\}=0.$
+Hence, $Y=[-1/6,0]$. We see that
+$G_{0}(y)=\frac{y+1}{y+7},G_{1}(y)=\frac{35(y+1)}{(6y+7)(y+7)},\textup{ and }G_{2}(y)=\frac{1}{6y+7}.$
+Hence,
+$\sup_{y\in[-1/6,0]}G_{0}(y)=\frac{1}{7},\textup{ and }\sup_{y\in[-1/6,0]}G_{2}(y)=\frac{1}{6}.$
+We also see that
+$H_{0}(y)=\frac{y-1}{8},H_{1}(y)=\frac{5y}{y+7},\textup{ and }H_{2}(y)=-\frac{1}{6y+7}.$
+Hence, $\textup{Fix}(H_{0})=\{-1/7\}$, $\textup{Fix}(H_{1})=\{0\}$ and $\textup{Fix}(H_{2})=\{-1/6\}$. Hence,
+${\bf G}(\textup{Fix}(H_{0}))=\left(\frac{1}{8},\frac{245}{344},\frac{7}{43}\right),{\bf G}(\textup{Fix}(H_{1}))=\left(\frac{1}{7},\frac{5}{7},\frac{1}{7}\right),\textup{ and }{\bf G}(\textup{Fix}(H_{2}))=\left(\frac{5}{41},\frac{175}{246},\frac{1}{6}\right).$
+Let
+$\widetilde{\mathcal{P}_{3}}:=\left\{(p_{0},p_{1},p_{2})\in\overline{\mathcal{P}_{3}}\ \middle|\ p_{0},p_{2}\in\left[\frac{7}{16},\frac{1}{2}\right]\right\}.$
+We now show that
+$s_{3}\left({\bf G}(y)|{\bf q}\right)\leq-0.74,\ y\in[-1/6,0],\ {\bf q}\in\widetilde{\mathcal{P}_{3}}.$ (4.3)
+For $q>0$, let $f_{q}(x):=x\log(x/q)$ for $x\in[0,q]$.
+Then, $f_{q}$ is increasing on $[0,q/e]$ and decreasing on $[q/e,1]$.
+Since ${\bf q}=(q_{0},q_{1},q_{2})\in\widetilde{\mathcal{P}_{3}}$, it holds that $q_{0},q_{2}\leq 1/2$ and $q_{1}\leq 1/8$.
+Therefore, for $y\in[-1/6,0]$,
+$s_{3}\left({\bf G}(y)|{\bf q}\right)\leq f_{1/2}(G_{0}(y))+f_{1/8}(G_{1}(y))+f_{1/2}(G_{2}(y))$
+$\leq 2f_{1/2}(1/6)+f_{1/8}(2/3)=\log 3-\frac{8}{3}\log 2\leq-0.74.$
+Thus we have (4.3).
+By Proposition 3.6, it holds that
+$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.74,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{3}}$.}$ (4.4)
+Let $\epsilon\in(0,1/4)$.
+Let $\{f_{0},f_{1},f_{2}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1,2$.
+Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}+r_{2}^{s}=1$.
+Assume that $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$.
+Then, by (4.4) and Lemma 3.5,
+$\dim_{H}\mu_{\varphi}\leq s-\frac{0.74}{\log(1/\epsilon)}.$
+We see that $\dfrac{0.74}{\log 20}>0.24$ and $\dfrac{\log 3}{\log 3-\log 1.1}<1.1$.
+Hence, (3.5) holds for $N=3$, $\epsilon=0.05$, $\epsilon_{2}=0.74$ and $\delta=0.1$.
+By Lemma 3.4, if $\{f_{0},f_{1},f_{2}\}$ is a contractive system such that $0.05\leq r_{i}\leq 0.95$, $i=0,1,2$, $r_{0}+r_{1}+r_{2}\leq 1.1$, and $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$, then,
+$\dim_{H}\mu_{\varphi}\leq s-\frac{0.74}{\log(1/\epsilon)}\leq\dfrac{\log 3}{\log 3-\log 1.1}-\frac{0.74}{\log 20}<1.$
+Now we can let $\{f_{0},f_{1},f_{2}\}$ be a compatible system such that $0.475\leq r_{0}\leq 0.5$, $0.05\leq r_{1}\leq 0.1$, and $0.475\leq r_{2}\leq 0.5$.
+Then, $1\leq r_{0}+r_{1}+r_{2}\leq 1.1$.
+The maps $f_{0}$, $f_{1}$ and $f_{2}$ can be non-affine.
+Since $s\leq\frac{\log 3}{\log 3-\log 1.1}\leq 1.1$, $\min\{r_{0}^{s},r_{2}^{s}\}\geq(0.475)^{1.1}\geq\frac{7}{16}$.
+Hence, $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$.
+Example 4.3.
+Let $N=4$. Let
+$g_{0}(x)=\frac{x}{3},\ g_{1}(x)=\frac{x+2}{6},\ g_{2}(x)=\frac{x+3}{6},\textup{ and }g_{3}(x)=\frac{x+2}{3}.$
+Then, $\alpha=\beta=0$, and hence, $Y=\{0\}$. We see that
+$G_{0}(0)=\frac{1}{3},\ G_{1}(0)=\frac{1}{6},\ G_{2}(0)=\frac{1}{6},\textup{ and }G_{3}(0)=\frac{1}{3}.$
+Let
+$\widetilde{\mathcal{P}_{4}}:=\left\{(p_{0},p_{1},p_{2},p_{3})\in\overline{\mathcal{P}_{4}}\ \middle|\ p_{0},p_{3}\in\left[0,\frac{1}{8}\right]\right\}.$
+Then, for ${\bf q}=(q_{0},q_{1},q_{2},q_{3})\in\widetilde{\mathcal{P}_{4}}$,
+$s_{4}\left({\bf G}(0)|{\bf q}\right)=\log 3+\frac{1}{3}\log 2+\frac{1}{6}(2\log q_{0}+\log q_{1}+\log q_{2}+2\log q_{3}).$
+Since $\sum_{i=0}^{3}q_{i}=1$, $q_{1}+q_{2}=1-q_{0}-q_{3}$, and hence, $\log q_{1}+\log q_{2}\leq 2\log((1-q_{0}-q_{3})/2)$. Therefore,
+$s_{4}\left({\bf G}(0)|{\bf q}\right)\leq\log 3+\frac{1}{3}\left(\log q_{0}+\log q_{3}+\log(1-q_{0}-q_{3})\right).$
+Let $f(x,y):=\log x+\log y+\log(1-x-y)$ for $x,y\in[0,1/8]$. Then,
+$\max_{x,y\in[0,1/8]}f(x,y)=f\left(\frac{1}{8},\frac{1}{8}\right)=\log 3-8\log 2.$
+Therefore,
+$s_{4}\left({\bf G}(0)|{\bf q}\right)\leq\frac{4}{3}\log\frac{3}{4}\leq-0.38,\ {\bf q}\in\widetilde{\mathcal{P}_{4}}.$
+By Proposition 3.6, it holds that
+$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.38,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{4}}$.}$ (4.5)
+Let $\epsilon\in(0,1/4)$.
+Let $\{f_{0},f_{1},f_{2},f_{3}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1,2,3$.
+Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}+r_{2}^{s}+r_{3}^{s}=1$.
+Assume that $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$.
+Then, by (4.5) and Lemma 3.5,
+$\dim_{H}\mu_{\varphi}\leq s-\frac{0.38}{\log(1/\epsilon)}.$
+We see that $\dfrac{0.38}{\log 10}>0.165$ and $1.1521.15<\dfrac{\log 4}{\log 4-\log 1.2}<$ 1.152.
+Hence, (3.5) holds for $N=4$, $\epsilon=0.1$, $\epsilon_{2}=0.38$ and $\delta=0.2$.
+By Lemma 3.4, if $\{f_{0},f_{1},f_{2},f_{3}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1,2,3$, $r_{0}+r_{1}+r_{2}+r_{3}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$, then,
+$\dim_{H}\mu_{\varphi}\leq s-\frac{0.38}{\log(1/\epsilon)}\leq\dfrac{\log 4}{\log 4-\log 1.2}-\frac{0.38}{\log 10}<1.$
+Now we can let $\{f_{0},f_{1},f_{2},f_{3}\}$ be a compatible system such that $0.1\leq r_{0}\leq 0.125$, $0.4\leq r_{1}\leq 0.475$, $0.4\leq r_{2}\leq 0.475$, $0.1\leq r_{3}\leq 0.125$.
+Then, $1\leq r_{0}+r_{1}+r_{2}+r_{3}\leq 1.2$.
+The maps $f_{0}$, $f_{1}$, $f_{2}$ and $f_{3}$ can be non-affine.
+Since $s\geq 1$, $\max\{r_{0}^{s},r_{3}^{s}\}\leq 1/8$ and hence $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$.

### v4

@@
 This framework contains de Rham’s functional equations.
-We give sufficient conditions for singularity of the solution with quantitative estimates in the case that the equation is driven by non-affine maps, in particular, linear fractional transformations.
+We give sufficient conditions for singularity of the solution with quantitative estimates in the case where the equation is driven by a family of non-affine maps and a family of linear fractional transformations.
 1. Introduction
-Let $N\geq 2$ and let $\mathcal{I}_{N}:=\{0,1,\cdots,N-1\}$.
+Let $N\geq 2$ and let $\mathcal{I}_{N}\coloneqq\{0,1,\ldots,N-1\}$.
 Let $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ be two families of $N$ maps on $[0,1]$ satisfying a compatibility condition.
-We consider the following conjugate equation for $\varphi:[0,1]\to[0,1]$:
-$g_{i}\circ\varphi=\varphi\circ f_{i},\ i\in\mathcal{I}_{N}.$ (1.1)
-See Section 2 below for the definition of the compatibility condition. An alternative expression is given by
-$\varphi(x)=\begin{cases}g_{0}(\varphi(f_{0}^{-1}(x)))&0\leq x\leq f_{0}(1)\\ \vdots&\\ g_{N-1}(\varphi(f_{N-1}^{-1}(x)))&f_{N-1}(0)\leq x\leq 1\end{cases}.$ (1.2)
+We consider the following conjugate equation for $\varphi\colon[0,1]\to[0,1]$:
+(1.1) $g_{i}\circ\varphi=\varphi\circ f_{i},\ i\in\mathcal{I}_{N}.$
+See the beginning of Section 2 below for the definition of the compatibility condition. An alternative expression is given by
+(1.2) $\varphi(x)=\begin{cases}g_{0}(\varphi(f_{0}^{-1}(x)))&0=f_{0}(0)\leq x\leq f_{0}(1)\\ \vdots&\\ g_{N-1}(\varphi(f_{N-1}^{-1}(x)))&f_{N-1}(0)\leq x\leq f_{N-1}(1)=1\end{cases}.$
 This is known as de Rham’s functional equation [9, 10]. This functional equation is generally fractal in nature, and solutions of (1.1) can be singular functions.
-Assume that $N=2$, $10<p<$ 1, $f_{0}(x)=x/2$, $f_{1}(x)=(x+1)/2$, $g_{0}(x)=px$ and $g_{1}(x)=(1-p)x+p$. Then, (1.2) is expressed as
+Assume that $N=2$, $0<p<1$, $f_{0}(x)=x/2$, $f_{1}(x)=(x+1)/2$, $g_{0}(x)=px$ and $g_{1}(x)=(1-p)x+p$. Then, (1.2) is expressed as
 $\varphi(x)=\begin{cases}p\varphi(2x)&0\leq x\leq 1/2\\ (1-p)\varphi(2x-1)+p&1/2\leq x\leq 1\end{cases}.$
-There exists a unique continuous strictly increasing solution $\varphi_{p}$ for this equation, and $\varphi_{p}$ is singular if $p\neq 1/2$ and $\varphi_{p}(x)=x$ if $p=1/2$.
-$\varphi_{p}$ is called the Lebesgue singular function.
-This function is also a solution of the functional equation by Matkowski and Wesołowski [16, 24, 25].
-Minkowski’s question-mark function [23] is also obtained by a solution of (1.1) for the case that $(f_{0}(x),f_{1}(x),g_{0}(x),g_{1}(x))=\left(x/(x+1),1/(2-x),x/2,(x+1)/2\right)$.
+There exists a unique continuous strictly increasing solution $\varphi_{p}$ for this equation, and the function $\varphi_{p}$ is singular if $p\neq 1/2$, and $\varphi_{p}(x)=x$ if $p=1/2$.
+See [34, Subsection 3.4] for the proof.
+The function $\varphi_{p}$ is often called the Lebesgue singular function.
+This function is sometimes referred to by the name of Riesz and Nágy ([1, 3]), Salem ([8]), or Takács ([3]).
+This function is also a solution of the functional equation by Matkowski and Wesołowski [24, 25].
+Minkowski’s question-mark function was initially considered in [23].
+This function can be obtained by a solution of (1.1) for the case where $f_{0}(x)=x/(x+1)$, $f_{1}(x)=1/(2-x)$, $g_{0}(x)=x/2$, and $g_{1}(x)=(x+1)/2$.
 It can also be expressed in the form
@@
 We refer the reader to [27, Subsection 3.4] for a more detailed review of de Rham’s functional equations.
-The main purpose of this paper is to investigate some properties of the solution of (1.1).
-Specifically, we give sufficient conditions for the solution to be singular.
+The main purpose of this paper is to investigate sufficient conditions for the solution of (1.1) to be singular.
+Necessary and sufficient conditions of $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ for the singularity of the solution have been considered in [4, 5, 12, 15, 18, 25].
+We consider the case where both $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are non-affine maps.
+This case is more difficult to analyze than the case where both $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are affine maps.
+In [27], the author dealt with the case where $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are linear fractional transformations and $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ are affine maps.
+Our results, Theorems 3.2 and 3.3 below, deal with the case where $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ are not necessarily affine maps.
+We do not assume the differentiability of $\{f_{i}\}_{i\in\mathcal{I}_{N}}$.
 Since $\varphi$ is an increasing function on $[0,1]$, we deal with the probability measure $\mu_{\varphi}$, which is the Borel probability measure on $[0,1]$ whose distribution function is $\varphi$.
-We consider the case that both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are non-affine.
-This case is more difficult to analyze than the case that both $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are affine.
-It is interesting to know a necessary and sufficient condition of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ for the singularity.
-This topic has been considered in [2, 3, 10, 13, 18, 25].
-In [27], the author dealt with the case that $\{g_{i}\}_{i}$ are linear fractional transformations and $\{f_{i}\}_{i}$ are affine.
-Our results, Theorems 3.2 and 3.3 below, deal with the case that $\{f_{i}\}_{i}$ are not necessarily affine.
-We do not assume the differentiability of $\{f_{i}\}_{i}$.
-For the proof, we use a martingale approach as in [27], but here we need more delicate quantitative estimates.
+We give sufficient conditions for $\dim_{H}\mu_{\varphi}<1$, which implies the singularity for $\varphi$.
+For the proof, we use an approach using martingales as in [27], but here we need more delicate quantitative estimates for $\dim_{H}\mu_{\varphi}$.
 The structure of this paper is as follows.
 In Section 2, we give some terminology and basic results of the conjugate equation on the unit interval.
-In Section 3, we deal with the singularity.
+In Section 3, we give our main results and their proofs.
 Finally, we give three numerical examples in Section 4.
@@
 We first discuss the existence and uniqueness of the solution of the conjugate equation (1.1).
-We say that a family of maps $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ is a compatible system or satisfies a compatibility condition if they are strictly increasing continuous functions such that $h_{0}(0)=0$, $h_{i-1}(1)=h_{i}(0)$, $11\leq i\leq N-$ 1, $h_{N-1}(1)=1$, and, a compatible system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system on $[0,1]$ if
-$\mathcal{D}_{h}:=\{h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j):i_{1},\cdots i_{n}\in\mathcal{I}_{N},j\in\{0,1\}\}$
+We say that a family of maps $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ is a compatible system or that it satisfies a compatibility condition if they are strictly increasing continuous functions such that $h_{0}(0)=0$, $h_{i-1}(1)=h_{i}(0)$, $1\leq i\leq N-1$, $h_{N-1}(1)=1$, and that a compatible system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system on $[0,1]$ if
+$\mathcal{D}_{h}\coloneqq\left\{h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j)\colon i_{1},\ldots,i_{n}\in\mathcal{I}_{N},j\in\{0,1\}\right\}$
 is dense in $[0,1]$.
@@
 Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be a D-system on $[0,1]$. Then, for every $x\in[0,1]$, there exists an infinite sequence $(i_{n})_{n}$ of $\mathcal{I}_{N}$ such that
-$x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1).$ (2.1)
-If $x\in\mathcal{D}_{h}\setminus\{0,1\}$, then, there exist two such infinite sequences. Otherwise, there exists exactly one such infinite sequence.
+(2.1) $x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1).$
+If $x\in\mathcal{D}_{h}\setminus\{0,1\}$, then there exist two such infinite sequences. Otherwise, there exists exactly one such infinite sequence.
 Proof.
 We first show that for every infinite sequence $(i_{n})_{n}$,
-$\lim_{n\to\infty}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)|=0.$ (2.2)
+(2.2) $\lim_{n\to\infty}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)|=0.$
 Assume this fails.
-By the definition of $\{h_{i}\}_{i}$, $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right)_{n}$ is increasing and $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right)_{n}$ is decreasing.
-Since $\mathcal{D}_{h}$ is dense in $[0,1]$, there exists $k_{1},\cdots,k_{N}$ and $j\in\{0,1\}$ such that
-$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)<h_{k_{1}}\circ\cdots\circ h_{k_{N}}(j)$
-$<\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
-This cannot occur. Hence (2.2) holds.
+By the definition of $\{h_{i}\}_{i\in\mathcal{I}_{N}}$, the sequence $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right)_{n}$ is increasing and $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right)_{n}$ is decreasing.
+Since $\mathcal{D}_{h}$ is dense in $[0,1]$, there exist $M\in\mathbb{N}$ and $k_{1},\ldots,k_{M}$ as well as $j\in\{0,1\}$ such that
+$h_{i_{1}}\circ\cdots\circ h_{i_{M}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)<h_{k_{1}}\circ\cdots\circ h_{k_{M}}(j)$
+$\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{M}}(1).$
+This cannot occur since the sequence $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j)\right)_{i_{1},\dots,i_{n},j}$ is in lexicographic order, more precisely, $h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j)\leq h_{i_{1}^{\prime}}\circ\cdots\circ h_{i_{n}^{\prime}}(j^{\prime})$ if and only if $\displaystyle j+\sum_{\ell=1}^{n}i_{\ell}N^{n-\ell}\leq j^{\prime}+\sum_{\ell=1}^{n}i_{\ell}^{\prime}N^{n-\ell}$.
+Hence (2.2) holds.
 Case 1.
@@
 Since $h_{0}(0)=0$ and $x>0$, we can assume that $i_{n}\geq 1$.
-Let $i_{k}:=0$ for every $k\geq n+1$.
-Then, (2.1) holds.
-By replacing $i_{n}$ with $i_{n}-1$ and let $i_{k}:=N-1$ for every $k\geq n+1$, then, (2.1) holds.
+Let $i_{k}\coloneqq 0$ for every $k\geq n+1$.
+Then, the first equality of (2.1) holds.
+By this and (2.2), the second equality of (2.1) holds.
+Thus (2.1) holds.
+By replacing $i_{n}$ with $i_{n}-1$ and letting $i_{k}\coloneqq N-1$ for every $k\geq n+1$, the equations in (2.1) hold.
 Assume that $x=h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)$.
 Since $h_{N-1}(1)=1$ and $x<1$, we can assume that $i_{n}\leq N-2$.
-Let $i_{k}:=N-1$ for every $k\geq n+1$.
+Let $i_{k}\coloneqq N-1$ for every $k\geq n+1$.
 Then, (2.1) holds.
-By replacing $i_{n}$ with $i_{n}+1$ and let $i_{k}:=0$ for every $k\geq n+1$, then, (2.1) holds.
+It remains valid if we replace $i_{n}$ with $i_{n}+1$ and let $i_{k}\coloneqq 0$ for every $k\geq n+1$.
+There are no more solutions of the infinite sequences satisfying (2.1) since the sequence $\left(h_{i_{1}}\circ\cdots\circ h_{i_{n}}(j)\right)_{i_{1},\dots,i_{n},j}$ is in lexicographic order.
 Case 2.
@@
 Then, there is no other choice than $i_{k}=N-1$ for every $k\geq 1$ for (2.1).
-Case 3. Assume $x\notin\mathcal{D}_{h}$. Then, for each $n\geq 1$, there exists a unique finite sequence $(i_{n,1},\cdots,i_{n,n})$ such that
+Case 3. Assume $x\notin\mathcal{D}_{h}$. Then, for each $n\geq 1$, there exists a unique finite sequence $(i_{n,1},\ldots,i_{n,n})$ such that
 $h_{i_{n,1}}\circ\cdots\circ h_{i_{n,n}}(0)<x<h_{i_{n,1}}\circ\cdots\circ h_{i_{n,n}}(1).$
-Then, $i_{n-1,k}=i_{n,k}$ for every $k\leq n-1$. Let $i_{k}:=i_{n,k}$. By (2.2), (2.1) holds. If (2.1) holds, then, for every $N\geq 1$,
-$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
+Then, $i_{n-1,k}=i_{n,k}$ for every $k\leq n-1$. Let $i_{k}\coloneqq i_{n,k}$. By (2.2), the equations in (2.1) hold. If (2.1) holds, then for every $M\geq 1$,
+$h_{i_{1}}\circ\cdots\circ h_{i_{M}}(0)\leq\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)=x=\lim_{n\to\infty}h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\leq h_{i_{1}}\circ\cdots\circ h_{i_{M}}(1).$
 Hence,
-$h_{i_{1}}\circ\cdots\circ h_{i_{N}}(0)<x<h_{i_{1}}\circ\cdots\circ h_{i_{N}}(1).$
+$h_{i_{1}}\circ\cdots\circ h_{i_{M}}(0)<x<h_{i_{1}}\circ\cdots\circ h_{i_{M}}(1).$
 Hence the uniqueness holds. ∎
@@
 Proof.
-We first see that
-$\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(j)\right)=g_{i_{1}}\circ\cdots\circ g_{i_{n}}(j),\ i_{1},\cdots,i_{n}\in\mathcal{I}_{N},\ j\in\{0,1\},$ (2.3)
-by induction in $n$. Hence, $\varphi(\mathcal{D}_{f})=\mathcal{D}_{g}$.
-Step 1.
+First we show the existence.
+Step 1. We define $\varphi(x)$ for $x\in\mathcal{D}_{f}$. Let
+(2.3) $\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(j)\right)\coloneqq g_{i_{1}}\circ\cdots\circ g_{i_{n}}(j),\ i_{1},\ldots,i_{n}\in\mathcal{I}_{N},\ j\in\{0,1\}.$
+This is well-defined since $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are compatible systems. Hence, $\varphi(\mathcal{D}_{f})=\mathcal{D}_{g}$.
+Step 2.
 We define $\varphi(x)$ for $x\notin\mathcal{D}_{f}$.
-By the assumption, there exists an infinite sequence $(i_{n})_{n}$ of $\mathcal{I}_{N}$ such that $x\in(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1))$.
+By the assumption and Lemma 2.1, there exists a unique infinite sequence $(i_{n})_{n}$ of $\mathcal{I}_{N}$ such that $x\in(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1))$.
 Since $\left[g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0),g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)\right]$ is decreasing with respect to $n$,
@@
 By the density of $\mathcal{D}_{g}$ in $[0,1]$, we can apply (2.2), and obtain that
-$\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1).$ (2.4)
-We let $\varphi(x):=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)$.
-Step 2. We show $\varphi$ is strictly increasing on $[0,1]$. By (2.3), $\varphi$ is strictly increasing on $\mathcal{D}_{f}$. Let $x<y$. By (2.4), for some $n$,
+$\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)=\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1).$
+We let $\displaystyle\varphi(x)\coloneqq\lim_{n\to\infty}g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)$.
+Step 3.
+We show that $\varphi$ is strictly increasing on $[0,1]$.
+By (2.3), $\varphi$ is strictly increasing on $\mathcal{D}_{f}$.
+Let $x<y$.
+By the density of $\mathcal{D}_{f}$ in $[0,1]$, we can apply (2.2), and obtain that
+$\lim_{n\to\infty}f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0)=\lim_{n\to\infty}f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1).$
+Hence it holds that for some $n$ and $i_{1},\dots,i_{n},j_{1},\dots,j_{n}\in\mathcal{I}_{N}$,
 $f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0)\leq x\leq f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)<f_{j_{1}}\circ\cdots\circ f_{j_{n}}(0)\leq y\leq f_{j_{1}}\circ\cdots\circ f_{j_{n}}(1).$
-By this and Step 1,
+By this and Step 2,
 $\varphi(x)\leq g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)<g_{j_{1}}\circ\cdots\circ g_{j_{n}}(0)\leq\varphi(y).$
-Step 3. We show that $\varphi$ is continuous on $[0,1]$. Assume that there is a discontinuous point $x_{0}\in(0,1)$ of $\varphi$. Then,
-$b_{0}:=\lim_{x\to x_{0}+0}\varphi(x)>\lim_{x\to x_{0}-0}\varphi(x)=:a_{0}.$
-By Step 2, $\varphi([0,1])\cap(a_{0},b_{0})=\emptyset$, in particular, $\mathcal{D}_{g}\cap(a_{0},b_{0})=\emptyset$.
+Step 4. We show that $\varphi$ is continuous on $[0,1]$. Assume that there is a discontinuous point $x_{0}\in(0,1)$ of $\varphi$. Then,
+$b_{0}\coloneqq\lim_{x\to x_{0}+0}\varphi(x)>\lim_{x\to x_{0}-0}\varphi(x)\eqqcolon a_{0}.$
+By Step 3, $\varphi([0,1])\cap(a_{0},b_{0})=\emptyset$, in particular, $\mathcal{D}_{g}\cap(a_{0},b_{0})=\emptyset$.
 This contradicts the density of $\mathcal{D}_{g}$ in $[0,1]$.
-The same argument is applied to the case that $x_{0}\in\{0,1\}$.
-∎
-It is natural to consider the inverse function of the solution $\varphi$. We easily see that
+The same argument is applied to the case where $x_{0}\in\{0,1\}$.
+Second we show the uniqueness.
+Step 1.
+We show that if $\varphi$ is a solution of (1.1), then $\varphi(j)=j,j\in\{0,1\}$.
+By (1.1) and the assumption that $f_{0}(0)=0$, $\varphi(0)=\varphi(f_{0}(0))=g_{0}(\varphi(0))$.
+Denote the $n$-fold composition of $g_{0}$ by $g_{0}^{n}$.
+Then, by induction in $n$, we see that $\varphi(0)=g_{0}^{n}(\varphi(0))$ for every $n\geq 1$.
+Since $g_{0}$ is an increasing function on $[0,1]$, $g_{0}^{n}$ is also an increasing function on $[0,1]$ for each $n$.
+Therefore, $0=g_{0}^{n}(0)\leq g_{0}^{n}(\varphi(0))\leq g_{0}^{n}(1)$.
+By this and (2.2),
+$\varphi(0)=\lim_{n\to\infty}g_{0}^{n}(\varphi(0))=\lim_{n\to\infty}g_{0}^{n}(1)=0.$
+In the same manner, we can show that $\varphi(1)=1$.
+Step 2. Since $\varphi(j)=j,j\in\{0,1\}$, we see that
+(2.4) $\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(j)\right)=g_{i_{1}}\circ\cdots\circ g_{i_{n}}(j),\ i_{1},\ldots,i_{n}\in\mathcal{I}_{N},\ j\in\{0,1\},$
+by induction in $n$.
+Since $\mathcal{D}_{f}$ is dense in $[0,1]$ and $\varphi$ is assumed to be continuous, we obtain the uniqueness. ∎
+We give an example of a compatible system $\{h_{0},h_{1}\}$ which is not a $D$-system. This is suggested by an anonymous referee.
+Example 2.4.
+Let
+$h_{0}(x)\coloneqq\begin{cases}\frac{1}{2}x&0\leq x\leq\frac{1}{4},\\ 2x-\frac{3}{8}&\frac{1}{4}\leq x\leq\frac{1}{2},\\ \frac{3}{8}x+\frac{7}{16}&\frac{1}{2}\leq x\leq 1,\end{cases}$
+and
+$h_{1}(x)\coloneqq\frac{3}{16}x+\frac{13}{16},\qquad 0\leq x\leq 1.$
+Then the functions $h_{0}$ and $h_{1}$ are strictly increasing and continuous. Moreover,
+$0=h_{0}(0)<h_{0}(1)=\frac{13}{16}=h_{1}(0)<h_{1}(1)=1.$
+Hence, $\{h_{0},h_{1}\}$ is a compatible system.
+We also observe that
+$h_{0}(x)\geq\frac{3}{8}\quad\text{for every }x\in\left[\frac{3}{8},1\right].$
+Therefore,
+$\lim_{n\to\infty}h_{0}^{n}(1)\geq\frac{3}{8}.$
+Since $\displaystyle\inf\mathcal{D}_{h}\setminus\{0\}=\lim_{n\to\infty}h_{0}^{n}(1)$, it follows that $\mathcal{D}_{h}\cap(0,\frac{3}{8})=\emptyset$.
+Consequently, $\mathcal{D}_{h}$ is not dense in $[0,1]$.
+Finally, we remark that $h_{0}$ has a fixed point other than zero; indeed, $\displaystyle h_{0}\!\left(\frac{3}{8}\right)=\frac{3}{8}$.
+It is natural to consider the inverse function of the solution $\varphi$. We easily see the following assertion.
 Lemma 2.5.
 Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are D-systems on $[0,1]$.
-Let $\varphi$ be a unique solution of (1.1).
+Let $\varphi$ be the unique solution of (1.1).
 Then, the inverse function $\varphi^{-1}$ of $\varphi$ is well-defined and satisfies the following conjugate equation
 $\varphi^{-1}\circ g_{i}=f_{i}\circ\varphi^{-1},\ i\in\mathcal{I}_{N},$
-which is equivalent with (1.1) where $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are exchanged.
-By [4, Proposition 3.1], if $\varphi$ is singular, then, $\varphi^{-1}$ is also singular.
+which is equivalent with (1.1) where $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are exchanged.
+By [6, Proposition 3.1], if $\varphi$ is singular, then $\varphi^{-1}$ is also singular.
 2.2. Weak contraction
-Hereafter, we assume more additional conditions for the D-systems $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$. We say that a function $h:[0,1]\to[0,1]$ is a contraction on $[0,1]$ if
-$\textup{Lip}(h):=\sup_{x\neq y}\left|\frac{h(x)-h(y)}{x-y}\right|<1.$
-$\textup{Lip}(h)$ is called the Lipschitz constant of $h$. We say that a function $h:[0,1]\to[0,1]$ is a weak contraction on $[0,1]$ if
-$\phi_{h}(t):=\lim_{s\to t+0}\sup\left\{|h(x)-h(y)|:|x-y|\leq s\right\}<t,\ \ t>0.$
-The following is well-known. See [5, 14].
+Hereafter, we assume additional conditions for the D-systems $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$.
+We say that a function $h\colon[0,1]\to[0,1]$ is a contraction on $[0,1]$ if
+$\|h\|_{\textup{Lip}}\coloneqq\sup_{x\neq y}\left|\frac{h(x)-h(y)}{x-y}\right|<1.$
+$\|h\|_{\textup{Lip}}$ is called the Lipschitz constant of $h$.
+For a function $h\colon[0,1]\to\mathbb{R}$, let
+(2.5) $\omega_{h}(s)\coloneqq\sup\left\{|h(x)-h(y)|\colon|x-y|\leq s\right\}.$
+We say that a function $h\colon[0,1]\to[0,1]$ is a weak contraction on $[0,1]$ if
+$\phi_{h}(t)\coloneqq\lim_{s\to t+0}\omega_{h}(s)<t,\ \ t>0.$
 Lemma 2.6.
-Let $h$ be a weak contraction. Then, $\lim_{n\to\infty}\phi_{h}^{n}(t)=0$ for every $t>0$ and $|h(x)-h(y)|\leq\phi_{h}(|x-y|)$ for $x,y\in[0,1]$.
+Let $h$ be a weak contraction. Then, $\displaystyle\lim_{n\to\infty}\phi_{h}^{n}(t)=0$ for every $t>0$ and $|h(x)-h(y)|\leq\phi_{h}(|x-y|)$ for $x,y\in[0,1]$.
 Proof.
-Let
-$\omega_{h}(s):=\sup\left\{|h(x)-h(y)|:|x-y|\leq s\right\}.$
-Since this function is increasing, $\phi_{h}$ is right-continuous and increasing on $[0,1]$. Hence, $|h(x)-h(y)|\leq\lim_{s\to|x-y|+0}\phi_{h}(s)=\phi_{h}(|x-y|)$.
+Since $\omega_{h}$ is increasing, $\phi_{h}$ is increasing and right-continuous on $[0,1]$. Hence,
+$|h(x)-h(y)|\leq\lim_{s\to|x-y|+0}\phi_{h}(s)=\phi_{h}(|x-y|).$
 By the assumption, for every $t>0$, $(\phi_{h}^{n}(t))_{n}$ is strictly decreasing.
-There exists the limit $\phi^{\infty}_{h}(t):=\lim_{n\to\infty}\phi_{h}^{n}(t)$.
-By the right-continuity of $\phi_{h}$, $\phi_{h}\left(\phi^{\infty}_{h}(t)\right)=\phi^{\infty}_{h}(t)$.
+Hence, the limit $\displaystyle\phi^{\infty}_{h}(t)\coloneqq\lim_{n\to\infty}\phi_{h}^{n}(t)$ exists.
+By the right-continuity of $\phi_{h}$, we see that $\displaystyle\phi_{h}\left(\phi^{\infty}_{h}(t)\right)=\phi^{\infty}_{h}(t)$.
 Hence, $\phi^{\infty}_{h}(t)=0$.
@@
 Lemma 2.7.
-Let $\{h_{i}\}_{i}$ be a compatible system on $[0,1]$ such that each $h_{i}$ is a weak contraction on $[0,1]$. Then, $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system.
-We say that a compatible system $\{h_{i}\}_{i}$ on $[0,1]$ is an (weakly) contractive system on $[0,1]$ if each $h_{i}$ is an (weak) contraction on $[0,1]$.
-Every weakly contractive system is a D-system.
+Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be a compatible system on $[0,1]$ such that each $h_{i}$ is a weak contraction on $[0,1]$. Then, $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is a D-system.
 Proof.
-Let $\phi:=\max_{i\in\mathcal{I}_{N}}\phi_{h_{i}}$.
-Then, $\phi$ is right-continuous [^1] The assumption that $N$ is finite would be important.
-on $[0,1]$, $\phi(t)<t$ for every $t>0$, and $\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\phi(|x-y|)$, $x,y\in[0,1]$.
-Then, we see that $\lim_{n\to\infty}\phi^{n}(t)=0$ for every $t>0$.
+Let $\displaystyle\phi\coloneqq\max_{i\in\mathcal{I}_{N}}\phi_{h_{i}}$.
+Since each $\phi_{h_{i}}$ is right-continuous, $\phi$ is right-continuous [^1] The assumption that $N$ is finite is important.
+on $[0,1]$, $\phi(t)<t$ for every $t>0$, and $\displaystyle\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\phi(|x-y|)$, $x,y\in[0,1]$.
+Then, we see that $\displaystyle\lim_{n\to\infty}\phi^{n}(t)=0$ for every $t>0$.
 Hence,
 $\left|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right|\leq\phi^{n}(1).$
-$\lim_{n\to\infty}\max_{i_{1},\cdots,i_{n}\in\mathcal{I}_{N}}|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)|=0.$
+$\lim_{n\to\infty}\max_{i_{1},\ldots,i_{n}\in\mathcal{I}_{N}}\left|h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)-h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0)\right|=0.$
 By the definition, for each $n\geq 1$,
-$[0,1]=\bigcup_{i_{1},\cdots,i_{n}\in\mathcal{I}_{N}}\left[h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0),h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right],$
+$0,1]=\bigcup_{i_{1},\ldots,i_{n}\in\mathcal{I}_{N}}\left[h_{i_{1}}\circ\cdots\circ h_{i_{n}}(0),h_{i_{1}}\circ\cdots\circ h_{i_{n}}(1)\right],$
 and their interiors are disjoint. ∎
+We say that a compatible system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ is a (weakly) contractive system on $[0,1]$ if each $h_{i}$ is a (weak) contraction on $[0,1]$.
+Every weakly contractive system is a D-system.
 2.3. Linear fractional transforms
 Definition 2.8.
-We say that a family of functions $\{h_{i}\}_{i}$ is an LF system on $[0,1]$ if $h_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, for each $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers such that
-(1) (well-defined) $c_{i}+d_{i}>0$, $d_{i}>0$, $i\in\mathcal{I}_{N}$.
-(2) (strictly increasing, weak contraction)
+We say that a family of functions $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ is an LF system on $[0,1]$ if $h_{i}(x)\coloneqq\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, for each $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers such that
+(1) (well-defined) It holds that $c_{i}+d_{i}>0$ and $d_{i}>0$ for $i\in\mathcal{I}_{N}$.
+(2) (strictly increasing, weak contraction) It holds that
 $0<a_{i}d_{i}-b_{i}c_{i}\leq\min\{d_{i}^{2},(c_{i}+d_{i})^{2}\},\ i\in\mathcal{I}_{N}.$
-(3) (compatibility) $b_{0}=0$, $\dfrac{a_{i}+b_{i}}{c_{i}+d_{i}}=\dfrac{b_{i+1}}{d_{i+1}}$, $i\in\mathcal{I}_{N}$, where we let $\dfrac{b_{N}}{d_{N}}:=1$.
-For condition (2), we may consider a stronger condition that
+(3) (compatibility) It holds that $b_{0}=0$ and $\dfrac{a_{i}+b_{i}}{c_{i}+d_{i}}=\dfrac{b_{i+1}}{d_{i+1}}$ for $i\in\mathcal{I}_{N}$, where we let $\dfrac{b_{N}}{d_{N}}\coloneqq 1$.
+(4) The sequence $\left(\dfrac{b_{i}}{d_{i}}\right)_{i}$ is strictly increasing in $i$.
+Condition (4) follows from the conditions (1), (2) and (3) but is frequently used. For condition (2), we may consider a stronger condition:
 (2’) $0<a_{i}d_{i}-b_{i}c_{i}<\min\{d_{i}^{2},(d_{i}+c_{i})^{2}\},\ i\in\mathcal{I}_{N}$.
-Since $h_{i}$ is invariant if we multiply a positive constant by each of $a_{i},b_{i},c_{i}$ and $d_{i}$, we may assume that $d_{i}=1$, or, $a_{i}d_{i}-b_{i}c_{i}=1$.
+Since $h_{i}$ is invariant if we multiply each of $a_{i},b_{i},c_{i}$ and $d_{i}$ by a positive constant, we may assume that $d_{i}=1$ or $a_{i}d_{i}-b_{i}c_{i}=1$.
+An LF system is a compatible system.
 Lemma 2.9.
-Every LF system is a weakly contractive system. If (2’) holds instead of (2), then, it is a contractive system.
+Every LF system is a weakly contractive system. If (2’) holds instead of (2), then it is a contractive system.
 Proof.
-Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF-system. We assume that $d_{i}=1$ for every $i$. By Lemma 2.6, it suffices to show that each $h_{i}$ is a weak contraction. We first remark that
-$|h_{i}(x)-h_{i}(y)|=\frac{a_{i}-b_{i}c_{i}}{(c_{i}x+1)(c_{i}y+1)}|x-y|.$
+Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF-system.
+We assume that $d_{i}=1$ for every $i\in\mathcal{I}_{N}$.
+By Lemma 2.7, it suffices to show that each $h_{i}$ is a weak contraction.
+We first remark that
+(2.6) $|h_{i}(x)-h_{i}(y)|=\frac{a_{i}-b_{i}c_{i}}{(c_{i}x+1)(c_{i}y+1)}|x-y|.$
 We consider three cases: (i) $c_{i}>0$, (ii) $c_{i}<0$, and (iii) $c_{i}=0$.
-(i) Assume $c_{i}>0$. Then, $a_{i}-b_{i}c_{i}\leq 1$ and
+(i) Assume $c_{i}>0$.
+We remark that $a_{i}-b_{i}c_{i}\leq 1$ and
 $|h_{i}(x)-h_{i}(y)|\leq\frac{a_{i}-b_{i}c_{i}}{c_{i}|x-y|+1}|x-y|\leq\frac{|x-y|}{c_{i}|x-y|+1}.$
-Hence,
+Hence, by (2.5),
 $\omega_{h_{i}}(s)\leq\frac{s}{c_{i}s+1},\ \ s\geq 0.$
@@
 $|h_{i}(x)-h_{i}(y)|\leq\frac{a_{i}-b_{i}c_{i}}{(c_{i}+1)(c_{i}(1-|x-y|)+1)}|x-y|\leq\frac{(1+c_{i})|x-y|}{-c_{i}|x-y|+1+c_{i}}.$
-Hence,
+Hence, by (2.5),
 $\omega_{h_{i}}(s)\leq\frac{(1+c_{i})s}{-c_{i}s+1+c_{i}},\ \ s\geq 0.$
@@
 (iii) Assume $c_{i}=0$.
-Then, $0<a_{i}\leq 1$ and $|h_{i}(x)-h_{i}(y)|=a_{i}|x-y|$.
-By the assumption for $a_{i},b_{i},c_{i}$, $h_{j}(1)>h_{j}(0)$ for every $j$.
-Hence, by $N\geq 2$, $a_{i}=h_{i}(1)-h_{i}(0)<1$.
+We remark that $0<a_{i}\leq 1$ and $|h_{i}(x)-h_{i}(y)|=a_{i}|x-y|$.
+By the assumption for $a_{i},b_{i},c_{i}$, it holds that $h_{j}(1)>h_{j}(0)$ for every $j$.
+By the fact that $N\geq 2$, we obtain that $a_{i}=h_{i}(1)-h_{i}(0)<1$.
 ∎
@@
 The technique used above is related to a characterization of Rakotch contractions. See [20, Lemma 2.10 (iii)].
-Let $(\mathcal{I}_{N})^{*}:=\cup_{n\in\mathbb{N}}(\mathcal{I}_{N})^{n}$.
+Let $\displaystyle(\mathcal{I}_{N})^{*}\coloneqq\bigcup_{n\in\mathbb{N}}(\mathcal{I}_{N})^{n}$.
+The length of $\sigma$ is the unique number $n\in\mathbb{N}$ such that $\sigma\in(\mathcal{I}_{N})^{n}$.
 We denote the length of $\sigma\in(\mathcal{I}_{N})^{*}$ by $|\sigma|$.
-For an LF system $\{h_{i}\}_{i}$ and $\sigma=(\sigma_{1},\cdots,\sigma_{n})\in(\mathcal{I}_{N})^{*}$, we let
-$h_{\sigma}(x):=h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
+For an LF system $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ and $\sigma=(\sigma_{1},\ldots,\sigma_{n})\in(\mathcal{I}_{N})^{*}$, we let
+$h_{\sigma}(x)\coloneqq h_{\sigma_{1}}\circ\cdots\circ h_{\sigma_{n}}(x),\ x\in[0,1].$
 Lemma 2.11.
-Let $\{h_{i}\}_{i}$ be an LF system. Then, there exists $c\in(0,1)$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
+Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF system. Then, there exists $c\in(0,1)$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
 $h_{\sigma}(1)-h_{\sigma}(0)\geq c^{|\sigma|}.$
 Proof.
-Let $A_{i}:=\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}$, $i\in\mathcal{I}_{N}$.
+Let $A_{i}\coloneqq\begin{pmatrix}a_{i}&b_{i}\\ c_{i}&d_{i}\end{pmatrix}$, $i\in\mathcal{I}_{N}$.
 We assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i\in\mathcal{I}_{N}$.
-Let $(i_{n})_{n}\in(\mathcal{I}_{N})^{\mathbb{N}}$.
-Let
+Let $(i_{n})_{n}\in(\mathcal{I}_{N})^{\mathbb{N}}$ and
 $\begin{pmatrix}p_{n}&q_{n}\\ r_{n}&s_{n}\end{pmatrix}=A_{i_{1}}\cdots A_{i_{n}},\ n\geq 1.$
-Let $C_{1}:=\max\{|a_{i}|,|b_{i}|,|c_{i}|,|d_{i}|:i\in\mathcal{I}_{N}\}$. Then, for every $n$, $|r_{n+1}|+|s_{n+1}|\leq C_{1}(|r_{n}|+|s_{n}|)$. Hence, $|r_{n}|+|s_{n}|\leq 2C_{1}^{n}$ for every $n$.
-Assume that $\sigma=(i_{1},\cdots,i_{n})$. Then, $|\sigma|=n$ and
-$h_{\sigma}(1)-h_{\sigma}(0)=\frac{1}{s_{n}(r_{n}+s_{n})}\geq\frac{1}{(|r_{n}|+|s_{n}|)^{2}}\geq\frac{1}{4|C_{1}|^{2n}}.$
+Let $C_{1}\coloneqq\max\{|a_{i}|,|b_{i}|,|c_{i}|,|d_{i}|\colon i\in\mathcal{I}_{N}\}$.
+Then, for every $n$, $|r_{n+1}|+|s_{n+1}|\leq 2C_{1}(|r_{n}|+|s_{n}|)$.
+Hence, $|r_{n}|+|s_{n}|\leq(2C_{1})^{n}$ for every $n$.
+Assume that $\sigma=(i_{1},\ldots,i_{n})$. Then, $|\sigma|=n$ and by using the equality $p_{n}s_{n}-q_{n}r_{n}=1$,
+$h_{\sigma}(1)-h_{\sigma}(0)=\frac{p_{n}+q_{n}}{r_{n}+s_{n}}-\frac{q_{n}}{s_{n}}$
+$\frac{1}{s_{n}(r_{n}+s_{n})}\geq\frac{1}{(|r_{n}|+|s_{n}|)^{2}}\geq\frac{1}{4^{n}|C_{1}|^{2n}}.$
 We give an upper bound for $h_{\sigma}(1)-h_{\sigma}(0)$.
 Lemma 2.12.
-Let $\{h_{i}\}_{i}$ be an LF system. Then, there exists $C>0$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
+Let $\{h_{i}\}_{i\in\mathcal{I}_{N}}$ be an LF system. Then, there exists $C>0$ such that for every $\sigma\in(\mathcal{I}_{N})^{*}$,
 $h_{\sigma}(1)-h_{\sigma}(0)\leq\frac{C}{|\sigma|}.$
 Proof.
-We first remark that
-$(c_{i}x+d_{i})(c_{i}y+d_{i})\geq 1+|c_{i}||x-y|,\ x,y\in[0,1].$ (2.5)
-We assume that $x\leq y$. Assume $c_{i}>0$. Then,
-$c_{i}y+d_{i}\geq c_{i}x+d_{i}\geq\min_{x\in[0,1]}c_{i}x+d_{i}=\min\{d_{i},c_{i}+d_{i}\}=d_{i}\geq 1.$
+In this proof, we assume that $a_{i}d_{i}-b_{i}c_{i}=1$ for each $i\in\mathcal{I}_{N}$. We first show that
+(2.7) $(c_{i}x+d_{i})(c_{i}y+d_{i})\geq 1+|c_{i}||x-y|,\ x,y\in[0,1].$
+We assume that $x\leq y$. Assume $c_{i}>0$. Then, by Definition 2.8 (2),
+$c_{i}y+d_{i}\geq c_{i}x+d_{i}\geq\min_{z\in[0,1]}c_{i}z+d_{i}=d_{i}\geq 1.$
 Hence,
 $(c_{i}x+d_{i})(c_{i}y+d_{i})=(c_{i}x+d_{i})^{2}+c_{i}(y-x)(c_{i}x+d_{i})\geq 1+c_{i}(y-x).$
-Assume $c_{i}<0$. Then,
-$(c_{i}x+d_{i})(c_{i}y+d_{i})=(c_{i}y+d_{i})^{2}-c_{i}(y-x)(c_{i}x+d_{i})\geq 1-c_{i}(y-x).$
-If $c_{i}=0$, then, $(c_{i}x+d_{i})(c_{i}y+d_{i})=d_{i}^{2}\geq 1=1+|c_{i}||x-y|$. Thus we see (2.5).
-By (2.5),
+Assume $c_{i}<0$. Then, by Definition 2.8 (2) and (1),
+$c_{i}x+d_{i}\geq c_{i}y+d_{i}\geq\min_{z\in[0,1]}c_{i}z+d_{i}=c_{i}+d_{i}=|c_{i}+d_{i}|\geq 1.$
+Hence,
+$(c_{i}x+d_{i})(c_{i}y+d_{i})=(c_{i}y+d_{i})^{2}-c_{i}(y-x)(c_{i}y+d_{i})\geq 1-c_{i}(y-x).$
+If $c_{i}=0$, then $(c_{i}x+d_{i})(c_{i}y+d_{i})=d_{i}^{2}\geq 1=1+|c_{i}||x-y|$. Thus we see (2.7).
+By (2.6) and (2.7),
 $|h_{i}(x)-h_{i}(y)|\leq\frac{|x-y|}{1+|c_{i}||x-y|}.$
-If $c_{i}=0$, then,
+If $c_{i}=0$, then
 $|h_{i}(x)-h_{i}(y)|=\frac{a_{i}}{d_{i}}|x-y|.$
-Since $h_{i}(1)-h_{i}(0)<1$, $a_{i}<d_{i}$. Let $C_{1}:=\min\{|c_{i}|:c_{i}\neq 0\}$ and $C_{2}:=\max\left\{\dfrac{a_{i}}{d_{i}}:c_{i}=0\right\}$. Then, $C_{1}>0$, $C_{2}<1$ and
-$\max_{i}|h_{i}(x)-h_{i}(y)|\leq\max\left\{\frac{|x-y|}{1+C_{1}|x-y|},C_{2}|x-y|\right\}.$
+Since $h_{i}(1)-h_{i}(0)<1$, it holds that $a_{i}<d_{i}$.
+Let $C_{1}\coloneqq\min\{|c_{i}|\colon c_{i}\neq 0\}$ and $C_{2}\coloneqq\max\left\{\dfrac{a_{i}}{d_{i}}\colon c_{i}=0\right\}$.
+Then, $C_{1}>0$, $C_{2}<1$ and
+$\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\max\left\{\frac{|x-y|}{1+C_{1}|x-y|},C_{2}|x-y|\right\}.$
 We can take $C_{3}\in(0,C_{1})$ such that
-$\max_{i}|h_{i}(x)-h_{i}(y)|\leq\frac{|x-y|}{1+C_{3}|x-y|},\ x,y\in[0,1].$
-Let $a_{n}:=\max\left\{h_{\sigma}(1)-h_{\sigma}(0):\sigma\in(\mathcal{I}_{N})^{*},|\sigma|=n\right\}$. Then, $a_{1}\leq\dfrac{1}{1+C_{3}}$ and $a_{n+1}\leq\dfrac{a_{n}}{1+C_{3}a_{n}},n\geq 1$. Hence,
-$a_{n}\leq\frac{1}{nC_{3}+1}\leq\frac{1}{nC_{3}}.$
+$\max_{i\in\mathcal{I}_{N}}|h_{i}(x)-h_{i}(y)|\leq\frac{|x-y|}{1+C_{3}|x-y|},\ x,y\in[0,1].$
+Let $A_{n}\coloneqq\max\left\{h_{\sigma}(1)-h_{\sigma}(0)\colon\sigma\in(\mathcal{I}_{N})^{*},|\sigma|=n\right\}$.
+Then, $A_{1}\leq\dfrac{1}{1+C_{3}}$ and $A_{n+1}\leq\dfrac{A_{n}}{1+C_{3}A_{n}},n\geq 1$.
+Hence,
+$A_{n}\leq\frac{1}{nC_{3}+1}\leq\frac{1}{nC_{3}}.\qed$
 Proposition 2.13 (Weak version of regularity).
-Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are both LF systems and let $\varphi$ be the solution of (1.1).
+Assume that $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are LF systems and let $\varphi$ be the solution of (1.1).
 Then, there exists a constant $C>0$ such that for every $y\in(0,1)$ and $r>0$ satisfying that $0\leq y-r<y+r\leq 1$,
 $\mu_{\varphi}\left([y-r,y+r]\right)\geq\exp(-C/r).$
-This might be regarded as a weak version of the regularity of $\mu_{\varphi}$ in the sense of Ullman-Stahl-Totik. See [22, Criterion 1.1]. However, the following proof does not yield the regularity.
+This might be regarded as a weak version of the regularity of $\mu_{\varphi}$ in the sense of Ullman–Stahl–Totik. See [22, Criterion 1.1]. However, the following proof does not yield the regularity.
 Proof.
-We apply Lemma 2.10 to $\{g_{i}\}_{i}$ and Lemma 2.11 to $\{f_{i}\}_{i}$.
+We apply Lemma 2.11 to $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ and Lemma 2.12 to $\{f_{i}\}_{i\in\mathcal{I}_{N}}$.
 Let $C$ be the constant appearing in Lemma 2.12.
-Take $y\in(0,1)$ and $r>0$ such that $0\leq y-r<y+r\leq 1$ arbitralily.
-Let $n$ be a natural number such that $\frac{C}{n}<r\leq\frac{C}{n-1}$.
-Then, by the compatibility condition and Lemma 2.11, there exist $i_{1},\cdots,i_{n}\in\mathcal{I}_{N}$ such that
-$[f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)]\subset[y-r,y+r].$
+Take $y\in(0,1)$ and $r>0$ such that $0\leq y-r<y+r\leq 1$ arbitrarily.
+Let $n$ be a natural number such that $\dfrac{C}{n}<r\leq\dfrac{C}{n-1}$.
+We let $\dfrac{C}{n-1}\coloneqq\infty$ for $n=1$.
+Then, by the compatibility condition and Lemma 2.12, there exist $i_{1},\ldots,i_{n}\in\mathcal{I}_{N}$ such that
+$\left[f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)\right]\subset[y-r,y+r].$
 Hence, by (2.4),
-$\mu_{\varphi}([y-r,y+r])\geq g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0).$
-Let $c$ be the constant appearing in Lemma 2.10. By Lemma 2.10,
-$\mu_{\varphi}([y-r,y+r])\geq c^{n}\geq c^{1+C/r}.$
+$\mu_{\varphi}\left([y-r,y+r]\right)\geq\mu_{\varphi}\left(\left[f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0),f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1)\right]\right)$
+$\varphi(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(1))-\varphi(f_{i_{1}}\circ\cdots\circ f_{i_{n}}(0))$
+$g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0).$
+By Lemma 2.11, there exists a constant $c\in(0,1)$ depending only on $\{g_{i}\}_{i}$ such that
+$g_{i_{1}}\circ\cdots\circ g_{i_{n}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{n}}(0)\geq c^{n}.$
+By the assumption for $n$, we obtain that $c^{n}\geq c^{1+C/r}$. Thus, we obtain that
+$\mu_{\varphi}([y-r,y+r])\geq c^{1+C/r}.\qed$
 3. Singularity
-In this section, we consider (1.1) under the setting that $\{g_{i}\}_{i}$ is an LF-system on $[0,1]$, and, $\{f_{i}\}_{i}$ is a weakly contractive D-system on $[0,1]$.
-By Lemma 2.8, both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are D-systems.
+In this section, we consider (1.1) under the setting that $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ is an LF-system on $[0,1]$, and, $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ is a weakly contractive D-system on $[0,1]$.
+By Lemma 2.7 and Lemma 2.9, both $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are D-systems.
 Hence, by Lemma 2.3, there exists a unique continuous strictly increasing solution $\varphi$ of (1.1).
-We consider singularity for the solution $\varphi$.
+We consider the singularity for the solution $\varphi$.
 We will give sufficient conditions for $\dim_{H}\mu_{\varphi}<1$.
-It is known that if $\dim_{H}\mu_{\varphi}<1$, then, $\varphi$ is a singular function on $[0,1]$, specifically, $\varphi$ is a strictly increasing and continuous function on $[0,1]$ such that its derivative is equal to zero almost everywhere.
-The main purpose of this section is to extend [27, Theorem 43] to the case that $f_{i}$ can be non-linear.
+It is known that if $\dim_{H}\mu_{\varphi}<1$, then $\mu_{\varphi}$ is singular with respect to the Lebesgue measure.
+By [28, Definition 9.2.2 and Theorem 9.2.4], this implies that $\varphi$ is a singular function on $[0,1]$, specifically, $\varphi$ is a strictly increasing and continuous function on $[0,1]$ such that its derivative is equal to zero almost everywhere.
+The main purpose of this section is to extend [27, Theorem 43] to the case where $f_{i}$ can be non-linear.
 In [27], the author dealt with many objects other than the conjugate functional equation, so the arguments were not specialized for the functional equation.
 We introduce some new notations and the formulations of our main results and some of the arguments in the proofs are different from [27, Theorem 43].
-These changes will make arguments more transparent.
-We assume that an LF-system $\{g_{i}\}_{i}$ is defined by $g_{i}(x):=\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1]$, $i\in\mathcal{I}_{N}$, where $a_{i},b_{i},c_{i},d_{i},i\in\mathcal{I}_{N}$, are real numbers satisfying the conditions (1), (2) and (3) in Definition 2.7.
+The author believes that these changes will make the arguments more transparent.
+We assume that an LF-system $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ is defined by $g_{i}(x)\coloneqq\dfrac{a_{i}x+b_{i}}{c_{i}x+d_{i}},x\in[0,1],(i\in\mathcal{I}_{N})$, where $a_{i},b_{i},c_{i},d_{i},(i\in\mathcal{I}_{N})$, are real numbers satisfying the conditions (1), (2) and (3) in Definition 2.8, specifically,
+(1) It holds that $c_{i}+d_{i}>0$ and $d_{i}>0$ for $i\in\mathcal{I}_{N}$.
+(2) It holds that
+$0<a_{i}d_{i}-b_{i}c_{i}\leq\min\{d_{i}^{2},(c_{i}+d_{i})^{2}\},\ i\in\mathcal{I}_{N}.$
+(3) It holds that $b_{0}=0$ and $\dfrac{a_{i}+b_{i}}{c_{i}+d_{i}}=\dfrac{b_{i+1}}{d_{i+1}}$ for $i\in\mathcal{I}_{N}$, where we let $\dfrac{b_{N}}{d_{N}}\coloneqq 1$.
 In this section, we assume $d_{i}=1$ for each $i\in\mathcal{I}_{N}$.
+We show that $(1-a_{i})^{2}+4b_{i}c_{i}\geq 0$.
+If $c_{i}\geq 0$, then this is true since $b_{i}\geq 0$.
+If $c_{i}<0$, then by the compatibility condition, $a_{i}+b_{i}\leq c_{i}+1$.
+Hence, $0\leq b_{i}-c_{i}\leq 1-a_{i}$, and then $(1-a_{i})^{2}\geq(b_{i}-c_{i})^{2}$.
+Hence, $(1-a_{i})^{2}+4b_{i}c_{i}\geq(b_{i}+c_{i})^{2}\geq 0$.
 Let
-$\alpha:=\min\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\middle|1\leq i\leq N-1\right\},$
-where we delete [^2] This is different from [27]. $\frac{c_{0}}{1-a_{0}}$ if $a_{0}=1$. Let
-$\beta:=\max\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\middle|1\leq i\leq N-1\right\},$
-where $\dfrac{c_{0}}{1-a_{0}}:=+\infty$ if $a_{0}=1$.
-Let $Y:=[\alpha,\beta]$. Let
-$G_{i}(y):=\frac{(a_{i}-b_{i}c_{i})(y+1)}{(b_{i}y+1)((a_{i}+b_{i})y+(c_{i}+1))}=\frac{(b_{i+1}-b_{i})(y+1)}{(b_{i+1}y+1)(b_{i}y+1)},\ y\in[\alpha,\beta].$
+$\alpha\coloneqq\min\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\,\middle|\,1\leq i\leq N-1\right\},$
+where we delete [^2] This is different from [27]. $\dfrac{c_{0}}{1-a_{0}}$ if $a_{0}=1$. Let
+$\beta\coloneqq\max\left\{0,\frac{c_{0}}{1-a_{0}},\frac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\,\middle|\,1\leq i\leq N-1\right\},$
+where $\dfrac{c_{0}}{1-a_{0}}\coloneqq+\infty$ if $a_{0}=1$.
+Let $Y\coloneqq[\alpha,\beta]$ and
+$G_{i}(y)\coloneqq\frac{(a_{i}-b_{i}c_{i})(y+1)}{(b_{i}y+1)((a_{i}+b_{i})y+(c_{i}+1))}=\frac{(b_{i+1}-b_{i})(y+1)}{(b_{i+1}y+1)(b_{i}y+1)},\ y\in Y,$
+as well as
+$H_{i}(y)\coloneqq\frac{a_{i}y+c_{i}}{b_{i}y+1},\ y\in Y,$
+where we let $b_{N}\coloneqq 1$. We remark that $G_{0}(y)=\dfrac{b_{1}(y+1)}{b_{1}y+1}$ and $G_{N-1}(y)=\dfrac{1-b_{N-1}}{b_{N-1}y+1}$.
 Let
-$H_{i}(y):=\frac{a_{i}y+c_{i}}{b_{i}y+1},\ y\in[\alpha,\beta].$
-Let $G_{i}(+\infty)$ and $H_{i}(+\infty)$ be the limits. Specifically,
-$G_{0}(+\infty):=1,G_{i}(+\infty):=0,H_{0}(+\infty):=+\infty,H_{i}(+\infty):=\frac{a_{i}}{b_{i}},1\leq i\leq N-1.$
+$G_{0}(+\infty)\coloneqq 1,G_{i}(+\infty)\coloneqq 0,H_{0}(+\infty)\coloneqq+\infty,H_{i}(+\infty)\coloneqq\frac{a_{i}}{b_{i}},1\leq i\leq N-1.$
+Then,
+$G_{i}(+\infty)=\lim_{y\to+\infty}G_{i}(y),\ H_{i}(+\infty)=\lim_{y\to+\infty}H_{i}(y),\ i\in\mathcal{I}_{N}.$
 The following is the same as [27, Lemma 27], so we will not repeat the proof.
 Lemma 3.1.
+We have the following:
 (i) $0\in Y\subset[-1,+\infty]$.
@@
 (iii) $\alpha\leq H_{i}(\alpha)\leq H_{i}(\beta)\leq\beta$, that is, $H_{i}(Y)\subset Y$.
-(iv) $\sum_{i\in\Sigma_{N}}G_{i}(y)=1,\ y\in Y.$
+(iv) $\displaystyle\sum_{i\in\mathcal{I}_{N}}G_{i}(y)=1,\ y\in Y.$
 By (iii), we can compose $H_{i}$ freely on $Y$.
-We consider the compactification when $\beta=+\infty$.
-Then, $Y=[\alpha,\beta]$ is compact, and $G_{i}:Y\to[0,1]$ and $H_{i}:Y\to Y$ are continuous.
+We consider the Alexandroff compactification when $\beta=+\infty$.
+Then, $Y=[\alpha,\beta]$ is compact, and $G_{i}\colon Y\to[0,1]$ and $H_{i}\colon Y\to Y$ are continuous.
 We remark that $Y,G_{i},H_{i}$ are defined by $(g_{i})_{i}$, not by $(f_{i})_{i}$.
-$\textup{Fix}(H_{i})$ denotes the set of fixed points of $H_{i}$ on $[\alpha,\beta]$. This is uniquely determined and
+We denote the set of fixed points of $H_{i}$ on $Y$ by $\textup{Fix}(H_{i})$.
+The set $\textup{Fix}(H_{i})$ is a singleton and
 $\textup{Fix}(H_{0})=\left\{\dfrac{c_{0}}{1-a_{0}}\right\},\ \textup{Fix}(H_{i})=\left\{\dfrac{a_{i}-1+\sqrt{(1-a_{i})^{2}+4b_{i}c_{i}}}{2b_{i}}\right\},\ i\geq 1.$
+If $a_{0}=1$, then $c_{0}>0$ and hence $\textup{Fix}(H_{0})=\{+\infty\}$.
+If $a_{0}=1$ and $c_{0}=0$, then, by using the assumption that $d_{0}=1$, we obtain $g_{0}(x)=x$ and hence $N=1$.
+This contradicts the assumption that $N\geq 2$.
 Let
-$\mathcal{P}_{N}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}>0,\sum_{i}p_{i}=1\right\}$
-and
-$\overline{\mathcal{P}_{N}}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}\geq 0,\sum_{i}p_{i}=1\right\}.$
+$\displaystyle\mathcal{P}_{N}$ $\displaystyle\coloneqq\left\{(p_{i})_{i\in\mathcal{I}_{N}}\colon p_{i}>0,\sum_{i}p_{i}=1\right\}\ \textup{ and}$
+$\displaystyle\overline{\mathcal{P}_{N}}$ $\displaystyle\coloneqq\left\{(p_{i})_{i\in\mathcal{I}_{N}}\colon p_{i}\geq 0,\sum_{i}p_{i}=1\right\}.$
 Let
-${\bf G}(y):=(G_{0}(y),\dots,G_{N-1}(y))\in\overline{\mathcal{P}_{N}},\ y\in Y.$
-Let
-$\mathfrak{p}_{i}:={\bf G}\left(\textup{Fix}(H_{i})\right)\in\overline{\mathcal{P}_{N}},\ i\in\mathcal{I}_{N}.$
-Let
-$\dim_{H}\mu_{\varphi}:=\inf\left\{\dim_{H}A:A\in\mathcal{B}([0,1]),\mu_{\varphi}(A)=1\right\}.$
-The following results extend [27, Theorem 43 (i)] to the case that $\{f_{i}\}_{i}$ may be non-affine.
+$\displaystyle{\bf G}(y)$ $\displaystyle\coloneqq(G_{0}(y),\dots,G_{N-1}(y))\in\overline{\mathcal{P}_{N}},\ y\in Y,$
+$\displaystyle\mathfrak{p}_{i}$ $\displaystyle\coloneqq{\bf G}\left(\textup{Fix}(H_{i})\right)\in\overline{\mathcal{P}_{N}},\ \ \ \ \ \ \ \ \ \ \ \ \ \ i\in\mathcal{I}_{N},\ \ \ \ \ \ \ \ \ \textup{ and,}$
+$\displaystyle\dim_{H}\mu_{\varphi}$ $\displaystyle\coloneqq\inf\left\{\dim_{H}A\colon A\in\mathcal{B}([0,1]),\mu_{\varphi}(A)=1\right\},$
+where $\mathcal{B}([0,1])$ is the Borel $\sigma$-algebra on $[0,1]$. The following results extend [27, Theorem 43 (i)] to the case where $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ may be non-affine maps.
 Theorem 3.2.
-If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, where we let $r_{i}:=\textup{Lip}(f_{i})$.
-The probability vectors $\mathfrak{p}_{i},i\in\mathcal{I}_{N}$, are determined only by $(\alpha,\beta,\{G_{i}\}_{i},\{H_{i}\}i)$ and date back to $\{g_{i}\}_{i}$.
-$\{f_{i}\}_{i}$ is not related to this.
-This contains the affine case that $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$.
-The reason why we have introduced $Y,G_{i}$ and $H_{i}$ will be clarified in the proof.
-We also deal with the case that |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$. Let the total distance between probability vectors be
-$\|{\bf p}-{\bf q}\|:=\sum_{i\in\mathcal{I}_{N}}|p_{i}-q_{i}|,\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
+If $\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|\geq 2$, then for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ such that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$ and $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, where we let $r_{i}\coloneqq\|f_{i}\|_{\textup{Lip}}$.
+The probability vectors $\mathfrak{p}_{i},i\in\mathcal{I}_{N}$, are determined only by $(\alpha,\beta,\{G_{i}\}_{i},\{H_{i}\}_{i})$ and depend only on $\{g_{i}\}_{i\in\mathcal{I}_{N}}$.
+The system $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ is not related to this.
+This contains the case where $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$.
+We also deal with the case where $\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$. Let the total variation distance between probability vectors be
+$\left\|{\bf p}-{\bf q}\right\|_{1}\coloneqq\sum_{i\in\mathcal{I}_{N}}|p_{i}-q_{i}|,\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
 Theorem 3.3.
-If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that it holds that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $(f_{i})_{i}$ on $[0,1]$ such that $\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, and $\|{\bf r}-\mathfrak{p}_{0}\|\geq\epsilon$, where we let $r_{i}:=\textup{Lip}(f_{i})$ and ${\bf r}:=(r_{0},\cdots,r_{N-1})$.
-If |$\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then, $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$.
-If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$, and $\mathfrak{p}_{0}={\bf q}$, then, the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$.
+If $\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$, then for every $\epsilon\in(0,1/2)$, there exists $\delta>0$ such that $\dim_{H}\mu_{\varphi}<1$ for every contractive system $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ on $[0,1]$ such that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}<1+\delta$, $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$, and $\|{\bf r}-\mathfrak{p}_{0}\|_{1}\geq\epsilon$, where we let $r_{i}\coloneqq\|f_{i}\|_{\textup{Lip}}$ and ${\bf r}\coloneqq(r_{0},\ldots,r_{N-1})$.
+We remark that $\left|\{\mathfrak{p}_{i}\}_{i\in\mathcal{I}_{N}}\right|=1$ if and only if $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$.
+If $f_{i}(x)=a_{i}x+(a_{0}+\cdots+a_{i-1})$ for each $i\in\mathcal{I}_{N}$, and $\mathfrak{p}_{0}={\bf q}$, then the solution $\varphi$ is smooth on $[0,1]$ and we also have an explicit expression for $\varphi$.
 See [27, Theorem 43 (ii)].
 3.1. Proofs
-We first give notations for symbolic dynamics. Let :=$\mathbb{N}:=\{1,2,\cdots\}$. We define a map $\pi:(\mathcal{I}_{N})^{\mathbb{N}}\to[0,1]$ by
-$\pi(i_{1}i_{2}\cdots):=\lim_{k\to\infty}f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0).$
-This is well-defined because each $f_{i}$ is a contraction and $\left(f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0)\right)_{k}$ is a Cauchy sequence.
-Let $\xi_{j}:(\mathcal{I}_{N})^{\mathbb{N}}\to\mathcal{I}_{N}$ be the projection to the $j$-th coordinate. Let
-[…]
+We first give notations for symbolic dynamics. Let $\mathbb{N}\coloneqq\{1,2,\ldots\}$. We define a map $\pi\colon(\mathcal{I}_{N})^{\mathbb{N}}\to[0,1]$ by
+$\pi(i_{1}i_{2}\ldots)\coloneqq\lim_{k\to\infty}f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0).$
+This is well-defined because each $f_{i}$ is a contraction and $\left(f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0)\right)_{k}$ is a Cauchy sequence since
+$\left|f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}\circ f_{i_{k+1}}(0)-f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0)\right|\leq\left(\max_{i\in\mathcal{I}_{N}}\|f_{i}\|_{\textup{Lip}}\right)^{k}.$
+Let $\xi_{j}\colon(\mathcal{I}_{N})^{\mathbb{N}}\to\mathcal{I}_{N}$ be the projection to the $j$-th coordinate. Let
+$I(i_{1},\ldots,i_{n})\coloneqq\left\{w\in(\mathcal{I}_{N})^{\mathbb{N}}\colon\xi_{j}(w)=i_{j},\ 1\leq j\leq n\right\}.$
 Then, by Lemma 2.1,
-$\pi\left(I(i_{1},\cdots,i_{k})\right)=\left[f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0),f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(1)\right].$
+$\pi\left(I(i_{1},\ldots,i_{k})\right)=\left[f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(0),f_{i_{1}}\circ f_{i_{2}}\circ\cdots\circ f_{i_{k}}(1)\right].$
 Hence, by (2.4),
-$\mu_{\varphi}\left(\pi(I(i_{1},\cdots,i_{k}))\right)$ $=\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(1)\right)-\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(0)\right)$
-$=g_{i_{1}}\circ\cdots\circ g_{i_{k}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{k}}(0).$
+$\displaystyle\mu_{\varphi}\left(\pi(I(i_{1},\ldots,i_{k}))\right)$ $\displaystyle=\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(1)\right)-\varphi\left(f_{i_{1}}\circ\cdots\circ f_{i_{k}}(0)\right)$
+$\displaystyle=g_{i_{1}}\circ\cdots\circ g_{i_{k}}(1)-g_{i_{1}}\circ\cdots\circ g_{i_{k}}(0).$
 For $y\in Y$, let $\nu_{y}$ be a probability measure on $(\mathcal{I}_{N})^{\mathbb{N}}$ such that $\nu_{y}(I(i))=G_{i}(y)$, $i\in\mathcal{I}_{N}$, and,
-$\frac{\nu_{y}(I(i_{1},\cdots,i_{n-1},i_{n}))}{\nu_{y}(I(i_{1},\cdots,i_{n-1}))}=G_{i_{n}}\circ H_{i_{n-1}}\circ\cdots\circ H_{i_{1}}(y),\ n\geq 2.$ (3.1)
-Since $\sum_{i\in\mathcal{I}_{N}}G_{i}(y)=1$ for each $y\in Y$, the consistency condition holds.
-Therefore, a probability measure on $(\mathcal{I}_{N})^{\mathbb{N}}$ satisfying (3.1) uniquely exists by the Kolmogorov extension theorem.
-For $i\in\mathcal{I}_{N}$, let $F_{i}:(\mathcal{I}_{N})^{\mathbb{N}}\to(\mathcal{I}_{N})^{\mathbb{N}}$ be a map defined by $F_{i}(z)=iz$. Now we see that
-$\nu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\nu_{H_{i}(y)}\circ F_{i}^{-1}.$ (3.2)
-Let $\mu_{y}:=\nu_{y}\circ\pi^{-1}$. Then, $\mu_{y}$ is a Borel probability measure on $[0,1]$. By $\pi\circ F_{i}=f_{i}\circ\pi$ and (3.2),
-$\mu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\mu_{H_{i}(y)}\circ f_{i}^{-1}.$ (3.3)
+(3.1) $\frac{\nu_{y}(I(i_{1},\ldots,i_{n-1},i_{n}))}{\nu_{y}(I(i_{1},\ldots,i_{n-1}))}=G_{i_{n}}\circ H_{i_{n-1}}\circ\cdots\circ H_{i_{1}}(y),\ n\geq 2.$
+Since $\displaystyle\sum_{i\in\mathcal{I}_{N}}G_{i}(y)=1$ for each $y\in Y$, the Kolmogorov consistency condition holds.
+Therefore, a probability measure on $(\mathcal{I}_{N})^{\mathbb{N}}$ satisfying (3.1) exists uniquely by the Kolmogorov extension theorem [32, Theorem 2.4.3].
+For $i\in\mathcal{I}_{N}$, let $F_{i}\colon(\mathcal{I}_{N})^{\mathbb{N}}\to(\mathcal{I}_{N})^{\mathbb{N}}$ be a map defined by $F_{i}(z)=iz$. Now we see that
+(3.2) $\nu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\nu_{H_{i}(y)}\circ F_{i}^{-1}.$
+Let $\mu_{y}\coloneqq\nu_{y}\circ\pi^{-1}$. Then, $\mu_{y}$ is a Borel probability measure on $[0,1]$. By $\pi\circ F_{i}=f_{i}\circ\pi$ and (3.2),
+(3.3) $\mu_{y}=\sum_{i\in\mathcal{I}_{N}}G_{i}(y)\mu_{H_{i}(y)}\circ f_{i}^{-1}.$
 We recall that $0\in Y=[\alpha,\beta]$. Then, by (3.3),
 $\mu_{0}=\sum_{i\in\mathcal{I}_{N}}G_{i}(0)\mu_{H_{i}(0)}\circ f_{i}^{-1}.$
-Since $H_{i}(0)\neq 0$ can occur, this does not define an iterated function system. (3.3) is in the framework of a place-dependent iterated function system [9].
+Since $H_{i}(0)\neq 0$ can occur, this does not define an iterated function system. The equation (3.3) is in the framework of a place-dependent iterated function system [11].
 By [27, Lemma 30],
-$\nu_{0}\circ\pi^{-1}=\mu_{0}=\mu_{\varphi}.$ (3.4)
-Let $\epsilon\in(0,1/2)$.
-We assume that $\epsilon\leq\min_{i\in\mathcal{I}_{N}}r_{i}\leq\max_{i\in\mathcal{I}_{N}}r_{i}\leq 1-\epsilon$.
-Since $\{f_{i}\}_{i}$ is a compatible system, $\sum_{i\in\mathcal{I}_{N}}r_{i}\geq 1$ and $r_{i}>0$ for each $i\in\mathcal{I}_{N}$.
-Let $s\geq 1$ is a constant such that $\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$.
+(3.4) $\nu_{0}\circ\pi^{-1}=\mu_{0}=\mu_{\varphi}.$
+Let $\epsilon\in(0,1/2)$. We assume that
+(3.5) $\epsilon\leq\min_{i\in\mathcal{I}_{N}}r_{i}\leq\max_{i\in\mathcal{I}_{N}}r_{i}\leq 1-\epsilon.$
+Since $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ is a compatible system, $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}\geq 1$ and $r_{i}>0$ for each $i\in\mathcal{I}_{N}$.
+Let $s\geq 1$ be the constant such that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$.
 As the following assertion shows, the constant $s$ is close to $1$ if $\delta>0$ is sufficiently small.
 Lemma 3.4.
-Let $\epsilon_{2}>0$. Then, for $\delta>0$ such that
-$\frac{\log N}{\log(N/(1+\delta))}<1+\frac{\epsilon_{2}}{\log 1/\epsilon},$ (3.5)
-it holds that $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$ for every choice of $(r_{i})_{i\in\mathcal{I}_{N}}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$.
+Let $\epsilon_{2}>0$. Then, for $\delta>0$ such that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}\leq 1+\delta$ and
+(3.6) $\frac{\log N}{\log(N/(1+\delta))}<1+\frac{\epsilon_{2}}{\log(1/\epsilon)},$
+it holds that
+$s\leq\frac{\log N}{\log(N/(1+\delta))}<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$
+for every choice of $(r_{i})_{i\in\mathcal{I}_{N}}\in\mathcal{P}_{N}$ such that $\epsilon\leq r_{i}\leq 1-\epsilon$ for each $i\in\mathcal{I}_{N}$.
 Proof.
-Let $u_{i}:=r_{i}/\sum_{i}r_{i}$. Then, $\sum_{i}u_{i}=1$. We see that $\sum_{i}r_{i}^{s}=1$ if and only if $\sum_{i}u_{i}^{s}=(\sum_{i}r_{i})^{-s}$. Hence, if $\sum_{i}r_{i}\leq 1+\delta$, then,
-$\frac{\log(\sum_{i}u_{i}^{s})}{s}\leq-\log(1+\delta).$
-$\sum_{i}u_{i}^{s}$ takes its minimum at $u_{0}=\cdots=u_{N-1}=\dfrac{1}{N}$.
+Let $\displaystyle u_{i}\coloneqq\frac{r_{i}}{\sum_{i\in\mathcal{I}_{N}}r_{i}}$.
+Then, $\displaystyle\sum_{i\in\mathcal{I}_{N}}u_{i}=1$.
+We see that $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}^{s}=1$ if and only if $\displaystyle\sum_{i\in\mathcal{I}_{N}}u_{i}^{s}=\left(\sum_{i}r_{i}\right)^{-s}$.
+Since $\displaystyle\sum_{i\in\mathcal{I}_{N}}r_{i}\leq 1+\delta$,
+$\frac{\log\left(\sum_{i\in\mathcal{I}_{N}}u_{i}^{s}\right)}{s}\geq-\log(1+\delta).$
+The sum $\displaystyle\sum_{i\in\mathcal{I}_{N}}u_{i}^{s}$ takes its minimum at $u_{0}=\cdots=u_{N-1}=\dfrac{1}{N}$.
 We see this by the method of Lagrange multipliers.
-The minimum is $N^{1-s}\leq 1$.
-Hence, $-\dfrac{s-1}{s}\log N\geq-\log(\sum_{i}r_{i})$, which is equivalent to $s<\dfrac{\log N}{\log(N/(1+\delta))}$.
-Therefore, if (3.5) holds, then, $s<1+\dfrac{\epsilon_{2}}{\log 1/\epsilon}$.
+The minimum of the sum is $N^{1-s}\leq 1$.
+Hence, $\displaystyle-\frac{s-1}{s}\log N\geq-\log\left(\sum_{i\in\mathcal{I}_{N}}r_{i}\right)$, which is equivalent to $s\leq\dfrac{\log N}{\log(N/(1+\delta))}$.
+Therefore, if (3.6) holds, then $\displaystyle s\leq\frac{\log N}{\log(N/(1+\delta))}<1+\frac{\epsilon_{2}}{\log(1/\epsilon)}$.
 ∎
 Let
-$R_{n}({\bf q};z):=\frac{\nu_{0}(I_{n}(z))}{q_{\xi_{1}(z)}\cdots q_{\xi_{n}(z)}},\ \ {\bf q}=(q_{i})_{i}\in\mathcal{P}_{N},z\in(\mathcal{I}_{N})^{\mathbb{N}}.$
-By a measure-theoretic argument, we can show that
+$R_{n}\left({\bf q};z\right)\coloneqq\frac{\nu_{0}\left(I(\xi_{1}(z),\ldots,\xi_{n}(z))\right)}{q_{\xi_{1}(z)}\cdots q_{\xi_{n}(z)}},\ \ {\bf q}=(q_{i})_{i}\in\mathcal{P}_{N},z\in(\mathcal{I}_{N})^{\mathbb{N}}.$
+By a measure-theoretic argument, we can show the following assertion:
 Lemma 3.5.
-Let $\widetilde{q}_{i}:=r_{i}^{s}>0$ and ${\bf\widetilde{q}}:=(\widetilde{q}_{i})_{i}\in\mathcal{P}_{N}$. If there exists $\epsilon_{2}>0$ such that
-$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf\widetilde{q}};z)}{n}\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$ (3.6)
-then,
+Let $\widetilde{q}_{i}\coloneqq r_{i}^{s}>0$ and ${\bf\widetilde{q}}\coloneqq(\widetilde{q}_{i})_{i}\in\mathcal{P}_{N}$. If there exists $\epsilon_{2}>0$ such that
+(3.7) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf\widetilde{q}};z)}{n}\leq-\epsilon_{2},\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
+then
 $\dim_{H}\mu_{\varphi}\leq s-\frac{\epsilon_{2}}{\log 1/\epsilon}.$
-We omit the proof of this assertion. See the proof of [27, Lemma 7]. (3.4) plays an important role.
-By Lemmas 3.4 and 3.5, we obtain that
+We omit the proof of this assertion. See the proof of [27, Lemma 7]. The equation (3.4) plays an important role.
+By Lemmas 3.4 and 3.5, we obtain the following assertion:
 Proposition 3.6.
 Assume that there exists $\epsilon_{2}>0$ such that for every ${\bf q}\in\mathcal{P}_{N}$,
-$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$ (3.7)
-Then, for $\delta=\delta(N,\epsilon,\epsilon_{2})>0$ satisfying (3.5), $\dim_{H}\mu_{\varphi}<1$.
+(3.8) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-\epsilon_{2},\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
+Then, for $\delta=\delta(N,\epsilon,\epsilon_{2})>0$ satisfying (3.6), it holds that $\dim_{H}\mu_{\varphi}<1$.
 Now it suffices to establish (3.8). We give some definitions for the relative entropy. Let
-$s_{N}\left({\bf p}|{\bf q}\right):=\sum_{i\in\mathcal{I}_{N}}p_{i}\log\frac{q_{i}}{p_{i}},\ \ {\bf p}=(p_{i})_{i},{\bf q}=(q_{i})_{i}\in\overline{\mathcal{P}_{N}}.$
-If $p_{i}=0$, then, $p_{i}\log\frac{q_{i}}{p_{i}}:=0$.
-$s_{N}$ is continuous on $\overline{\mathcal{P}_{N}}$.
-Then, by Jensen’s inequality, $s_{N}({\bf p}|{\bf q})\leq 0$.
+$s_{N}\left({\bf p}|{\bf q}\right)\coloneqq\sum_{i\in\mathcal{I}_{N}}p_{i}\log\frac{q_{i}}{p_{i}},\ \ {\bf p}=(p_{i})_{i\in\mathcal{I}_{N}},{\bf q}=(q_{i})_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}}.$
+If $p_{i}=0$, then we let $p_{i}\log\dfrac{q_{i}}{p_{i}}\coloneqq 0$.
+This is also valid for $q_{i}=0$.
+For each ${\bf q}\in\mathcal{P}_{N}$, the function $s_{N}(\cdot|{\bf q})$ is continuous on $\overline{\mathcal{P}_{N}}$.
+If ${\bf q}\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$, $s_{N}$ can take -$\infty$.
+By Jensen’s inequality, $s_{N}\left({\bf p}|{\bf q}\right)\leq 0$.
 It holds that $s_{N}\left({\bf p}|{\bf q}\right)=0$ if and only if ${\bf p}={\bf q}$.
 For $y\in Y,z\in(\mathcal{I}_{N})^{\mathbb{N}}$, let
-$p_{j}(y;z):=\left(G_{i}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(y)\right)_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}},\ j\geq 1.$
+$p_{j}(y;z)\coloneqq\left(G_{i}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(y)\right)_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}_{N}},\ j\geq 1.$
 and
-$p_{0}(y;z):={\bf G}(y)\in\overline{\mathcal{P}_{N}}.$
-By a martingale argument, we can show that
+$p_{0}(y;z)\coloneqq{\bf G}(y)\in\overline{\mathcal{P}_{N}}.$
+By a martingale argument, we can show the following assertion:
 Proposition 3.7.
 It holds that for every ${\bf q}\in\mathcal{P}_{N}$,
-$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}=\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}(p_{j}(0;z)|{\bf q}),\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
+$\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}=\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}(p_{j}(0;z)|{\bf q}),\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
 We omit the proof of this assertion. See the proof of [27, Proposition 6]. It is important to assume that ${\bf q}\in\mathcal{P}_{N}$, not ${\bf q}\in\overline{\mathcal{P}_{N}}$.
-For $\epsilon\in(0,1)$, let $\overline{\mathcal{P}_{N}(\epsilon)}:=\left\{(p_{i})_{i\in\mathcal{I}_{N}}:p_{i}\geq\epsilon,\sum_{i}p_{i}=1\right\}$.
+For $\epsilon\in(0,1)$, let
+$\overline{\mathcal{P}_{N}(\epsilon)}\coloneqq\left\{(p_{i})_{i\in\mathcal{I}_{N}}\colon p_{i}\geq\epsilon,\sum_{i\in\mathcal{I}_{N}}p_{i}=1\right\}.$
 Then, $\overline{\mathcal{P}_{N}(\epsilon)}\subset\mathcal{P}_{N}\subset\overline{\mathcal{P}_{N}}$.
-Recall Lemma 3.4. If
+Let $\widetilde{q}_{i}\coloneqq r_{i}^{s}>0$ and ${\bf\widetilde{q}}\coloneqq(\widetilde{q}_{i})_{i}\in\mathcal{P}_{N}$. Recall Lemma 3.4. If
 $\frac{\log N}{\log(N/(1+\delta))}\leq 1+\frac{\log 2}{\log(1/\epsilon)},$
-then, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. In other words, ${\bf\widetilde{q}}$ is apart from the boundary $\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$
-Assume that
+then, by (3.5), $\displaystyle r_{i}^{s}\geq\epsilon^{s}\geq\epsilon^{1+\frac{\log 2}{\log(1/\epsilon)}}=\dfrac{\epsilon}{2}$ and hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
+In other words, ${\bf\widetilde{q}}$ has positive distance from the boundary $\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
+Assume the following assertion:
 Proposition 3.8.
-If |$|\{\mathfrak{p}_{i}\}_{i}|\geq 2$, then, there exists $\epsilon_{2}>0$ such that for every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$,
-$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\epsilon_{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
+If $\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$, then there exists $\epsilon_{2}>0$ such that for every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$,
+$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\epsilon_{2},\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
 Theorem 3.2 follows from Propositions 3.6, 3.7 and 3.8.
@@
 We first give some further definitions for quantities of the relative entropy.
+We remark that the diameter of $\overline{\mathcal{P}}_{N}$ with respect to $\|\cdot\|_{1}$ is equal to $2$.
+By the triangle inequality, it holds that for ${\bf p}=(p_{i})_{i\in\mathcal{I}_{N}},{\bf q}=(q_{i})_{i\in\mathcal{I}_{N}}\in\overline{\mathcal{P}}_{N}$,
+$\left\|{\bf p}-{\bf q}\right\|_{1}=\sum_{i\in\mathcal{I}_{N}}|p_{i}-q_{i}|\leq\sum_{i\in\mathcal{I}_{N}}(p_{i}+q_{i})=2.$
+If ${\bf p}=(1,0,\dots,0)$ and ${\bf q}=(0,\dots,0,1)$, then $\left\|{\bf p}-{\bf q}\right\|_{1}=2$.
 Lemma 3.9.
-Let
-$V\left({\bf q};\epsilon\right):=\max\left\{s_{N}\left({\bf p}|{\bf q}\right)\middle|{\bf p}\in\overline{\mathcal{P}_{N}},\|{\bf p}-{\bf q}\|\geq\epsilon\right\},\ \epsilon>0.$
+For ${\bf q}\in\mathcal{P}_{N}$, let
+$V\left({\bf q};\epsilon^{\prime}\right)\coloneqq\max\left\{s_{N}\left({\bf p}|{\bf q}\right)\middle|{\bf p}\in\overline{\mathcal{P}_{N}},\|{\bf p}-{\bf q}\|_{1}\geq\epsilon^{\prime}\right\},\ \epsilon^{\prime}\in(0,1].$
 Then,
-(i) $V\left({\bf q};\epsilon\right)<0$.
-(ii) $V\left({\bf q};\epsilon\right)$ is decreasing with respect to $\epsilon$.
-(iii) $V\left({\bf q};\epsilon\right)$ is right-continuous with respect to $\epsilon$.
-(iv) $V\left({\bf q};\epsilon\right)$ is continuous with respect to ${\bf q}$.
+(i) $V\left({\bf q};\epsilon^{\prime}\right)<0$.
+(ii) $V\left({\bf q};\epsilon^{\prime}\right)$ is decreasing with respect to $\epsilon^{\prime}$.
+(iii) $V\left({\bf q};\epsilon^{\prime}\right)$ is right-continuous with respect to $\epsilon^{\prime}$.
+(iv) $V\left({\bf q};\epsilon^{\prime}\right)$ is continuous with respect to ${\bf q}$.
 Proof.
-(i) and (ii) are easy to see.
-(iii) Assume that for some $\epsilon_{0}\in(0,1/3)$, $V\left({\bf q};\epsilon\right)$ is not right-continuous.
-Then, $V\left({\bf q};\epsilon_{0}\right)>\lim_{\epsilon\to\epsilon_{0}+0}V\left({\bf q};\epsilon\right)$.
+Assertions (i) and (ii) are easy to see.
+(iii) Assume that for some $\epsilon_{0}\in(0,1/3)$, $V\left({\bf q};\epsilon^{\prime}\right)$ is not right-continuous.
+Then, $\displaystyle V\left({\bf q};\epsilon_{0}\right)>\lim_{\epsilon^{\prime}\to\epsilon_{0}+0}V\left({\bf q};\epsilon^{\prime}\right)$.
 Assume that $V\left({\bf q};\epsilon_{0}\right)=s_{N}\left({\bf p}|{\bf q}\right)$.
-If $\|{\bf p}-{\bf q}\|>\epsilon_{0}$, then, this cannot occur.
-Assume $\|{\bf p}-{\bf q}\|=\epsilon_{0}$.
-Then, there exists a sequence $\{{\bf p}^{(n)}\}_{n}$ in $\overline{\mathcal{P}_{N}}$ such that $\|{\bf p}^{(n)}-{\bf q}\|\geq\epsilon_{0}+N^{-n}$ for each $n$ and $\lim_{n\to\infty}\|{\bf p}^{(n)}-{\bf q}\|=\epsilon_{0}$.
+If $\left\|{\bf p}-{\bf q}\right\|_{1}>\epsilon_{0}$, then this cannot occur.
+Assume $\left\|{\bf p}-{\bf q}\right\|_{1}=\epsilon_{0}$.
+Then, there exists a sequence $\{{\bf p}^{(n)}\}_{n}$ in $\overline{\mathcal{P}_{N}}$ such that $\left\|{\bf p}^{(n)}-{\bf q}\right\|_{1}\geq\epsilon_{0}+N^{-n}$ for each $n$ and $\displaystyle\lim_{n\to\infty}\left\|{\bf p}^{(n)}-{\bf q}\right\|_{1}=\epsilon_{0}$.
 Hence,
@@
 (iv) follows from (iii) and
-$V\left({\bf q};\epsilon+\|{\bf p}-{\bf q}\|\right)\leq V\left({\bf p};\epsilon\right)$
-By Lemma 3.9, we can let
-$V(\epsilon):=\max_{{\bf q}\in\overline{\mathcal{P}_{N}}}V\left({\bf q};\epsilon\right)<0.$
-We now give a quantity defined by $Y,\{G_{i}\}_{i}$ and $\{H_{i}\}_{i}$.
+$V\left({\bf q};\epsilon^{\prime}+\left\|{\bf p}-{\bf q}\right\|_{1}\right)\leq V\left({\bf p};\epsilon^{\prime}\right)\qed$
+Setting
+$V_{\epsilon,1}(\epsilon^{\prime})\coloneqq\max_{{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}}V\left({\bf q};\epsilon^{\prime}\right),\ \ \epsilon^{\prime}\in(0,1],$
+we obtain by Lemma 3.9 and the compactness of $\overline{\mathcal{P}_{N}(\epsilon/2)}$ that
+$V_{\epsilon,1}(\epsilon^{\prime})<0.$
 Lemma 3.10.
 Let
-$W_{i}\left({\bf q},y\right):=\max\left\{\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|,\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|\right\},\ {\bf q}\in\overline{\mathcal{P}_{N}},y\in Y,i\in\mathcal{I}_{N}.$
+$W_{i}\left({\bf q},y\right)\coloneqq\max\left\{\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|,\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|\right\},\ {\bf q}\in\overline{\mathcal{P}_{N}},y\in Y,i\in\mathcal{I}_{N}.$
 Then,
@@
 (ii) $W_{i}({\bf q},y)=0$ if and only if ${\bf q}=\mathfrak{p}_{i}$ and $y=\textup{Fix}(H_{i})$.
-This assertion is easily shown.
-If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then, $s_{N}\left({\bf G}(y)|{\bf q}\right)<0$.
+This assertion is easily shown and we omit the proof.
+If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|>0$, then $s_{N}\left({\bf G}(y)|{\bf q}\right)<0$.
 However, in our setting, we cannot exclude the possibility that $|G_{k}(y)-q_{k}|=0$ for every $k\in\mathcal{I}_{N}$.
-Assume that |$\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$.
-If $\max_{k\in\mathcal{I}_{N}}|G_{k}(y)-q_{k}|$ is close to 00, then, $\max_{k\in\mathcal{I}_{N}}|G_{k}(H_{i}(y))-q_{k}|$ is strictly positive.
+Assume that $\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$.
+If $\displaystyle\max_{k\in\mathcal{I}_{N}}\left|G_{k}(y)-q_{k}\right|$ is close to $0$, then $\displaystyle\max_{k\in\mathcal{I}_{N}}\left|G_{k}(H_{i}(y))-q_{k}\right|$ is strictly positive.
 Hence, $W_{i}\left({\bf q},y\right)>0$.
-The following corresponds to [27, Lemma 8]. Informally speaking, if $W_{i}({\bf q},y)$ is positive uniformly to $y$, then, a sum of two successive terms of $s_{N}$ is negative.
+Since the diameter of $\overline{\mathcal{P}}_{N}$ is equal to $2$, we obtain $0\leq W_{i}\left({\bf q},y\right)\leq 2$.
+The following corresponds to [27, Lemma 8]. Informally speaking, if $W_{i}({\bf q},y)$ is positive uniformly in $y$, then a sum of two successive terms of $s_{N}$ is negative.
 Lemma 3.11.
-Let ${\bf q}\in\mathcal{P}_{N}$. Let $i_{1}\in\mathcal{I}_{N}$. Assume that for some $\epsilon_{1}>0$,
-$\inf_{y\in Y}W_{i_{1}}\left({\bf q},y\right)\geq\epsilon_{1}.$ (3.8)
-Assume that $X_{j+1}(z)=i_{1}$. Then,
-$s_{N}\left(p_{j}(0;z)|{\bf q}\right)+s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)\leq V(\epsilon_{1})<0.$
+Let ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. Let $i_{1}\in\mathcal{I}_{N}$. Assume that for some $\epsilon_{1}\in(0,1]$,
+(3.9) $\inf_{y\in Y}W_{i_{1}}\left({\bf q},y\right)\geq\epsilon_{1}.$
+Assume that $z\in\left(\mathcal{I}_{N}\right)^{\mathbb{N}}$ and $\xi_{j+1}(z)=i_{1}$. Then,
+$s_{N}\left(p_{j}(0;z)|{\bf q}\right)+s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)\leq V_{\epsilon,1}(\epsilon_{1})<0.$
 Proof.
-The proof is essentially the same as the proof of [27, Lemma 8] [^3] It contains a typo.. Since $X_{j+1}(z)=i_{1}$,
-$p_{j}(0;z)={\bf G}\left(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0)\right),$
+The proof is essentially the same as the proof of [27, Lemma 8] [^3] The proof contains a typo.
+In Case 2 in the proof, $G_{i+l}(h_{i+l}(y;x))$ should be replaced with $G_{i_{l}}(h_{i+l}(y;x))$..
+Since $\xi_{j+1}(z)=i_{1}$,
+$p_{j}(0;z)={\bf G}\left(H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(0)\right),$
 and
-$p_{j+1}(0;z)={\bf G}\left(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0)\right).$
+$p_{j+1}(0;z)={\bf G}\left(H_{i_{1}}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(0)\right).$
 Let ${\bf q}=(q_{k})_{k}$. Then, by (3.9), either
-$\left|G_{k_{1}}(H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))-q_{k_{1}}\right|\geq\epsilon_{1},\textup{ for some }k_{1}\in\mathcal{I}_{N},$
-or,
-$\left|G_{k_{2}}(H_{i_{1}}\circ H_{X_{j}(z)}\circ\cdots\circ H_{X_{1}(z)}(0))-q_{k_{2}}\right|\geq\epsilon_{1},\textup{ for some }k_{2}\in\mathcal{I}_{N},$
+$\left|G_{k_{1}}(H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(0))-q_{k_{1}}\right|\geq\epsilon_{1},\textup{ for some }k_{1}\in\mathcal{I}_{N}$
+or
+$\left|G_{k_{2}}(H_{i_{1}}\circ H_{\xi_{j}(z)}\circ\cdots\circ H_{\xi_{1}(z)}(0))-q_{k_{2}}\right|\geq\epsilon_{1},\textup{ for some }k_{2}\in\mathcal{I}_{N},$
 holds.
-If the former holds, then, $s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$.
-If the latter holds, then, $s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$.
+We remark that at least one of them holds.
+If the former holds, then $s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$.
+If the latter holds, then $s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{1}\right)$.
 Now it suffices to recall that $s_{N}$ takes non-positive values.
 ∎
-Let :=|$\ell:=|\{\mathfrak{p}_{i}\}_{i}|\geq 2$.
-We assume that $\{\mathfrak{p}_{i}\}_{i}=\{\mathfrak{p}_{j_{k}}:1\leq k\leq\ell\}$ where $10\leq j_{1}<\cdots<j_{\ell}\leq N-$ 1.
-Let $\delta_{0}:=\min\{\|\mathfrak{p}_{j_{k_{1}}}-\mathfrak{p}_{j_{k_{2}}}\|:k_{1}\neq k_{2}\}>0$.
-Let $B_{j}:=\left\{{\bf p}\in\overline{\mathcal{P}_{N}}:\|{\bf p}-\mathfrak{p}_{j}\|\leq\delta_{0}/4\right\}$.
-$B_{j}=B_{k}$ can occur even if $j\neq k$.
-Define a coloring map $\mathscr{A}:\overline{\mathcal{P}_{N}}\to\mathcal{I}_{N}$ by
-$\mathscr{A}\left({\bf p}\right):=\begin{cases}j_{2}&{\bf p}\in B_{j_{1}}\\ \vdots&\\ j_{\ell}&{\bf p}\in B_{j_{\ell-1}}\\ j_{1}&{\bf p}\in B_{j_{\ell}}\\ 0&\textup{ otherwise}\end{cases}.$
-Under these definitions, we see that
+Let $\ell\coloneqq\left|\{\mathfrak{p}_{i}\}_{i}\right|\geq 2$.
+We assume that $\{\mathfrak{p}_{i}\}_{i}=\{\mathfrak{p}_{j_{k}}\colon 1\leq k\leq\ell\}$ where $0\leq j_{1}<\cdots<j_{\ell}\leq N-1$.
+Let
+$\displaystyle\delta_{0}$ $\displaystyle\coloneqq\min\left\{\left\|\mathfrak{p}_{j_{k_{1}}}-\mathfrak{p}_{j_{k_{2}}}\right\|_{1}\colon k_{1}\neq k_{2}\right\}>0,\ \ \textup{ and }$
+$\displaystyle B_{j}$ $\displaystyle\coloneqq\left\{{\bf p}\in\overline{\mathcal{P}_{N}}\colon\left\|{\bf p}-\mathfrak{p}_{j}\right\|_{1}\leq\delta_{0}/4\right\}.$
+The equality $B_{j}=B_{k}$ can occur even if $j\neq k$. Define a coloring map $\mathscr{A}\colon\overline{\mathcal{P}_{N}}\to\mathcal{I}_{N}$ by
+$\mathscr{A}\left({\bf p}\right)\coloneqq\begin{cases}j_{2}&{\bf p}\in B_{j_{1}}\\ \vdots&\\ j_{\ell}&{\bf p}\in B_{j_{\ell-1}}\\ j_{1}&{\bf p}\in B_{j_{\ell}}\\ 0&\textup{ otherwise}\end{cases}.$
+Under these definitions, we see the following assertion:
 Lemma 3.12.
-$\widetilde{\epsilon_{1}}:=\inf\left\{W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\middle|{\bf q}\in\overline{\mathcal{P}_{N}},y\in Y\right\}>0.$
+$\widetilde{\epsilon_{1}}\coloneqq\inf\left\{W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\middle|{\bf q}\in\overline{\mathcal{P}_{N}},y\in Y\right\}>0.$
 Proof.
 Fix ${\bf q}\in\overline{\mathcal{P}_{N}}$. By Lemma 3.10,
-$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{2}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|.$
-If ${\bf G}(y)\notin B_{\mathscr{A}({\bf q})}$, then, $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}={\bf G}\left(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})\right)$ and hence, $y$ is not contained in an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$.
-Since $G_{j}$ and $H_{\mathscr{A}\left({\bf q}\right)}$ are continuous, $\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$ is continuous with respect to $y$ on $Y$.
+$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{2}\max_{j\in\mathcal{I}_{N}}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|.$
+Let
+$\widetilde{B}_{j}\coloneqq\left\{{\bf p}\in\overline{\mathcal{P}_{N}}\colon\left\|{\bf p}-\mathfrak{p}_{j}\right\|_{1}\leq\delta_{0}/8\right\},\ j\in\mathcal{I}_{N}.$
+If ${\bf G}(y)\notin\widetilde{B}_{\mathscr{A}({\bf q})}$, then $y\neq\mathfrak{p}_{\mathscr{A}({\bf q})}={\bf G}\left(\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})\right)$ and hence, there exists an open interval $J_{\mathscr{A}({\bf q})}$ in $Y$ containing $\textup{Fix}(H_{\mathscr{A}\left({\bf q}\right)})$ but not $y$.
+Since $G_{j}$ and $H_{\mathscr{A}\left({\bf q}\right)}$ are continuous, $\displaystyle\max_{j\in\mathcal{I}_{N}}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$ is continuous with respect to $y$ on $Y$.
 Hence,
-$\inf_{{\bf G}(y)\notin B_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
-$\geq\inf_{y\notin J_{\mathscr{A}({\bf q})}}\max_{j}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|=:\epsilon_{\mathscr{A}\left({\bf q}\right)}>0.$
-Assume that ${\bf G}(y)\in B_{\mathscr{A}({\bf q})}$. Since ${\bf q}\notin B_{\mathscr{A}({\bf q})}$,
-$W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\frac{1}{N}\sum_{j}|G_{j}(y)-q_{j}|\geq\frac{\delta_{0}}{4N}.$
+$\inf_{{\bf G}(y)\notin\widetilde{B}_{\mathscr{A}({\bf q})}}\max_{j\in\mathcal{I}_{N}}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|$
+$\geq\inf_{y\notin J_{\mathscr{A}({\bf q})}}\max_{j\in\mathcal{I}_{N}}\left|G_{j}(H_{\mathscr{A}\left({\bf q}\right)}(y))-G_{j}(y)\right|=:\epsilon_{\mathscr{A}\left({\bf q}\right)}>0.$
+Assume that ${\bf G}(y)\in\widetilde{B}_{\mathscr{A}({\bf q})}$.
+First, we consider the case where $\displaystyle{\bf q}\in\bigcup_{j\in\mathcal{I}_{N}}B_{j}$.
+Then, there exists a unique $i\in\{1,\dots,\ell\}$ such that ${\bf q}\in B_{j_{i}}$ and $\mathscr{A}({\bf q})\neq j_{i}$.
+By the triangle inequality,
+$\left\|G(y)-{\bf q}\right\|_{1}\geq\left\|\mathfrak{p}_{\mathscr{A}({\bf q})}-\mathfrak{p}_{j_{i}}\right\|_{1}-\left\|G(y)-\mathfrak{p}_{\mathscr{A}({\bf q})}\right\|_{1}-\left\|{\bf q}-\mathfrak{p}_{j_{i}}\right\|_{1}\geq\frac{\delta_{0}}{2}.$
+Second, we consider the case where $\displaystyle{\bf q}\notin\bigcup_{j\in\mathcal{I}_{N}}B_{j}$. Then, ${\bf q}\notin B_{\mathscr{A}({\bf q})}$. By the triangle inequality,
+$\left\|G(y)-{\bf q}\right\|_{1}\geq\left\|\mathfrak{p}_{\mathscr{A}({\bf q})}-{\bf q}\right\|_{1}-\left\|\mathfrak{p}_{\mathscr{A}({\bf q})}-G(y)\right\|_{1}\geq\frac{\delta_{0}}{8}.$
+Thus we obtain that
+$W_{\mathscr{A}\left({\bf q}\right)}\left({\bf q},y\right)\geq\frac{1}{N}\sum_{j\in\mathcal{I}_{N}}|G_{j}(y)-q_{j}|\geq\frac{\delta_{0}}{8N}.$
 Hence,
-$\inf_{y\in Y}W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\min\left\{\frac{\delta_{0}}{4N},\epsilon_{j_{k}}\middle|1\leq k\leq\ell\right\}=:\tilde{\epsilon_{1}}.$
+$\inf_{y\in Y}W_{\mathscr{A}\left({\bf q}\right)}({\bf q},y)\geq\min\left\{\frac{\delta_{0}}{8N},\epsilon_{j_{k}}\middle|1\leq k\leq\ell\right\}=:\widetilde{\epsilon_{1}}$$\qed$
 We discuss the frequency of occurrence of some elements in $\mathcal{I}_{N}$ in $(\xi_{n})_{n}$.
 Lemma 3.13.
-Let $I\subset\mathcal{I}_{N}$ be a non-empty set. Let $c_{I}:=\inf_{y\in Y}\sum_{i\in I}G_{i}(y)$. Then,
-$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n:\xi_{j}(z)\in I\}|}{n}\geq c_{I},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
+Let $I\subset\mathcal{I}_{N}$ be a non-empty set. Let $\displaystyle c_{I}\coloneqq\inf_{y\in Y}\sum_{i\in I}G_{i}(y)$. Then,
+$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n\colon\xi_{j}(z)\in I\}|}{n}\geq c_{I},\ \textup{ $\displaystyle\nu_{0}$-a.s. $\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
 This corresponds to [27, Lemma 9], but the formulation differs.
 Proof.
-If $c_{I}=0$, then, the assertion is obvious.
+If $c_{I}=0$, then the assertion is obvious.
 Assume $c_{I}>0$.
+We regard $\{\xi_{j}\}_{j\geq 1}$ as random variables with respect to the probability measure $\nu_{0}$.
 We denote the expectation with respect to $\nu_{0}$ by $E^{\nu_{0}}$.
-Let $\mathcal{F}_{m}:=\sigma(\xi_{1},\cdots,\xi_{m})$ for $m\geq 1$ and $\mathcal{F}_{0}$ be the trivial $\sigma$-algebra on $(\mathcal{I}_{N})^{\mathbb{N}}$.
+Let $\mathcal{F}_{m}\coloneqq\sigma(\xi_{1},\ldots,\xi_{m})$ for $m\geq 1$ and $\mathcal{F}_{0}$ be the trivial $\sigma$-algebra on $(\mathcal{I}_{N})^{\mathbb{N}}$.
 We denote the indicator function of $A\subset(\mathcal{I}_{N})^{\mathbb{N}}$ by ${\bf 1}_{A}$.
-Let $M_{0}:=0$ and
-$M_{n}-M_{n-1}:={\bf 1}_{\{\xi_{n}\in I\}}-E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{n}\in I\}}|\mathcal{F}_{n-1}\right],n\geq 1.$
-Then, $\{M_{n},\mathcal{F}_{n}\}_{n}$ is a martingale.
-By (3.1), $E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{n}\in I\}}|\mathcal{F}_{n-1}\right]\geq c_{I}$, $\nu_{0}$-a.s. Therefore, for every $\eta\in(0,1)$,
-$\nu_{0}\left(\frac{|\{1\leq j\leq n:\xi_{j}\in I\}|}{n}\leq(1-\eta)c_{I}\right)\leq\nu_{0}\left(M_{n}\leq-n\eta c_{I}\right).$
-We remark that $|M_{n}-M_{n-1}|\leq 1$ for each $n\geq 1$. By Azuma’s inequality ([1], [31, E14.2]),
+Let $M_{0}\coloneqq 0$ and
+$M_{j}-M_{j-1}\coloneqq{\bf 1}_{\{\xi_{j}\in I\}}-E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{j}\in I\}}|\mathcal{F}_{j-1}\right],\ j\geq 1.$
+Then, $\left\{M_{n},\mathcal{F}_{n}\right\}_{n}$ is a martingale.
+By (3.1), for every $j\geq 1$, $E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{j}\in I\}}|\mathcal{F}_{j-1}\right]\geq c_{I}$, $\nu_{0}$-a.s. Then, we see that
+$\frac{1}{n}\sum_{i=1}^{n}{\bf 1}_{I}(\xi_{i})=\frac{M_{n}}{n}+\frac{1}{n}\sum_{i=1}^{n}E^{\nu_{0}}\left[{\bf 1}_{\{\xi_{j}\in I\}}|\mathcal{F}_{j-1}\right]\geq\frac{M_{n}}{n}+c_{I},\textup{ $\displaystyle\nu_{0}$-a.s.}$
+Therefore, it holds that for every $\eta\in(0,1)$,
+$\displaystyle\nu_{0}\left(\frac{|\{1\leq j\leq n\colon\xi_{j}\in I\}|}{n}\leq(1-\eta)c_{I}\right)$ $\displaystyle\leq\nu_{0}\left(\frac{M_{n}}{n}+c_{I}\leq(1-\eta)c_{I}\right)$
+$\displaystyle=\nu_{0}\left(M_{n}\leq-n\eta c_{I}\right).$
+We remark that $|M_{n}-M_{n-1}|\leq 1$ for each $n\geq 1$. By Azuma’s inequality ([2], [33, E14.2]),
 $\nu_{0}\left(M_{n}\leq-n\eta c_{I}\right)\leq\exp\left(-\frac{n\eta^{2}c_{I}^{2}}{2}\right).$
 Using $c_{I}>0$ and Borel-Cantelli’s lemma,
-$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n:\xi_{j}(z)\in I\}|}{n}\geq(1-\eta)c_{I},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
-The assertion follows if $\eta\to+0$. ∎
+$\liminf_{n\to\infty}\frac{|\{1\leq j\leq n\colon\xi_{j}(z)\in I\}|}{n}\geq(1-\eta)c_{I},\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
+The assertion follows for $\eta\to+0$. ∎
 We proceed to the proof of Proposition 3.8. The proof is divided into two cases. The following is easily shown.
 Lemma 3.14.
+We have the following assertions:
 (i) $a_{0}\leq 1$ and $b_{N-1}+c_{N-1}\geq 0$.
-(ii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then, $\alpha=-1,\beta=+\infty$.
-(iii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$, then, $\alpha>-1,\beta=+\infty$.
-(iv) If $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$, then, $\alpha=-1,\beta<+\infty$.
-(v) If $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$, then, $\alpha>-1,\beta<+\infty$.
-(vi) If $a_{0}=1$, then, $\textup{Fix}(H_{0})=\{+\infty\}$ and $\mathfrak{p}_{0}=(1,0,\cdots,0)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
-(vii) If $b_{N-1}+c_{N-1}=0$, then, $\textup{Fix}(H_{N-1})=\{-1\}$ and $\mathfrak{p}_{N-1}=(0,\cdots,0,1)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
+(ii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then $\alpha=-1,\beta=+\infty$.
+(iii) If $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$, then $\alpha>-1,\beta=+\infty$.
+(iv) If $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$, then $\alpha=-1,\beta<+\infty$.
+(v) If $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$, then $\alpha>-1,\beta<+\infty$.
+(vi) If $a_{0}=1$, then $\textup{Fix}(H_{0})=\{+\infty\}$ and $\mathfrak{p}_{0}=(1,0,\ldots,0)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
+(vii) If $b_{N-1}+c_{N-1}=0$, then $\textup{Fix}(H_{N-1})=\{-1\}$ and $\mathfrak{p}_{N-1}=(0,\ldots,0,1)\in\overline{\mathcal{P}_{N}}\setminus\mathcal{P}_{N}$.
 $\alpha=-1$ or $\beta=+\infty$ occurs according to the values of $a_{0}$ and $b_{N-1}+c_{N-1}$.
-Case 1. We first deal with the case that $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$.
-In this case, we do not need to assume ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
+Case 1. We first deal with the case where $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$.
 Proof of Proposition 3.8, Case 1.
 By the assumption that $a_{0}<1$ and $b_{N-1}+c_{N-1}>0$ and Lemma 3.14 (v), $-1<\alpha\leq\beta<+\infty$. Hence,
-$10<\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)\leq\sup_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)<$ 1 .
-Fix ${\bf q}\in\overline{\mathcal{P}_{N}}$.
-Let $\widetilde{c}:=\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)$.
-Then, $c_{\{\mathscr{A}({\bf q})\}}\geq\widetilde{c}>0$ and by Lemma 3.13 for $I=\{\mathscr{A}({\bf q})\}$,
-$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}{\bf 1}_{\{\xi_{k}=\mathscr{A}({\bf q})\}}(z)\geq\widetilde{c},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
+$0<\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)\leq\sup_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)<1.$
+Fix ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$. Let $\displaystyle\widetilde{c}\coloneqq\inf_{y\in Y,i\in\mathcal{I}_{N}}G_{i}(y)$. Then,
+$c_{\{\mathscr{A}({\bf q})\}}=\inf_{y\in Y}G_{\mathscr{A}({\bf q})}(y)\geq\widetilde{c}>0,$
+and by Lemma 3.13 for $I=\{\mathscr{A}({\bf q})\}$,
+$\liminf_{n\to\infty}\frac{1}{n}\sum_{k=1}^{n}{\bf 1}_{\{\xi_{k}=\mathscr{A}({\bf q})\}}(z)\geq\widetilde{c},\ \textup{ $\displaystyle\nu_{0}$-a.s.
+$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
 The number $i_{1}\in\mathcal{I}_{N}$ in Lemma 3.11 will be chosen according to the position of ${\bf q}$.
-By applying Lemma 3.11 to $i_{1}=\mathscr{A}({\bf q})$ and $\epsilon_{1}=\widetilde{\epsilon_{1}}$ where $\widetilde{\epsilon_{1}}$ is the constant in Lemma 3.12, we see that if $X_{i+1}(z)=\mathscr{A}({\bf q})$, then,
-$s_{N}\left(p_{i}(0;z)|{\bf q}\right)+s_{N}\left(p_{i+1}(0;z)|{\bf q}\right)\leq V(\widetilde{\epsilon_{1}})<0.$
-Now the assertion follows if we let $\epsilon_{2}:=\dfrac{\widetilde{c}\ V(\widetilde{\epsilon_{1}})}{2}$. ∎
-Case 2. We now deal with the case that $a_{0}=1$ or $b_{N-1}+c_{N-1}=0$.
-If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then, $c_{\{0\}}=c_{\{N-1\}}=0$, so we consider $c_{\{0,N-1\}}$.
-In the following, we do not use Lemma 3.12 or the assumption that |$|\{\mathfrak{p}_{i}\}_{i}|\geq 2$.
-On the other hand, we assume ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
+By applying Lemma 3.11 to $i_{1}=\mathscr{A}({\bf q})$ and $\epsilon_{1}=\widetilde{\epsilon_{1}}$ where $\widetilde{\epsilon_{1}}$ is the constant in Lemma 3.12, we see that if $\xi_{i+1}(z)=\mathscr{A}({\bf q})$, then
+$s_{N}\left(p_{i}(0;z)|{\bf q}\right)+s_{N}\left(p_{i+1}(0;z)|{\bf q}\right)\leq V_{\epsilon,1}(\widetilde{\epsilon_{1}})<0.$
+Now the assertion follows if we let $\epsilon_{2}\coloneqq-\dfrac{\widetilde{c}\ V_{\epsilon,1}(\widetilde{\epsilon_{1}})}{2}>0$. ∎
+Case 2. We now deal with the case where $a_{0}=1$ or $b_{N-1}+c_{N-1}=0$.
+If $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$, then $c_{\{0\}}=c_{\{N-1\}}=0$, so we consider $c_{\{0,N-1\}}$.
+In the following, we use neither Lemma 3.12 nor the assumption that $\left|\left\{\mathfrak{p}_{i}\right\}_{i}\right|\geq 2$.
 Proof of Proposition 3.8, Case 2.
-Case 2.1. We first consider the case that $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$.
-By Lemma 3.14, $\alpha>-1$ and $\beta=+\infty$, $H_{0}(y)=y+c_{0}$, and $\textup{Fix}(H_{0})=\{+\infty\}$.
-$\mathfrak{p}_{0}=(1,0,\cdots,0)\notin\mathcal{P}_{N}$.
+We consider the following three cases.
+Case 2.1. We consider the case where $a_{0}=1$ and $b_{N-1}+c_{N-1}>0$.
+By Lemma 3.14, we see that $\alpha>-1$ and $\beta=+\infty$.
+It holds that $H_{0}(y)=y+c_{0}$, and hence $\textup{Fix}(H_{0})=\{+\infty\}$.
+Hence, $\mathfrak{p}_{0}={\bf G}(\textup{Fix}(H_{0}))=(1,0,\ldots,0)\notin\mathcal{P}_{N}$.
 On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\geq 1$.
-Since $G_{0}(+\infty)=G_{0}(H_{0}(+\infty))=1$ and $q_{0}<1-\epsilon/2$,
-$\inf_{y\in Y}W_{0}({\bf q},y)>c_{0}(\epsilon/2)>0,$
-where the lower bound $c_{0}(\epsilon/2)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
-Since $\alpha>-1$, $\inf_{y\in Y}G_{0}(y)\geq c>0$. By using Lemma 3.11 and Lemma 3.13 for $I=\{0\}$, we have the assertion.
-Case 2.2. We secondly consider the case that $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$.
-We recall that $\alpha=-1$ and $\beta<+\infty$, By Lemma 3.14, $H_{N-1}(y)=\dfrac{(1-2b_{N-1})y+b_{N-1}}{1-b_{N-1}}$, and $\textup{Fix}(H_{N-1})=\{-1\}$.
-$\mathfrak{p}_{N-1}=(0,\cdots,0,1)\notin\mathcal{P}_{N}$.
+Since $G_{0}(+\infty)=G_{0}(H_{0}(+\infty))=1$ and $q_{0}\leq 1-\epsilon/2$,
+$\inf_{y\in Y}W_{0}({\bf q},y)\geq\widetilde{c_{0}}(\epsilon/2),$
+where the positive lower bound $\widetilde{c_{0}}(\epsilon/2)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
+Since $\alpha>-1$, it holds that $\displaystyle\inf_{y\in Y}G_{0}(y)\geq c$ for some $c>0$. By using Lemma 3.11 and Lemma 3.13 for $I=\{0\}$, we have the assertion.
+Case 2.2. We consider the case where $a_{0}<1$ and $b_{N-1}+c_{N-1}=0$.
+By Lemma 3.14, we see that $\alpha=-1$ and $\beta<+\infty$.
+It holds that $H_{N-1}(y)=\dfrac{(1-2b_{N-1})y+b_{N-1}}{1-b_{N-1}}$, and hence $\textup{Fix}(H_{N-1})=\{-1\}$.
+Hence, $\mathfrak{p}_{N-1}={\bf G}(\textup{Fix}(H_{N-1}))=(0,\ldots,0,1)\notin\mathcal{P}_{N}$.
 On the other hand, $\mathfrak{p}_{i}\in\mathcal{P}_{N}$ for $i\leq N-2$.
 Since $G_{N-1}(-1)=G_{N-1}(H_{N-1}(-1))=1$ and $q_{N-1}<1-\epsilon/2$,
-$\inf_{y\in Y}W_{N-1}({\bf q},y)>c_{N-1}(\epsilon/2)>0,$
-where the lower bound $c_{N-1}(\epsilon)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
-Since $\beta<+\infty$, $\inf_{y\in Y}G_{N-1}(y)\geq c>0$.
+$\inf_{y\in Y}W_{N-1}({\bf q},y)\geq\widetilde{c_{N-1}}(\epsilon/2),$
+where the positive lower bound $\widetilde{c_{N-1}}(\epsilon)$ is uniform with respect to ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
+Since $\beta<+\infty$, it holds that $\displaystyle\inf_{y\in Y}G_{N-1}(y)\geq c$ for some $c>0$.
 The rest of the argument is the same as in the above case.
-Case 2.3. We finally consider the case that $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$.
-Since $\textup{Fix}(H_{0})=\{+\infty\}$, $\textup{Fix}(H_{N-1})=\{-1\}$, $q_{0}<1-\epsilon/2$ and $q_{N-1}<1-\epsilon/2$, it holds that
-$\inf_{y\in Y}W_{0}({\bf q},y)>c_{0}(\epsilon/2)>0$
+Case 2.3. We consider the case where $a_{0}=1$ and $b_{N-1}+c_{N-1}=0$.
+Since $\textup{Fix}(H_{0})=\{+\infty\}$, $\textup{Fix}(H_{N-1})=\{-1\}$, $q_{0}\leq 1-\epsilon/2$ and $q_{N-1}\leq 1-\epsilon/2$, it holds that
+$\inf_{y\in Y}W_{0}({\bf q},y)>\widetilde{c_{0}}(\epsilon/2)>0$
 and
-$\inf_{y\in Y}W_{N-1}({\bf q},y)>c_{N-1}(\epsilon/2)>0.$
+$\inf_{y\in Y}W_{N-1}({\bf q},y)>\widetilde{c_{N-1}}(\epsilon/2)>0.$
 Since
-$G_{0}(y)+G_{N-1}(y)=\frac{b_{1}b_{N-1}y^{2}+2b_{1}y1+b_{1}-b_{N-1}}{(b_{1}y+1)(b_{N-1}y+1)},$
+$G_{0}(y)+G_{N-1}(y)=\frac{b_{1}b_{N-1}y^{2}+2b_{1}y+1+b_{1}-b_{N-1}}{(b_{1}y+1)(b_{N-1}y+1)},$
 it holds that
@@
 By using Lemma 3.11 and Lemma 3.13 for $I=\{0,N-1\}$, we have the assertion. ∎
-Now the proof of Theorem 3.2 completes.
+Now the proof of Theorem 3.2 is completed.
 Remark 3.15.
 (i) In the proof of [27, Theorem 43], the details of Case 2 were not given.
-(ii) Assume that $\{g_{i}\}_{i}$ are affine functions such that $g_{i}(x)=b_{i}x+(b_{0}+\cdots+b_{i-1})$.
+(ii) Assume that $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are affine maps such that $g_{i}(x)=b_{i}x+(b_{0}+\cdots+b_{i-1})$.
 Since $Y=\{0\}$, there is no need to introduce $H_{i}$, and $W_{i}({\bf q},y)=\max_{k}|G_{k}(y)-q_{k}|=\max_{k}|b_{k}-q_{k}|$.
-However, $\mathscr{p}_{0}=\cdots=\mathscr{p}_{N-1}=(b_{i})_{i}$, and hence this is not in the scope of Theorem 3.2.
+However, it holds that $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}=(b_{i})_{i}$, and hence this is not in the scope of Theorem 3.2.
 This case is considered in Theorem 3.3.
 We now proceed to the proof of Theorem 3.3.
-By (3.6), if we take $\delta>0$ such that
-$\frac{\log N}{\log(N/(1+\delta))}\leq 1+\frac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)},$ (3.9)
-then, $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$.
-Assume that $s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$.
-Since $\min_{i}r_{i}\geq\epsilon$, $\sum_{i}|r_{i}-r_{i}^{s}|\leq\epsilon^{1-s}-1\leq\epsilon/2$.
-Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$, where we let $\widetilde{q}_{i}:=r_{i}^{s}$ for each $i\in\mathcal{I}_{N}$ and ${\bf\widetilde{q}}:=\left(\widetilde{q}_{i}\right)_{i\in\mathcal{I}_{N}}$.
-Since $\epsilon\leq 1/2$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\epsilon/2$ for each $i\in\mathcal{I}_{N}$.
+By Lemma 3.4, if we take $\delta>0$ such that
+(3.10) $\frac{\log N}{\log(N/(1+\delta))}\leq 1+\frac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)},$
+then $\displaystyle s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$.
+Assume that $\displaystyle s\leq 1+\dfrac{\log\left(1+\epsilon/2\right)}{\log(1/\epsilon)}$.
+Since $\displaystyle\min_{i\in\mathcal{I}_{N}}r_{i}\geq\epsilon$, $\displaystyle\sum_{i\in\mathcal{I}_{N}}|r_{i}-r_{i}^{s}|\leq\epsilon^{1-s}-1\leq\epsilon/2$.
+Hence, $\left\|{\bf\widetilde{q}}-\mathfrak{p}_{0}\right\|_{1}\geq\epsilon/2$, where we let $\widetilde{q}_{i}\coloneqq r_{i}^{s}$ for each $i\in\mathcal{I}_{N}$ and ${\bf\widetilde{q}}\coloneqq\left(\widetilde{q}_{i}\right)_{i\in\mathcal{I}_{N}}$.
+Since $\epsilon\in(0,1)$, $\widetilde{q}_{i}\geq\epsilon^{s}\geq\dfrac{\epsilon}{1+\epsilon/2}\geq\epsilon/2$ for each $i\in\mathcal{I}_{N}$.
 Hence, ${\bf\widetilde{q}}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$.
 By Lemma 3.10 (i) and (ii), it holds that
-$\epsilon_{3}:=\inf\left\{W_{i}({\bf q},y)\middle|i\in\mathcal{I}_{N},y\in Y,{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2\right\}>0.$
-By Lemma 3.11, we see that for every $j\geq 0$, every $z\in(\mathcal{I}_{N})^{\mathbb{N}}$, and every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$ with $\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2$,
+$\epsilon_{3}\coloneqq\inf\left\{W_{i}({\bf q},y)\middle|i\in\mathcal{I}_{N},y\in Y,{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|_{1}\geq\epsilon/2\right\}>0.$
+By Lemma 3.11, we see that for every $j\geq 0$, every $z\in(\mathcal{I}_{N})^{\mathbb{N}}$, and every ${\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}$ with $\left\|{\bf q}-\mathfrak{p}_{0}\right\|_{1}\geq\epsilon/2$,
 $s_{N}\left(p_{j}(0;z)|{\bf q}\right)+s_{N}\left(p_{j+1}(0;z)|{\bf q}\right)\leq V\left({\bf q};\epsilon_{3}\right).$
-By Lemma 3.9, we can let
-$V_{\epsilon}(\epsilon_{3}):=\max\left\{V\left({\bf q};\epsilon_{3}\right)\middle|{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|\geq\epsilon/2\right\}<0.$
+By Lemma 3.9, we see that
+$V_{\epsilon,2}(\epsilon_{3})\coloneqq\max\left\{V\left({\bf q};\epsilon_{3}\right)\middle|{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)},\left\|{\bf q}-\mathfrak{p}_{0}\right\|_{1}\geq\epsilon/2\right\}<0.$
 Then,
-$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-\frac{V_{\epsilon}(\epsilon_{3})}{2},\ \textup{ $\nu_{0}$-a.s.$z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
-Assume that $\delta>0$ satisfies not only (3.9) but also (3.5) for $\epsilon_{2}:=V_{\epsilon}(\epsilon_{3})/2$. Then, by Propositions 3.6 and 3.7, $\dim_{H}\mu_{\varphi}<1$.
+$\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq\frac{V_{\epsilon,2}(\epsilon_{3})}{2},\ \textup{ $\displaystyle\nu_{0}$-a.s.
+$\displaystyle z\in(\mathcal{I}_{N})^{\mathbb{N}}$.}$
+Assume that $\delta>0$ satisfies not only (3.10) but also (3.6) for $\epsilon_{2}\coloneqq V_{\epsilon}(\epsilon_{3})/2$. Then, by Propositions 3.6 and 3.7, $\dim_{H}\mu_{\varphi}<1$.
 Remark 3.16.
-(i) We do not need to use Lemma 3.13. For every small $\epsilon>0$ and every $i\in\mathcal{I}_{N}$,
+(i) If $\mathfrak{p}_{0}=\cdots=\mathfrak{p}_{N-1}$, then $\mathfrak{p}_{0}\in\mathcal{P}_{N}$.
+We show this.
+Since $G_{0}(y)=\dfrac{b_{1}(y+1)}{b_{1}y+1}$ and $0<b_{1}<1$, it is injective on $Y$.
+Hence, $\textup{Fix}(H_{0})=\cdots=\textup{Fix}(H_{N-1})$.
+Hence, $\displaystyle\frac{c_{0}}{1-a_{0}}=\frac{a_{1}-1+\sqrt{(1-a_{1})^{2}+4b_{1}c_{1}}}{2b_{1}}<\infty$.
+By Definition 2.8 and the assumption that $d_{0}=1$, we see that $a_{0}\in(0,1)$ and $\dfrac{a_{0}}{c_{0}+1}=b_{1}\in(0,1)$.
+Hence, $\displaystyle\frac{c_{0}}{1-a_{0}}>-1$.
+Recall that $\displaystyle G_{i}(y)=\frac{(b_{i+1}-b_{i})(y+1)}{(b_{i}y+1)(b_{i+1}y+1)},i\in\mathcal{I}_{N}$.
+Since $0=b_{0}<b_{1}<\cdots<b_{N-1}<b_{N}=1$, $G_{i}(y)>0$ for every $y\in(-1,+\infty)$ and every $i\in\mathcal{I}_{N}$.
+Hence, $\displaystyle\mathfrak{p}_{0}=\left(G_{i}\left(\frac{c_{0}}{1-a_{0}}\right)\right)_{i\in\mathcal{I}_{N}}\in\mathcal{P}_{N}$.
+(ii) In the proof of Theorem 3.3, we do not need to use Lemma 3.13, since we have the estimate that $\epsilon_{3}>0$. We remark that for every small $\epsilon>0$ and every $i\in\mathcal{I}_{N}$,
 $\min\left\{W_{i}({\bf q},y)\middle|y\in Y,{\bf q}\in\overline{\mathcal{P}_{N}(\epsilon/2)}\right\}=0,$
-by Lemma 3.10 (ii).
-(ii) The case that both of $\{f_{i}\}_{i}$ and $\{g_{i}\}_{i}$ are affine is easily handled by the dimension formula.
-However, if some of the functions $\{f_{i},g_{i}\}_{i}$ are non-affine, then, it is difficult to judge the singularity by a generalized dimension formula in [9].
-4. Example
-We give three examples that $\{g_{i}\}_{i}$ is an LF system and $\{f_{i}\}_{i}$ is non-affine and the solution $\varphi$ is singular. The result depends on the quantitative estimates in Section 3.
-We first deal with the case that $N=2$. Then,
+because, by (i) above, we can let ${\bf q}={\bf G}\left(\textup{Fix}(H_{i})\right)=\mathfrak{p}_{i}=\mathfrak{p}_{j},j\in\mathcal{I}_{N}$ and $y=\textup{Fix}(H_{i})$.
+(iii) The case where both $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ and $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ are affine maps is easily handled by the dimension formula.
+However, if some of the functions $\{f_{i},g_{i}\}_{i}$ are non-affine maps, then it is difficult to judge the singularity by the generalized dimension formula in [11, Corollary 3.5].
+4. Examples
+We give three examples assuming that $\{g_{i}\}_{i\in\mathcal{I}_{N}}$ is an LF system and $\{f_{i}\}_{i\in\mathcal{I}_{N}}$ are non-affine maps and the solution $\varphi$ is singular.
+The result depends on the quantitative estimates in Section 3.
+We assume that $d_{i}=1$ for every $i\in\mathcal{I}_{N}$.
+We first deal with the case where $N=2$. Then,
 $G_{0}(y)=\frac{b_{1}(y+1)}{b_{1}y+1},\ H_{0}(y)=a_{0}y+c_{0},\ H_{1}(y)=\frac{(1-b_{1}+c_{1})y+c_{1}}{b_{1}y+1}.$
@@
 Let
-$g_{0}(x):=\frac{x}{-x+6},\textup{ and }g_{1}(x):=\frac{3x+1}{5-x}.$
+$g_{0}(x)\coloneqq\frac{x}{-x+6}=\frac{\frac{x}{6}}{-\frac{x}{6}+1},\textup{ and }g_{1}(x)\coloneqq\frac{3x+1}{5-x}=\frac{\frac{3}{5}x+\frac{1}{5}}{-\frac{x}{5}+1}.$
 Then, $g_{0}$ is contractive, and $g_{1}$ is weakly contractive but not contractive. We see that
@@
 Hence, $\textup{Fix}(H_{0})=\{-1/5\}$ and $\textup{Fix}(H_{1})=\{-1\}$. Hence,
-${\bf G}(\textup{Fix}(H_{0}))=\left(\frac{1}{6},\frac{5}{6}\right)\textup{ and }{\bf G}(\textup{Fix}(H_{1}))=(0,1).$
-Let $\widetilde{\mathcal{P}_{2}}:=\left\{(p_{0},p_{1})\in\mathcal{P}_{2}:p_{0}\geq 4/5\right\}$. Then,
-$|G_{0}(y)-q_{0}|=|G_{1}(y)-q_{1}|\geq\frac{3}{5},\ y\in[-1,0],\ {\bf q}=(q_{0},q_{1})\in\widetilde{\mathcal{P}_{2}}.$
-Hence,
-$s_{2}\left({\bf G}(y)|{\bf q}\right)\leq V\left({\bf q};\frac{6}{5}\right),\ y\in[-1,0],\ {\bf q}\in\widetilde{\mathcal{P}_{2}}.$ (4.1)
-Since $s_{2}({\bf p}|{\bf q})=s_{2}\left((p_{0},1-p_{0})|(q_{0},1-q_{0})\right)$ is increasing with respect to $p_{0}$ on $[0,q_{0}]$,
-$V\left({\bf q};\frac{6}{5}\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|(q_{0},1-q_{0})\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|\left(\frac{4}{5},\frac{1}{5}\right)\right)\leq-\frac{6}{5}\log 2\leq-0.83.$
+${\bf G}\left(\textup{Fix}(H_{0})\right)=\left(\frac{1}{6},\frac{5}{6}\right)\textup{ and }{\bf G}(\textup{Fix}(H_{1}))=(0,1).$
+Let $\widetilde{\mathcal{P}_{2}}\coloneqq\left\{(p_{0},p_{1})\in\mathcal{P}_{2}\colon p_{0}\geq 4/5\right\}$. Then,
+$\left|G_{0}(y)-q_{0}\right|=\left|G_{1}(y)-q_{1}\right|\geq\frac{3}{5},\ y\in[-1,0],\ {\bf q}=(q_{0},q_{1})\in\widetilde{\mathcal{P}_{2}}$.
+Since $s_{2}\left({\bf p}|{\bf q}\right)=s_{2}\left((p_{0},1-p_{0})|(q_{0},1-q_{0})\right)$ is increasing with respect to $p_{0}$ on $[0,q_{0}]$,
+$\displaystyle s_{2}\left({\bf G}(y)|{\bf q}\right)\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|(q_{0},1-q_{0})\right)$ $\displaystyle\leq s_{2}\left(\left(\frac{1}{5},\frac{4}{5}\right)\middle|\left(\frac{4}{5},\frac{1}{5}\right)\right)$
+(4.1) $\displaystyle\leq-\frac{6}{5}\log 2\leq-0.83.$
 By Proposition 3.7, it holds that
-$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.83,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{2}}$.}$ (4.2)
-Let $\epsilon\in(0,1/4)$.
-Let $\{f_{0},f_{1}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1$.
-Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}=1$.
+(4.2) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}=\limsup_{n\to\infty}\frac{1}{n}\sum_{j=0}^{n-1}s_{N}\left(p_{j}(0;z)|{\bf q}\right)\leq-0.83,\ \textup{ $\displaystyle\nu_{0}$-a.s.
+$\displaystyle z$, \ $\displaystyle{\bf q}\in\widetilde{\mathcal{P}_{2}}$.}$
+Let $\epsilon\in(0,1/4)$ and let $\{f_{0},f_{1}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}\coloneqq\|f_{i}\|_{\textup{Lip}}$, $i=0,1$.
+Let $s\geq 1$ be the constant such that $r_{0}^{s}+r_{1}^{s}=1$.
 Assume that $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$.
@@
 $\dim_{H}\mu_{\varphi}\leq s-\frac{0.83}{\log(1/\epsilon)}.$
-We see that $\dfrac{0.83}{\log 10}\geq 0.36$ and $\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$.
-Hence, (3.5) holds for $N=2$, $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$.
-By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, $r_{0}+r_{1}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$, then,
+By Lemma 3.4, if $\{f_{0},f_{1}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1$, and $r_{0}+r_{1}\leq 1.2$, then $s\leq\dfrac{\log 2}{\log 2-\log 1.2}$.
+Since we see that $\dfrac{0.83}{\log 10}\geq 0.36$ and $\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$, equation (3.6) holds for $N=2$, $\epsilon=0.1$, $\epsilon_{2}=0.83$ and $\delta=0.2$.
+Therefore, if $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$, then
 $\dim_{H}\mu_{\varphi}\leq s-\frac{0.83}{\log(1/\epsilon)}\leq\dfrac{\log 2}{\log 2-\log 1.2}-\frac{0.83}{\log 10}<1.$
@@
 Then, $r_{0}+r_{1}\leq 1.2$.
-The maps $f_{0}$ and $f_{1}$ can be non-affine.
-Since $s\leq\frac{\log 2}{\log 2-\log 1.2}\leq 1.357$, $r_{0}^{s}\geq(0.85)^{1.357}\geq 0.8$.
-Hence, $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$.
+The maps $f_{0}$ and $f_{1}$ can be non-affine maps.
+Since $s\leq\dfrac{\log 2}{\log 2-\log 1.2}\leq 1.357$, we have $r_{0}^{s}\geq(0.85)^{1.357}\geq 0.8$.
+Hence, $(r_{0}^{s},r_{1}^{s})\in\widetilde{\mathcal{P}_{2}}$ and we see that $\dim_{H}\mu_{\varphi}<1$.
 In this case, we see that $a_{0}<1$ and $b_{1}+c_{1}=0$.
 Hence, this case corresponds to Case 2.2.
-However, we do not need to apply Lemma 3.11 since we have the upper bound for the relative entropy in (4.1).
+However, we do not need to apply Lemma 3.11 since we have the direct upper bound for the relative entropy in (4.1).
 Example 4.2.
 Let $N=3$. Let
-$g_{0}(x):=\frac{x}{8-x},\ g_{1}(x):=\frac{5x+1}{7},\textup{ and }g_{2}(x):=\frac{6}{7-x}.$
+$g_{0}(x)\coloneqq\frac{x}{8-x},\ g_{1}(x)\coloneqq\frac{5x+1}{7},\textup{ and }g_{2}(x)\coloneqq\frac{6}{7-x}.$
 Then,
@@
 Hence, $\textup{Fix}(H_{0})=\{-1/7\}$, $\textup{Fix}(H_{1})=\{0\}$ and $\textup{Fix}(H_{2})=\{-1/6\}$. Hence,
-${\bf G}(\textup{Fix}(H_{0}))=\left(\frac{1}{8},\frac{245}{344},\frac{7}{43}\right),{\bf G}(\textup{Fix}(H_{1}))=\left(\frac{1}{7},\frac{5}{7},\frac{1}{7}\right),\textup{ and }{\bf G}(\textup{Fix}(H_{2}))=\left(\frac{5}{41},\frac{175}{246},\frac{1}{6}\right).$
+${\bf G}(\textup{Fix}(H_{0}))=\left(\frac{1}{8},\frac{245}{344},\frac{7}{43}\right),{\bf G}(\textup{Fix}(H_{1}))=\left(\frac{1}{7},\frac{5}{7},\frac{1}{7}\right),\textup{ and }$
+${\bf G}(\textup{Fix}(H_{2}))=\left(\frac{5}{41},\frac{175}{246},\frac{1}{6}\right).$
 Let
-$\widetilde{\mathcal{P}_{3}}:=\left\{(p_{0},p_{1},p_{2})\in\overline{\mathcal{P}_{3}}\ \middle|\ p_{0},p_{2}\in\left[\frac{7}{16},\frac{1}{2}\right]\right\}.$
+$\widetilde{\mathcal{P}_{3}}\coloneqq\left\{(p_{0},p_{1},p_{2})\in\overline{\mathcal{P}_{3}}\ \middle|\ p_{0},p_{2}\in\left[\frac{7}{16},\frac{1}{2}\right]\right\}.$
 We now show that
-$s_{3}\left({\bf G}(y)|{\bf q}\right)\leq-0.74,\ y\in[-1/6,0],\ {\bf q}\in\widetilde{\mathcal{P}_{3}}.$ (4.3)
-For $q>0$, let $f_{q}(x):=x\log(x/q)$ for $x\in[0,q]$.
+(4.3) $s_{3}\left({\bf G}(y)|{\bf q}\right)\leq-0.74,\ y\in[-1/6,0],\ {\bf q}\in\widetilde{\mathcal{P}_{3}}$.
+For $q>0$, let $f_{q}(x)\coloneqq x\log(q/x)$ for $x\in[0,1]$.
 Then, $f_{q}$ is increasing on $[0,q/e]$ and decreasing on $[q/e,1]$.
+Furthermore, for each $x\geq 0$, $f_{q}(x)$ is increasing with respect to $q$.
 Since ${\bf q}=(q_{0},q_{1},q_{2})\in\widetilde{\mathcal{P}_{3}}$, it holds that $q_{0},q_{2}\leq 1/2$ and $q_{1}\leq 1/8$.
+We also see that $\max\{G_{0}(y),G_{2}(y)\}\leq\dfrac{1}{6}<\dfrac{1}{2e}$ and $\dfrac{1}{8e}<\dfrac{2}{3}\leq G_{1}(y)$.
 Therefore, for $y\in[-1/6,0]$,
-$s_{3}\left({\bf G}(y)|{\bf q}\right)\leq f_{1/2}(G_{0}(y))+f_{1/8}(G_{1}(y))+f_{1/2}(G_{2}(y))$
-$\leq 2f_{1/2}(1/6)+f_{1/8}(2/3)=\log 3-\frac{8}{3}\log 2\leq-0.74.$
+$\displaystyle s_{3}\left({\bf G}(y)|{\bf q}\right)$ $\displaystyle=f_{q_{0}}(G_{0}(y))+f_{q_{1}}(G_{1}(y))+f_{q_{2}}(G_{2}(y))$
+$\displaystyle\leq f_{1/2}(G_{0}(y))+f_{1/8}(G_{1}(y))+f_{1/2}(G_{2}(y))$
+$\displaystyle\leq 2f_{1/2}(1/6)+f_{1/8}(2/3)=\log 3-\frac{8}{3}\log 2\leq-0.74.$
 Thus we have (4.3).
-By Proposition 3.6, it holds that
-$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.74,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{3}}$.}$ (4.4)
-Let $\epsilon\in(0,1/4)$.
-Let $\{f_{0},f_{1},f_{2}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1,2$.
-Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}+r_{2}^{s}=1$.
-Assume that $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$.
-Then, by (4.4) and Lemma 3.5,
+By Proposition 3.7,
+(4.4) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-0.74,\ \textup{ $\displaystyle\nu_{0}$-a.s. $\displaystyle z$, \ $\displaystyle{\bf q}\in\widetilde{\mathcal{P}_{3}}$.}$
+[…]
 $\dim_{H}\mu_{\varphi}\leq s-\frac{0.74}{\log(1/\epsilon)}.$
-We see that $\dfrac{0.74}{\log 20}>0.24$ and $\dfrac{\log 3}{\log 3-\log 1.1}<1.1$.
-Hence, (3.5) holds for $N=3$, $\epsilon=0.05$, $\epsilon_{2}=0.74$ and $\delta=0.1$.
-By Lemma 3.4, if $\{f_{0},f_{1},f_{2}\}$ is a contractive system such that $0.05\leq r_{i}\leq 0.95$, $i=0,1,2$, $r_{0}+r_{1}+r_{2}\leq 1.1$, and $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$, then,
+By Lemma 3.4, if $\{f_{0},f_{1},f_{2}\}$ is a contractive system such that $0.05\leq r_{i}\leq 0.95$, $i=0,1,2$, and $r_{0}+r_{1}+r_{2}\leq 1.1$, then $s\leq\dfrac{\log 3}{\log 3-\log 1.1}$.
+Since we see that $\dfrac{0.74}{\log 20}>0.24$ and $\dfrac{\log 3}{\log 3-\log 1.1}<1.1$, equation (3.6) holds for $N=3$, $\epsilon=0.05$, $\epsilon_{2}=0.74$ and $\delta=0.1$.
+Therefore, if $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$, then
 $\dim_{H}\mu_{\varphi}\leq s-\frac{0.74}{\log(1/\epsilon)}\leq\dfrac{\log 3}{\log 3-\log 1.1}-\frac{0.74}{\log 20}<1.$
@@
 Then, $1\leq r_{0}+r_{1}+r_{2}\leq 1.1$.
-The maps $f_{0}$, $f_{1}$ and $f_{2}$ can be non-affine.
-Since $s\leq\frac{\log 3}{\log 3-\log 1.1}\leq 1.1$, $\min\{r_{0}^{s},r_{2}^{s}\}\geq(0.475)^{1.1}\geq\frac{7}{16}$.
-Hence, $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$.
+The maps $f_{0}$, $f_{1}$ and $f_{2}$ can be non-affine maps.
+Since $\displaystyle s\leq\frac{\log 3}{\log 3-\log 1.1}\leq 1.1$, we have $\min\{r_{0}^{s},r_{2}^{s}\}\geq(0.475)^{1.1}\geq\frac{7}{16}$.
+Hence, $(r_{0}^{s},r_{1}^{s},r_{2}^{s})\in\widetilde{\mathcal{P}_{3}}$ and $\dim_{H}\mu_{\varphi}<1$.
 Example 4.3.
@@
 Let
-$\widetilde{\mathcal{P}_{4}}:=\left\{(p_{0},p_{1},p_{2},p_{3})\in\overline{\mathcal{P}_{4}}\ \middle|\ p_{0},p_{3}\in\left[0,\frac{1}{8}\right]\right\}.$
+$\widetilde{\mathcal{P}_{4}}\coloneqq\left\{(p_{0},p_{1},p_{2},p_{3})\in\overline{\mathcal{P}_{4}}\ \middle|\ p_{0},p_{3}\in\left[0,\frac{1}{8}\right]\right\}.$
 Then, for ${\bf q}=(q_{0},q_{1},q_{2},q_{3})\in\widetilde{\mathcal{P}_{4}}$,
 $s_{4}\left({\bf G}(0)|{\bf q}\right)=\log 3+\frac{1}{3}\log 2+\frac{1}{6}(2\log q_{0}+\log q_{1}+\log q_{2}+2\log q_{3}).$
-Since $\sum_{i=0}^{3}q_{i}=1$, $q_{1}+q_{2}=1-q_{0}-q_{3}$, and hence, $\log q_{1}+\log q_{2}\leq 2\log((1-q_{0}-q_{3})/2)$. Therefore,
+Since $\displaystyle\sum_{i=0}^{3}q_{i}=1$, $q_{1}+q_{2}=1-q_{0}-q_{3}$, and hence, $\log q_{1}+\log q_{2}\leq 2\log((1-q_{0}-q_{3})/2)$. Therefore,
 $s_{4}\left({\bf G}(0)|{\bf q}\right)\leq\log 3+\frac{1}{3}\left(\log q_{0}+\log q_{3}+\log(1-q_{0}-q_{3})\right).$
-Let $f(x,y):=\log x+\log y+\log(1-x-y)$ for $x,y\in[0,1/8]$. Then,
+Let $f(x,y)\coloneqq\log x+\log y+\log(1-x-y)$ for $x,y\in[0,1/8]$. Then,
 $\max_{x,y\in[0,1/8]}f(x,y)=f\left(\frac{1}{8},\frac{1}{8}\right)=\log 3-8\log 2.$
@@
 $s_{4}\left({\bf G}(0)|{\bf q}\right)\leq\frac{4}{3}\log\frac{3}{4}\leq-0.38,\ {\bf q}\in\widetilde{\mathcal{P}_{4}}$.
-By Proposition 3.6, it holds that
-$\limsup_{n\to\infty}\frac{\log R_{n}({\bf q};z)}{n}\leq-0.38,\ \textup{ $\nu_{0}$-a.s.$z$, \ ${\bf q}\in\widetilde{\mathcal{P}_{4}}$.}$ (4.5)
-Let $\epsilon\in(0,1/4)$.
-Let $\{f_{0},f_{1},f_{2},f_{3}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}:=\textup{Lip}(f_{i})$, $i=0,1,2,3$.
-Let $s\geq 1$ be a constant such that $r_{0}^{s}+r_{1}^{s}+r_{2}^{s}+r_{3}^{s}=1$.
+By Proposition 3.7,
+(4.5) $\limsup_{n\to\infty}\frac{-\log R_{n}({\bf q};z)}{n}\leq-0.38,\ \textup{ $\displaystyle\nu_{0}$-a.s.$\displaystyle z$, \ $\displaystyle{\bf q}\in\widetilde{\mathcal{P}_{4}}$.}$
+Let $\epsilon\in(0,1/4)$ and let $\{f_{0},f_{1},f_{2},f_{3}\}$ be a contractive system such that $\epsilon\leq r_{i}\leq 1-\epsilon$, where $r_{i}\coloneqq\|f_{i}\|_{\textup{Lip}}$, $i=0,1,2,3$.
+Let $s\geq 1$ be the constant such that $r_{0}^{s}+r_{1}^{s}+r_{2}^{s}+r_{3}^{s}=1$.
 Assume that $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$.
@@
 $\dim_{H}\mu_{\varphi}\leq s-\frac{0.38}{\log(1/\epsilon)}.$
-We see that $\dfrac{0.38}{\log 10}>0.165$ and $1.1521.15<\dfrac{\log 4}{\log 4-\log 1.2}<$ 1.152.
-Hence, (3.5) holds for $N=4$, $\epsilon=0.1$, $\epsilon_{2}=0.38$ and $\delta=0.2$.
-By Lemma 3.4, if $\{f_{0},f_{1},f_{2},f_{3}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1,2,3$, $r_{0}+r_{1}+r_{2}+r_{3}\leq 1.2$, and $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$, then,
+By Lemma 3.4, if $\{f_{0},f_{1},f_{2},f_{3}\}$ is a contractive system such that $0.1\leq r_{i}\leq 0.9$, $i=0,1,2,3$, and $r_{0}+r_{1}+r_{2}+r_{3}\leq 1.2$, then $s\leq\dfrac{\log 4}{\log 4-\log 1.2}$.
+Since we see that $\dfrac{0.38}{\log 10}>0.165$ and $1.15<\dfrac{\log 4}{\log 4-\log 1.2}<1.152$, equation (3.6) holds for $N=4$, $\epsilon=0.1$, $\epsilon_{2}=0.38$ and $\delta=0.2$.
+Therefore, if $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$, then
 $\dim_{H}\mu_{\varphi}\leq s-\frac{0.38}{\log(1/\epsilon)}\leq\dfrac{\log 4}{\log 4-\log 1.2}-\frac{0.38}{\log 10}<1.$
@@
 Then, $1\leq r_{0}+r_{1}+r_{2}+r_{3}\leq 1.2$.
-The maps $f_{0}$, $f_{1}$, $f_{2}$ and $f_{3}$ can be non-affine.
-Since $s\geq 1$, $\max\{r_{0}^{s},r_{3}^{s}\}\leq 1/8$ and hence $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$.
+The maps $f_{0}$, $f_{1}$, $f_{2}$ and $f_{3}$ can be non-affine maps.
+Since $s\geq 1$, we have $\max\{r_{0}^{s},r_{3}^{s}\}\leq 1/8$.
+Hence, $(r_{0}^{s},r_{1}^{s},r_{2}^{s},r_{3}^{s})\in\widetilde{\mathcal{P}_{4}}$, and $\dim_{H}\mu_{\varphi}<1$.